Jefferson Lab HALOG ARCHIVES Logbook for December 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 308936 12/31/09 14:58 R. Michaels Compton DAQ work towards a trigger supervisor
  • 308935 12/24/09 06:03 auto_snap Left Arm HV snapshot
  • 308934 12/24/09 06:03 auto_snap Happex Charge Asym snapshot
  • 308933 12/24/09 06:03 auto_snap Hall A Tools snapshot
  • 308932 12/24/09 06:03 auto_snap Compton Polarimeter snapshot
  • 308931 12/24/09 06:03 auto_snap Beam Position and Current snapshot
  • 308930 12/24/09 06:03 auto_snap Accelerator Energies snapshot
  • 308929 12/24/09 02:03 auto_snap Left Arm HV snapshot
  • 308928 12/24/09 02:02 auto_snap Happex Charge Asym snapshot
  • 308927 12/24/09 02:02 auto_snap Hall A Tools snapshot
  • 308926 12/24/09 02:02 auto_snap Compton Polarimeter snapshot
  • 308925 12/24/09 02:02 auto_snap Beam Position and Current snapshot
  • 308924 12/24/09 02:02 auto_snap Accelerator Energies snapshot
  • Wednesday

  • 308923 12/23/09 22:02 auto_snap Left Arm HV snapshot
  • 308922 12/23/09 22:02 auto_snap Happex Charge Asym snapshot
  • 308921 12/23/09 22:02 auto_snap Hall A Tools snapshot
  • 308920 12/23/09 22:02 auto_snap Compton Polarimeter snapshot
  • 308919 12/23/09 22:02 auto_snap Beam Position and Current snapshot
  • 308918 12/23/09 22:02 auto_snap Accelerator Energies snapshot
  • 308917 12/23/09 18:02 auto_snap Left Arm HV snapshot
  • 308916 12/23/09 18:02 auto_snap Happex Charge Asym snapshot
  • 308915 12/23/09 18:02 auto_snap Hall A Tools snapshot
  • 308914 12/23/09 18:02 auto_snap Compton Polarimeter snapshot
  • 308913 12/23/09 18:02 auto_snap Beam Position and Current snapshot
  • 308912 12/23/09 18:02 auto_snap Accelerator Energies snapshot
  • 308911 12/23/09 16:47 R. Subedi Detector HV turned off
  • 308910 12/23/09 14:01 auto_snap Left Arm HV snapshot
  • 308909 12/23/09 14:01 auto_snap Happex Charge Asym snapshot
  • 308908 12/23/09 14:01 auto_snap Hall A Tools snapshot
  • 308907 12/23/09 14:01 auto_snap Compton Polarimeter snapshot
  • 308906 12/23/09 14:01 auto_snap Beam Position and Current snapshot
  • 308905 12/23/09 14:01 auto_snap Accelerator Energies snapshot
  • 308904 12/23/09 10:01 auto_snap Left Arm HV snapshot
  • 308903 12/23/09 10:01 auto_snap Happex Charge Asym snapshot
  • 308902 12/23/09 10:01 auto_snap Hall A Tools snapshot
  • 308901 12/23/09 10:01 auto_snap Compton Polarimeter snapshot
  • 308900 12/23/09 10:01 auto_snap Beam Position and Current snapshot
  • 308899 12/23/09 10:01 auto_snap Accelerator Energies snapshot
  • 308898 12/23/09 06:01 auto_snap Left Arm HV snapshot
  • 308897 12/23/09 06:01 auto_snap Happex Charge Asym snapshot
  • 308896 12/23/09 06:01 auto_snap Hall A Tools snapshot
  • 308895 12/23/09 06:01 auto_snap Compton Polarimeter snapshot
  • 308894 12/23/09 06:00 auto_snap Beam Position and Current snapshot
  • 308893 12/23/09 06:00 auto_snap Accelerator Energies snapshot
  • 308892 12/23/09 02:00 auto_snap Left Arm HV snapshot
  • 308891 12/23/09 02:00 auto_snap Happex Charge Asym snapshot
  • 308890 12/23/09 02:00 auto_snap Hall A Tools snapshot
  • 308889 12/23/09 02:00 auto_snap Compton Polarimeter snapshot
  • 308888 12/23/09 02:00 auto_snap Beam Position and Current snapshot
  • 308887 12/23/09 02:00 auto_snap Accelerator Energies snapshot
  • Tuesday

  • 308886 12/22/09 22:00 auto_snap Left Arm HV snapshot
  • 308885 12/22/09 22:00 auto_snap Happex Charge Asym snapshot
  • 308884 12/22/09 22:00 auto_snap Hall A Tools snapshot
  • 308883 12/22/09 22:00 auto_snap Compton Polarimeter snapshot
  • 308882 12/22/09 22:00 auto_snap Beam Position and Current snapshot
  • 308881 12/22/09 22:00 auto_snap Accelerator Energies snapshot
  • 308880 12/22/09 18:00 auto_snap Left Arm HV snapshot
  • 308879 12/22/09 18:00 auto_snap Happex Charge Asym snapshot
  • 308878 12/22/09 17:59 auto_snap Hall A Tools snapshot
  • 308877 12/22/09 17:59 auto_snap Compton Polarimeter snapshot
  • 308876 12/22/09 17:59 auto_snap Beam Position and Current snapshot
  • 308875 12/22/09 17:59 auto_snap Accelerator Energies snapshot
  • 308874 12/22/09 13:59 auto_snap Left Arm HV snapshot
  • 308873 12/22/09 13:59 auto_snap Happex Charge Asym snapshot
  • 308872 12/22/09 13:59 auto_snap Hall A Tools snapshot
  • 308871 12/22/09 13:59 auto_snap Compton Polarimeter snapshot
  • 308870 12/22/09 13:59 auto_snap Beam Position and Current snapshot
  • 308869 12/22/09 13:59 auto_snap Accelerator Energies snapshot
  • 308868 12/22/09 12:06 Party goers PVDIS end of run party now !
  • 308867 12/22/09 11:32 hsmith check list
  • 308866 12/22/09 09:59 auto_snap Left Arm HV snapshot
  • 308865 12/22/09 09:59 auto_snap Happex Charge Asym snapshot
  • 308864 12/22/09 09:59 auto_snap Hall A Tools snapshot
  • 308863 12/22/09 09:59 auto_snap Compton Polarimeter snapshot
  • 308862 12/22/09 09:59 auto_snap Beam Position and Current snapshot
  • 308861 12/22/09 09:59 auto_snap Accelerator Energies snapshot
  • 308860 12/22/09 08:14 Meekins Cryotarget Warm up
  • 308858 12/22/09 07:50 R. Subedi Hall in restricted access
  • 308857 12/22/09 07:14 Kai target alarm
  • 308856 12/22/09 05:59 auto_snap Left Arm HV snapshot
  • 308855 12/22/09 05:58 auto_snap Happex Charge Asym snapshot
  • 308854 12/22/09 05:58 auto_snap Hall A Tools snapshot
  • 308853 12/22/09 05:58 auto_snap Compton Polarimeter snapshot
  • 308852 12/22/09 05:58 auto_snap Beam Position and Current snapshot
  • 308851 12/22/09 05:58 auto_snap Accelerator Energies snapshot
  • 308850 12/22/09 05:16 Kai Pan move target to Empty
  • 308849 12/22/09 05:04 A. Camsonne Edet to garage
  • 308848 12/22/09 05:02 Holmstrom Owl Shift Summary
  • 308847 12/22/09 05:01 Kai Pan Experiment ends at 5:00am
  • 308846 12/22/09 05:00 Holmstrom Beam Gone
  • 308845 12/22/09 04:59 Holmstrom Turn off VDCs and HRSs
  • 308842 12/22/09 04:11 A. Camsonne Signal over background Compton
  • 308841 12/22/09 03:27 A. Camsonne Compton electron detector moved to position 12 mm
  • 308840 12/22/09 03:15 A. Camsonne Current Edet at 7 mm 10 uA
  • 308839 12/22/09 02:54 A. Camsonne Moving electron detector to position 7 mm
  • 308838 12/22/09 02:54 Holmstrom Q2 Measurements & Compton
  • 308836 12/22/09 02:53 apar Start_Parity_Run_14827__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,VDC on,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308834 12/22/09 02:53 apar End_of_Parity_Run_14827__CODA_Config=Pvdis
  • 308833 12/22/09 02:51 A. Camsonne Requested 10 uA by closing the slits
  • 308832 12/22/09 02:47 A. Camsonne Requested 50 uA
  • 308831 12/22/09 02:45 A. Camsonne Compton electron detector current at 17 mm 50 uA
  • 308830 12/22/09 02:42 apar End_of_Parity_Run_14826__CODA_Config=Pvdis
  • 308827 12/22/09 02:29 apar Start_Parity_Run_14826__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,VDC on,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308822 12/22/09 02:28 apar End_of_Parity_Run_14825__CODA_Config=Pvdis
  • 308821 12/22/09 02:15 A. Camsonne ETROC power cycled
  • 308820 12/22/09 02:02 A. Camsonne Compton electron detector moved to position 17 mm for check out
  • 308819 12/22/09 01:59 A. Camsonne Compton Electron detector current garage position
  • 308818 12/22/09 01:58 auto_snap Left Arm HV snapshot
  • 308817 12/22/09 01:58 auto_snap Happex Charge Asym snapshot
  • 308816 12/22/09 01:58 auto_snap Hall A Tools snapshot
  • 308815 12/22/09 01:58 auto_snap Compton Polarimeter snapshot
  • 308814 12/22/09 01:58 auto_snap Beam Position and Current snapshot
  • 308813 12/22/09 01:58 auto_snap Accelerator Energies snapshot
  • 308812 12/22/09 01:48 apar End_of_Parity_Run_14824__CODA_Config=Pvdis
  • 308811 12/22/09 01:41 Holmstrom Tools with VDCs
  • 308810 12/22/09 01:34 apar Start_Parity_Run_14825__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,VDC on,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308805 12/22/09 01:17 Holmstrom VDC
  • 308804 12/22/09 01:15 Holmstrom Turning on VDCs
  • 308801 12/22/09 00:34 apar Start_Parity_Run_14824__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308798 12/22/09 00:33 apar End_of_Parity_Run_14823__CODA_Config=Pvdis
  • 308796 12/22/09 00:02 Deng swing shift summary
  • Monday

  • 308795 12/21/09 23:50 Deng Important!!--All stuff disappeared in adaql2 machine suddenly
  • 308792 12/21/09 23:36 apar Start_Parity_Run_14823__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308791 12/21/09 22:50 apar Start_Parity_Run_14822__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308786 12/21/09 22:47 apar End_of_Parity_Run_14821__CODA_Config=Pvdis
  • 308785 12/21/09 22:05 Deng Beam current lowered to 105uA due to lower target heat power
  • 308784 12/21/09 22:02 narayan Target: Heater going to 0W
  • 308783 12/21/09 21:58 auto_snap Left Arm HV snapshot
  • 308782 12/21/09 21:58 auto_snap Happex Charge Asym snapshot
  • 308781 12/21/09 21:58 auto_snap Hall A Tools snapshot
  • 308780 12/21/09 21:58 auto_snap Compton Polarimeter snapshot
  • 308779 12/21/09 21:58 auto_snap Beam Position and Current snapshot
  • 308778 12/21/09 21:57 auto_snap Accelerator Energies snapshot
  • 308777 12/21/09 21:52 apar Start_Parity_Run_14821__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308772 12/21/09 21:47 apar End_of_Parity_Run_14820__CODA_Config=Pvdis
  • 308771 12/21/09 21:02 apar Start_Parity_Run_14820__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308766 12/21/09 21:00 apar End_of_Parity_Run_14819__CODA_Config=Pvdis
  • 308765 12/21/09 20:55 narayan Target: running good
  • 308764 12/21/09 20:11 narayan Compton rates and High Voltage
  • 308763 12/21/09 20:04 apar Start_Parity_Run_14819__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308758 12/21/09 20:01 apar End_of_Parity_Run_14818__CODA_Config=Pvdis
  • 308757 12/21/09 19:06 apar Start_Parity_Run_14818__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308752 12/21/09 19:02 apar End_of_Parity_Run_14817__CODA_Config=Pvdis
  • 308751 12/21/09 18:57 Deng hallA general tools and bcm script chart snap shot
  • 308750 12/21/09 18:07 apar Start_Parity_Run_14817__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308745 12/21/09 18:04 apar End_of_Parity_Run_14816__CODA_Config=Pvdis
  • 308744 12/21/09 17:57 auto_snap Left Arm HV snapshot
  • 308743 12/21/09 17:57 auto_snap Happex Charge Asym snapshot
  • 308742 12/21/09 17:57 auto_snap Hall A Tools snapshot
  • 308741 12/21/09 17:57 auto_snap Compton Polarimeter snapshot
  • 308740 12/21/09 17:57 auto_snap Beam Position and Current snapshot
  • 308739 12/21/09 17:57 auto_snap Accelerator Energies snapshot
  • 308738 12/21/09 17:55 A. Camsonne Compton electron detector rates still too high
  • 308737 12/21/09 17:09 apar Start_Parity_Run_14816__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308732 12/21/09 17:06 apar End_of_Parity_Run_14815__CODA_Config=Pvdis
  • 308731 12/21/09 16:10 apar Start_Parity_Run_14815__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308728 12/21/09 16:06 apar Halfwave plate IN, Slug 28
  • 308724 12/21/09 16:03 apar End_of_Parity_Run_14814__CODA_Config=Pvdis
  • 308723 12/21/09 16:03 jpchen Day shift summary
  • 308720 12/21/09 15:07 apar Start_Parity_Run_14814__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308717 12/21/09 15:06 apar End_of_Parity_Run_14813__CODA_Config=Pvdis
  • 308716 12/21/09 14:30 jpchen BCM/BPM charts
  • 308715 12/21/09 14:29 R. Michaels spot++
  • 308714 12/21/09 14:28 jpchen Tools screen
  • 308713 12/21/09 14:27 R. Michaels target snapshots
  • 308708 12/21/09 14:10 apar Start_Parity_Run_14813__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308707 12/21/09 14:09 apar End_of_Parity_Run_14812__CODA_Config=Pvdis
  • 308706 12/21/09 13:57 auto_snap Left Arm HV snapshot
  • 308705 12/21/09 13:57 auto_snap Happex Charge Asym snapshot
  • 308704 12/21/09 13:57 auto_snap Hall A Tools snapshot
  • 308703 12/21/09 13:57 auto_snap Compton Polarimeter snapshot
  • 308702 12/21/09 13:57 auto_snap Beam Position and Current snapshot
  • 308701 12/21/09 13:57 auto_snap Accelerator Energies snapshot
  • 308698 12/21/09 13:13 apar Start_Parity_Run_14812__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308695 12/21/09 13:12 apar End_of_Parity_Run_14811__CODA_Config=Pvdis
  • 308694 12/21/09 13:02 R. Michaels L-HRS effic and pi-rej
  • 308693 12/21/09 13:00 R. Michaels R-HRS effic and pi-rej
  • 308692 12/21/09 12:48 R. Michaels question re: time-diff macro plots
  • 308691 12/21/09 12:27 R. Michaels auto abu restarted
  • 308690 12/21/09 12:23 Armstrong CODA crash; restart
  • 308689 12/21/09 12:20 jpchen HV alarm, turned Compton HVs off
  • 308686 12/21/09 12:16 apar Start_Parity_Run_14811__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308683 12/21/09 12:10 apar End_of_Parity_Run_14809__CODA_Config=Pvdis
  • 308680 12/21/09 11:15 apar Start_Parity_Run_14809__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308677 12/21/09 11:13 apar End_of_Parity_Run_14808__CODA_Config=Pvdis
  • 308676 12/21/09 10:59 Armstrong Scalers snapshots
  • 308675 12/21/09 10:35 A. Camsonne Compton background not good at high current
  • 308672 12/21/09 10:18 apar Start_Parity_Run_14808__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=after target recovery complete,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308671 12/21/09 10:18 R. Michaels Qsq for R-HRS and L-HRS
  • 308670 12/21/09 10:15 jpchen spot++, Q2, now VDC off for production
  • 308669 12/21/09 10:13 A. Camsonne Compton rates OK at low current
  • 308660 12/21/09 09:57 auto_snap Left Arm HV snapshot
  • 308659 12/21/09 09:57 auto_snap Happex Charge Asym snapshot
  • 308658 12/21/09 09:57 auto_snap Hall A Tools snapshot
  • 308657 12/21/09 09:56 auto_snap Compton Polarimeter snapshot
  • 308656 12/21/09 09:56 auto_snap Beam Position and Current snapshot
  • 308655 12/21/09 09:56 auto_snap Accelerator Energies snapshot
  • 308654 12/21/09 09:13 jpchen Target ready.
  • 308653 12/21/09 09:10 Armstrong Target recovered
  • 308652 12/21/09 08:37 R. Michaels D2 vent for 10 sec
  • 308651 12/21/09 08:31 R. Michaels target woes. Move to Ta
  • 308650 12/21/09 08:23 Armstrong Left Collimation to Open
  • 308649 12/21/09 08:12 Kai sieve in "6 msr" now
  • 308648 12/21/09 08:06 Kai move sieve to be "6 msr 0.6188m"
  • 308646 12/21/09 08:04 Kai owl shift summary
  • 308645 12/21/09 07:36 Al Tobias Target Warming Up: 27.6K Little CryoFLow!!!!!!!!
  • 308644 12/21/09 07:35 Kai target warms up even in 2uA, call cryo again
  • 308635 12/21/09 07:16 apar End_of_Parity_Run_14807__CODA_Config=Pvdis
  • 308634 12/21/09 07:11 Kai turn on VDC to take 2uA beam
  • 308633 12/21/09 07:07 Al Tobias Target: Cryoflow from CHL (4K) decreasing to Hall A LD2 CryoTarget
  • 308631 12/21/09 07:00 Al Tobias Target: Cryo Flow is Lower! Beam OFF for now and I opened JT Valve on click to 66.9%
  • 308630 12/21/09 07:00 Kai we tried 80uA beam, but target still got problem
  • 308629 12/21/09 06:49 Al Tobias Target: CryoFlows Low!
  • 308628 12/21/09 06:42 Al Tobias Asked MCC to shut beam momentarily due to Target CryoFlow decreasing ...
  • 308622 12/21/09 06:06 apar End_of_Parity_Run_14806__CODA_Config=Pvdis
  • 308621 12/21/09 05:56 auto_snap Left Arm HV snapshot
  • 308620 12/21/09 05:56 auto_snap Happex Charge Asym snapshot
  • 308619 12/21/09 05:56 auto_snap Hall A Tools snapshot
  • 308618 12/21/09 05:56 auto_snap Compton Polarimeter snapshot
  • 308617 12/21/09 05:56 auto_snap Beam Position and Current snapshot
  • 308616 12/21/09 05:56 auto_snap Accelerator Energies snapshot
  • 308615 12/21/09 05:38 Al Tobias Target Screens
  • 308614 12/21/09 05:34 Kai Jin Huang saw the halog about sieve position and called in
  • 308613 12/21/09 05:22 R. Subedi Sieve OPEN
  • 308612 12/21/09 05:11 Kai update golden rootfile to be left #26414, right#5934
  • 308611 12/21/09 05:10 apar Start_Parity_Run_14806__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308606 12/21/09 05:08 apar End_of_Parity_Run_14805__CODA_Config=Pvdis
  • 308605 12/21/09 04:51 Kai Question about the position of the sieve
  • 308604 12/21/09 04:12 apar Start_Parity_Run_14805__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308599 12/21/09 04:09 apar End_of_Parity_Run_14804__CODA_Config=Pvdis
  • 308598 12/21/09 03:40 Al Tobias Target: JT Valve opened 1-step to 65.4%
  • 308597 12/21/09 03:39 Kai low beam cut setting
  • 308596 12/21/09 03:27 Kai Right arm all magnet ready
  • 308595 12/21/09 03:25 J. Rubin New Prescalers
  • 308594 12/21/09 03:13 J. Rubin spot++ (L-HRS run #26413)
  • 308593 12/21/09 03:13 M Friend Compton status
  • 308592 12/21/09 03:10 Al Tobias Moved Target to LD2
  • 308591 12/21/09 03:07 Kai magnet status on the right arm
  • 308589 12/21/09 03:05 Kai finish calibration, go to production
  • 308587 12/21/09 02:55 apar Start_Parity_Run_14804__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308585 12/21/09 02:41 apar End_of_Parity_Run_14803__CODA_Config=Pvdis
  • 308583 12/21/09 02:41 Kai global electron rate
  • 308582 12/21/09 02:39 Kai Left scaler3,4
  • 308581 12/21/09 02:38 Kai Left scaler1,2 rate reading
  • 308580 12/21/09 02:37 Kai sieve pattern for #26411, sieve in, raster on 6*6, LD2
  • 308579 12/21/09 02:13 Al Tobias Target Charts
  • 308577 12/21/09 02:06 apar Start_Parity_Run_14803__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm Al Dummy
  • 308575 12/21/09 02:02 Kai right arm Q2 quenched, while Dipole is not recovered yet also
  • 308574 12/21/09 01:58 Al Tobias Target moved to Aluminum DUMMY
  • 308573 12/21/09 01:56 auto_snap Left Arm HV snapshot
  • 308572 12/21/09 01:56 auto_snap Happex Charge Asym snapshot
  • 308571 12/21/09 01:56 auto_snap Hall A Tools snapshot
  • 308570 12/21/09 01:56 auto_snap Compton Polarimeter snapshot
  • 308569 12/21/09 01:56 auto_snap Beam Position and Current snapshot
  • 308568 12/21/09 01:56 auto_snap Accelerator Energies snapshot
  • 308567 12/21/09 01:51 Kai turn off the VDC, move sieve out
  • 308564 12/21/09 01:45 Kai sieve patter for #26410, sieve in, raster off, carbon
  • 308561 12/21/09 01:32 Al Tobias Moved Target to LD2 w/ Beam Compensation ON
  • 308560 12/21/09 01:30 Al Tobias I turned on Beam Compensation
  • 308559 12/21/09 01:29 Kai finish carbon calibration run, go to LD2
  • 308558 12/21/09 01:26 Al Tobias I turned on Beam Compensation
  • 308555 12/21/09 01:21 Kai sieve pattern for #26409, raster on
  • 308554 12/21/09 01:07 Al Tobias Target: Does Beam Compensation have to be turned ON?
  • 308553 12/21/09 01:03 Al Tobias LD2 Temps & CryoFlows
  • 308548 12/21/09 00:43 Kai snapshot of hall A GT
  • 308547 12/21/09 00:25 Al Tobias LD2 Temp & CryoFlows
  • 308546 12/21/09 00:25 Kai move sieve IN
  • 308541 12/21/09 00:15 Al Tobias Target Moved to OPTICS
  • 308538 12/21/09 00:11 R. Subedi Correct Q2 for left HRS
  • 308537 12/21/09 00:04 apar End_of_Parity_Run_14802__CODA_Config=Pvdis
  • 308534 12/21/09 00:01 ZAFAR AHMED Owl shift summary
  • Sunday

  • 308533 12/20/09 23:53 R. Michaels Q2 for L-HRS
  • 308531 12/20/09 23:44 R. Subedi Spot check with the Q2 run
  • 308530 12/20/09 23:43 R. Michaels beam recovery procedure
  • 308523 12/20/09 23:37 R. Subedi left Q2 tripped 2nd time
  • 308522 12/20/09 23:29 R. Subedi spot++ looks good
  • 308521 12/20/09 23:28 R. Michaels put in LD2 target
  • 308518 12/20/09 22:48 R. Michaels aiming towards production: magnets, etc
  • 308517 12/20/09 22:41 apar Start_Parity_Run_14802__CODA_Config=Pvdis_Run_type=Test,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308516 12/20/09 22:39 R. Michaels Cavity studies, started, then aborted
  • 308515 12/20/09 22:30 apar Start_Parity_Run_14801__CODA_Config=Pvdis_Run_type=Test,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308513 12/20/09 22:28 ZAFAR AHMED All pregains of cavities changed to 1
  • 308512 12/20/09 22:28 apar End_of_Parity_Run_14800__CODA_Config=Pvdis
  • 308511 12/20/09 22:28 R. Michaels reduce JT valve
  • 308510 12/20/09 22:16 R. Michaels cryoflows very unstable
  • 308509 12/20/09 22:10 R. Michaels access to turn on heaters
  • 308508 12/20/09 22:04 R. Michaels sudden drop of target temp !
  • 308507 12/20/09 21:55 auto_snap Left Arm HV snapshot
  • 308506 12/20/09 21:55 auto_snap Happex Charge Asym snapshot
  • 308505 12/20/09 21:55 auto_snap Hall A Tools snapshot
  • 308504 12/20/09 21:55 auto_snap Compton Polarimeter snapshot
  • 308503 12/20/09 21:55 auto_snap Beam Position and Current snapshot
  • 308502 12/20/09 21:55 auto_snap Accelerator Energies snapshot
  • 308501 12/20/09 21:50 R. Michaels Cavity studies, part 0
  • 308499 12/20/09 21:41 apar Start_Parity_Run_14800__CODA_Config=Pvdis_Run_type=Test,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308497 12/20/09 21:33 R. Michaels cavity scan result and plan
  • 308496 12/20/09 21:30 R. Subedi Run plan tonight , status update: hrs magnets ready, target need to wait
  • 308495 12/20/09 21:11 apar End_of_Parity_Run_14799__CODA_Config=Pvdis
  • 308492 12/20/09 20:48 Jack (Techs) Magnets
  • 308491 12/20/09 20:28 ZAFAR AHMED Cryo level in magnets is good.
  • 308490 12/20/09 19:23 ZAFAR AHMED Target will be ready by tomorrow morning
  • 308489 12/20/09 19:15 Bo Target Status Update
  • 308488 12/20/09 19:01 ZAFAR AHMED Target
  • 308487 12/20/09 18:33 apar Start_Parity_Run_14799__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308484 12/20/09 18:31 apar End_of_Parity_Run_14798__CODA_Config=Pvdis
  • 308481 12/20/09 18:12 R. Subedi Status Update: things look promising
  • 308480 12/20/09 17:55 auto_snap Left Arm HV snapshot
  • 308479 12/20/09 17:55 auto_snap Happex Charge Asym snapshot
  • 308478 12/20/09 17:55 auto_snap Hall A Tools snapshot
  • 308477 12/20/09 17:55 auto_snap Compton Polarimeter snapshot
  • 308476 12/20/09 17:55 auto_snap Beam Position and Current snapshot
  • 308475 12/20/09 17:55 auto_snap Accelerator Energies snapshot
  • 308474 12/20/09 17:13 R. Subedi Cause of Cryo-problem (copied from ELOG): the culprit of longer down time
  • 308471 12/20/09 16:29 R. Michaels amplifiers added for cavities
  • 308470 12/20/09 16:15 B. Quinn Day shift summary
  • 308469 12/20/09 16:14 Jack (Techs) Right Arm vacuum pump
  • 308468 12/20/09 16:07 Jack (Techs) Two problems
  • 308466 12/20/09 13:55 auto_snap Left Arm HV snapshot
  • 308465 12/20/09 13:55 auto_snap Happex Charge Asym snapshot
  • 308464 12/20/09 13:55 auto_snap Hall A Tools snapshot
  • 308463 12/20/09 13:55 auto_snap Compton Polarimeter snapshot
  • 308462 12/20/09 13:54 auto_snap Beam Position and Current snapshot
  • 308461 12/20/09 13:54 auto_snap Accelerator Energies snapshot
  • 308460 12/20/09 09:54 auto_snap Left Arm HV snapshot
  • 308459 12/20/09 09:54 auto_snap Happex Charge Asym snapshot
  • 308458 12/20/09 09:54 auto_snap Hall A Tools snapshot
  • 308457 12/20/09 09:54 auto_snap Compton Polarimeter snapshot
  • 308456 12/20/09 09:54 auto_snap Beam Position and Current snapshot
  • 308455 12/20/09 09:54 auto_snap Accelerator Energies snapshot
  • 308454 12/20/09 09:30 R. Subedi Status Update of Cryo-related problem
  • 308453 12/20/09 08:22 apar Start_Parity_Run_14798__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308451 12/20/09 08:07 Kai Pan owl shift summary
  • 308450 12/20/09 07:38 Al Tobias Target 56K+ asked for Controlled Access
  • 308449 12/20/09 07:27 Al Tobias Target still warm: 38K+
  • 308448 12/20/09 06:46 R. Michaels Low-I cavity tests, part III
  • 308447 12/20/09 06:45 apar End_of_Parity_Run_14797__CODA_Config=Pvdis
  • 308435 12/20/09 06:27 R. Subedi Magnet cryo-problem (Tech)
  • 308433 12/20/09 06:25 Al Tobias Nope, LD2 Target not quite back
  • 308431 12/20/09 06:19 apar Start_Parity_Run_14797__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=off,comment_text=Low current cavity test (@ 0.1uA, Raster Off),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308429 12/20/09 06:17 apar End_of_Parity_Run_14796__CODA_Config=Pvdis
  • 308427 12/20/09 06:16 R. Michaels Low-I cavity tests, part II
  • 308418 12/20/09 06:01 Al Tobias Target is coming back: LD2 23.2K and lowering.
  • 308415 12/20/09 05:58 apar Start_Parity_Run_14796__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=off,comment_text=Low current cavity test (@ 0.5uA, Raster Off),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308411 12/20/09 05:54 auto_snap Left Arm HV snapshot
  • 308410 12/20/09 05:54 apar End_of_Parity_Run_14795__CODA_Config=Pvdis
  • 308409 12/20/09 05:54 auto_snap Happex Charge Asym snapshot
  • 308408 12/20/09 05:54 auto_snap Hall A Tools snapshot
  • 308407 12/20/09 05:54 auto_snap Compton Polarimeter snapshot
  • 308406 12/20/09 05:54 auto_snap Beam Position and Current snapshot
  • 308405 12/20/09 05:54 auto_snap Accelerator Energies snapshot
  • 308401 12/20/09 05:49 Kai all magnets trip
  • 308397 12/20/09 05:45 Al Tobias Turned on Target Heater/PID and fan=15Hz
  • 308394 12/20/09 05:40 Al Tobias Target Cooling Down?
  • 308390 12/20/09 05:34 Al Tobias Leaving LD2 Target alone for Now: Still no CryoFlow
  • 308389 12/20/09 05:32 R. Michaels Low-I cavity tests, part I
  • 308379 12/20/09 04:48 apar Start_Parity_Run_14795__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=off,comment_text=Low current cavity test (@ 1uA, Raster Off),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308375 12/20/09 04:42 apar End_of_Parity_Run_14794__CODA_Config=Pvdis
  • 308372 12/20/09 04:38 apar Start_Parity_Run_14794__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Low current cavity test (@ 1uA),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308371 12/20/09 04:37 apar End_of_Parity_Run_14793__CODA_Config=Pvdis
  • 308369 12/20/09 04:32 apar Start_Parity_Run_14793__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Low current cavity test,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308367 12/20/09 04:11 Al Tobias Target CryoFlow situation
  • 308366 12/20/09 04:00 apar End_of_Parity_Run_14792__CODA_Config=Pvdis
  • 308365 12/20/09 03:51 R. Michaels access request to reboot cavity crate
  • 308364 12/20/09 03:47 R. Michaels Lumi Oversampling Run Summary
  • 308363 12/20/09 03:43 apar Start_Parity_Run_14792__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (no oversampling -- baseline),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308362 12/20/09 03:42 R. Michaels Return to Ovs=0. Run 14972
  • 308361 12/20/09 03:41 apar End_of_Parity_Run_14791__CODA_Config=Pvdis
  • 308360 12/20/09 03:28 apar Start_Parity_Run_14791__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 10),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308359 12/20/09 03:27 R. Michaels run 14791 with ovs=9
  • 308358 12/20/09 03:26 apar End_of_Parity_Run_14790__CODA_Config=Pvdis
  • 308357 12/20/09 03:25 R. Michaels sometimes wrong oversamp in db
  • 308356 12/20/09 03:11 apar End_of_Parity_Run_14787__CODA_Config=Pvdis
  • 308355 12/20/09 03:11 R. Michaels Ovs=7 run 14790
  • 308354 12/20/09 03:10 apar Start_Parity_Run_14790__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 8),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308353 12/20/09 03:08 apar End_of_Parity_Run_14789__CODA_Config=Pvdis
  • 308352 12/20/09 03:07 R. Michaels Ovs=1 run 14789
  • 308351 12/20/09 03:06 Al Tobias Cryoflow to Target: Update
  • 308350 12/20/09 02:58 apar Start_Parity_Run_14789__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 10),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308349 12/20/09 02:57 apar End_of_Parity_Run_14788__CODA_Config=Pvdis
  • 308348 12/20/09 02:56 R. Michaels Ovs=9 runs, cannot revive BMW
  • 308347 12/20/09 02:50 apar Start_Parity_Run_14788__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 10),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308346 12/20/09 02:40 apar Start_Parity_Run_14787__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 10),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308345 12/20/09 02:38 apar End_of_Parity_Run_14786__CODA_Config=Pvdis
  • 308344 12/20/09 02:33 apar Start_Parity_Run_14786__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling factor 10),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308343 12/20/09 02:32 apar End_of_Parity_Run_14785__CODA_Config=Pvdis
  • 308342 12/20/09 02:31 R. Michaels Run 14784 with ovs=3
  • 308341 12/20/09 02:19 apar Start_Parity_Run_14785__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling '3'),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308340 12/20/09 02:16 apar End_of_Parity_Run_14784__CODA_Config=Pvdis
  • 308339 12/20/09 02:13 apar Start_Parity_Run_14784__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target (oversampling '3'),dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308338 12/20/09 02:10 apar End_of_Parity_Run_14783__CODA_Config=Pvdis
  • 308337 12/20/09 02:08 R. Michaels ovs = 0 run 14783
  • 308336 12/20/09 01:57 apar Start_Parity_Run_14783__CODA_Config=Pvdis_Run_type=Other,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Oversampling test on tantalum target,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=Thick Tantalum
  • 308335 12/20/09 01:54 auto_snap Left Arm HV snapshot
  • 308334 12/20/09 01:54 auto_snap Happex Charge Asym snapshot
  • 308333 12/20/09 01:53 auto_snap Hall A Tools snapshot
  • 308332 12/20/09 01:53 auto_snap Compton Polarimeter snapshot
  • 308331 12/20/09 01:53 auto_snap Beam Position and Current snapshot
  • 308330 12/20/09 01:53 auto_snap Accelerator Energies snapshot
  • 308329 12/20/09 01:49 R. Michaels near-term run plan
  • 308328 12/20/09 01:48 Al Tobias Target is at Thick Tantalum
  • 308327 12/20/09 01:00 Al Tobias Target Warm: Return JT Valve to original running setting
  • 308326 12/20/09 00:54 jpchen cryo flow dropped from 9 g/s to near 0 ~10 pm
  • 308325 12/20/09 00:50 R. Subedi Run Coordinator Report for Dec 19, 2009
  • 308324 12/20/09 00:38 jpchen Target temperature back to around 50K
  • 308323 12/20/09 00:14 jpchen target status and vacuum
  • Saturday

  • 308321 12/19/09 23:59 Armstrong Shift Summary - Swing shift
  • 308320 12/19/09 23:29 jpchen turned off fan adn HP heater.
  • 308319 12/19/09 23:10 jpchen What to do when cryo flow become unstable and dropping
  • 308318 12/19/09 22:45 Armstrong target status.
  • 308317 12/19/09 22:30 Armstrong Target has boiled off
  • 308316 12/19/09 22:20 silwal haD2_P_Ballast_R alarm
  • 308315 12/19/09 22:17 silwal cryo flow rate below 2g/s, target temp close to 50 K, reduced the fan freq to 15 Hz
  • 308310 12/19/09 22:02 silwal cryo flow rate down to about 2.5g/s from 4.0
  • 308309 12/19/09 22:02 R. Subedi Slug 27 started
  • 308308 12/19/09 22:00 silwal haScatt_Chamber_Vac_R alarms: couple of times
  • 308307 12/19/09 21:58 silwal target moved to optics
  • 308304 12/19/09 21:53 auto_snap Left Arm HV snapshot
  • 308303 12/19/09 21:53 auto_snap Happex Charge Asym snapshot
  • 308302 12/19/09 21:53 auto_snap Hall A Tools snapshot
  • 308301 12/19/09 21:53 auto_snap Compton Polarimeter snapshot
  • 308300 12/19/09 21:53 auto_snap Beam Position and Current snapshot
  • 308299 12/19/09 21:53 auto_snap Accelerator Energies snapshot
  • 308292 12/19/09 21:42 silwal target moved to dummy aluminum
  • 308289 12/19/09 21:28 Armstrong New Magnet settings
  • 308286 12/19/09 21:21 silwal target moved to LD2
  • 308269 12/19/09 19:33 silwal target charts: cryo flow rate fluctuations (due to fluctuation in weather conditions)
  • 308268 12/19/09 19:32 silwal target to optics (carbon)
  • 308263 12/19/09 19:19 apar End_of_Parity_Run_14782__CODA_Config=Pvdis
  • 308260 12/19/09 18:40 apar Start_Parity_Run_14782__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=left arm Q1 and D at 3.66 GeV/c. Q2 and Q3 at 3.3a49 GeV/c, right arm Q1 at 3.13 GeV/c, other magnets in rhrs at 3.1 GeV/c,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308259 12/19/09 18:37 A. Camsonne Compton EPICS reboot
  • 308258 12/19/09 18:35 Jack (Techs) Gas Shed brief alarms
  • 308257 12/19/09 18:25 apar Start_Parity_Run_14781__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=left arm Q1 and D at 3.66 GeV/c. Q2 and Q3 at 3.3a49 GeV/c, right arm Q1 at 3.13 GeV/c, other magnets in rhrs at 3.1 GeV/c,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308254 12/19/09 18:22 Armstrong New HRS magnet settings
  • 308253 12/19/09 18:16 apar End_of_Parity_Run_14780__CODA_Config=Pvdis
  • 308250 12/19/09 17:53 auto_snap Left Arm HV snapshot
  • 308249 12/19/09 17:53 auto_snap Happex Charge Asym snapshot
  • 308248 12/19/09 17:53 auto_snap Hall A Tools snapshot
  • 308247 12/19/09 17:53 auto_snap Compton Polarimeter snapshot
  • 308246 12/19/09 17:53 auto_snap Beam Position and Current snapshot
  • 308245 12/19/09 17:53 auto_snap Accelerator Energies snapshot
  • 308244 12/19/09 17:40 A. Camsonne Compton background rates high
  • 308241 12/19/09 17:35 apar Start_Parity_Run_14780__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3660,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308240 12/19/09 17:31 A. Camsonne Compton electron detector current position 7mm 2 uA
  • 308239 12/19/09 17:29 Armstrong Q2 done, go to 100 uA
  • 308236 12/19/09 17:25 R. Subedi Sieve pattern with raster OFF
  • 308235 12/19/09 17:22 silwal target graphs
  • 308234 12/19/09 17:22 R. Subedi Sieve OUT
  • 308231 12/19/09 17:21 silwal moved target to LD
  • 308228 12/19/09 16:57 R. Subedi Sieve slit from 26389
  • 308222 12/19/09 16:13 R. Subedi Sieve IN
  • 308213 12/19/09 16:00 A. Deur Day shift summary
  • 308208 12/19/09 15:54 Kiadtisak target snapshot
  • 308207 12/19/09 15:51 R. Subedi New calibration kin begins
  • 308206 12/19/09 15:41 Kiadtisak Target moved to optics
  • 308203 12/19/09 15:39 apar End_of_Parity_Run_14779__CODA_Config=Pvdis
  • 308198 12/19/09 15:11 apar Start_Parity_Run_14779__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308197 12/19/09 15:10 apar End_of_Parity_Run_14778__CODA_Config=Pvdis
  • 308194 12/19/09 14:14 apar Start_Parity_Run_14778__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308191 12/19/09 14:13 apar End_of_Parity_Run_14777__CODA_Config=Pvdis
  • 308190 12/19/09 13:52 auto_snap Left Arm HV snapshot
  • 308189 12/19/09 13:52 auto_snap Happex Charge Asym snapshot
  • 308188 12/19/09 13:52 auto_snap Hall A Tools snapshot
  • 308187 12/19/09 13:52 auto_snap Compton Polarimeter snapshot
  • 308186 12/19/09 13:52 auto_snap Beam Position and Current snapshot
  • 308185 12/19/09 13:52 auto_snap Accelerator Energies snapshot
  • 308182 12/19/09 13:16 apar Start_Parity_Run_14777__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308179 12/19/09 13:15 apar End_of_Parity_Run_14776__CODA_Config=Pvdis
  • 308176 12/19/09 12:19 apar Start_Parity_Run_14776__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308173 12/19/09 12:18 apar End_of_Parity_Run_14775__CODA_Config=Pvdis
  • 308172 12/19/09 11:46 R. Subedi Left hrs coda fine now
  • 308170 12/19/09 11:41 R. Subedi Left hrs coda problem, trying to get it back
  • 308165 12/19/09 11:22 apar Start_Parity_Run_14775__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308162 12/19/09 11:20 apar End_of_Parity_Run_14774__CODA_Config=Pvdis
  • 308161 12/19/09 11:11 A. Deur Automatic BTA not working
  • 308159 12/19/09 10:53 Kiadtisak Target fill pressure
  • 308158 12/19/09 10:44 Kiadtisak Sieve pattern for 26334
  • 308155 12/19/09 10:24 apar Start_Parity_Run_14774__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308152 12/19/09 10:23 apar End_of_Parity_Run_14773__CODA_Config=Pvdis
  • 308151 12/19/09 10:04 A. Deur screenshot of left scaler display
  • 308150 12/19/09 10:02 A. deur DAQ problem and left scalers
  • 308149 12/19/09 10:01 R. Subedi Left scalers back from remote reset of TS
  • 308148 12/19/09 09:52 auto_snap Left Arm HV snapshot
  • 308147 12/19/09 09:52 auto_snap Happex Charge Asym snapshot
  • 308146 12/19/09 09:52 auto_snap Hall A Tools snapshot
  • 308145 12/19/09 09:52 auto_snap Compton Polarimeter snapshot
  • 308144 12/19/09 09:52 auto_snap Beam Position and Current snapshot
  • 308143 12/19/09 09:52 auto_snap Accelerator Energies snapshot
  • 308139 12/19/09 09:27 apar Start_Parity_Run_14773__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308136 12/19/09 09:26 apar End_of_Parity_Run_14772__CODA_Config=Pvdis
  • 308135 12/19/09 08:57 A. Deur Left Scaler problem
  • 308134 12/19/09 08:42 R. Michaels re: left arm scalers
  • 308133 12/19/09 08:39 A. Deur beam stripchart snapshot
  • 308128 12/19/09 08:33 A. Deur Hall A tools
  • 308127 12/19/09 08:31 Kiadtisak Target Screenshot
  • 308126 12/19/09 08:30 apar Start_Parity_Run_14772__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308125 12/19/09 08:28 apar End_of_Parity_Run_14771__CODA_Config=Pvdis
  • 308124 12/19/09 08:23 A. Deur routine spot++
  • 308123 12/19/09 07:57 Kai owl shift summary
  • 308121 12/19/09 07:45 Al Tobias Target Screens
  • 308120 12/19/09 07:41 Al Tobias Gas Shed Airflow Alarm again ...
  • 308119 12/19/09 07:32 apar Start_Parity_Run_14771__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308114 12/19/09 07:31 apar End_of_Parity_Run_14770__CODA_Config=Pvdis
  • 308113 12/19/09 07:02 Al Tobias Gas Shed Airflow Alarm
  • 308110 12/19/09 06:35 apar Start_Parity_Run_14770__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308107 12/19/09 06:34 apar End_of_Parity_Run_14769__CODA_Config=Pvdis
  • 308106 12/19/09 05:52 auto_snap Left Arm HV snapshot
  • 308105 12/19/09 05:52 auto_snap Happex Charge Asym snapshot
  • 308104 12/19/09 05:52 auto_snap Hall A Tools snapshot
  • 308103 12/19/09 05:51 auto_snap Compton Polarimeter snapshot
  • 308102 12/19/09 05:51 auto_snap Beam Position and Current snapshot
  • 308101 12/19/09 05:51 auto_snap Accelerator Energies snapshot
  • 308100 12/19/09 05:39 apar Start_Parity_Run_14769__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308099 12/19/09 05:27 apar Start_Parity_Run_14768__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308094 12/19/09 05:26 apar End_of_Parity_Run_14767__CODA_Config=Pvdis
  • 308093 12/19/09 05:04 Al Tobias Target Status: Loop1 Cryo
  • 308092 12/19/09 04:58 Al Tobias Compton DAQ reboot
  • 308091 12/19/09 04:30 apar Start_Parity_Run_14767__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308086 12/19/09 04:28 apar End_of_Parity_Run_14766__CODA_Config=Pvdis
  • 308085 12/19/09 04:10 J. Rubin spot++
  • 308084 12/19/09 04:02 Al Tobias Target Fill Pressure just dropped 1.5psi
  • 308083 12/19/09 03:33 R. Subedi Left xscaler not responding, data look ok
  • 308082 12/19/09 03:33 apar Start_Parity_Run_14766__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308077 12/19/09 03:31 apar End_of_Parity_Run_14765__CODA_Config=Pvdis
  • 308075 12/19/09 03:18 Al Tobias Detectors Flammable Gas Sensors
  • 308071 12/19/09 02:47 A. Camsonne Compton electron detector current position 10 mm 100 uA
  • 308070 12/19/09 02:35 apar Start_Parity_Run_14765__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308067 12/19/09 02:34 Kai flip the HWP from IN to OUT
  • 308066 12/19/09 02:33 apar Halfwave plate OUT, Slug 26
  • 308065 12/19/09 02:30 J. Rubin L-HRS showing 100% deadtime, but doesn't make sense...
  • 308064 12/19/09 02:30 apar End_of_Parity_Run_14764__CODA_Config=Pvdis
  • 308061 12/19/09 02:05 apar Start_Parity_Run_14764__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308057 12/19/09 01:56 Kai (not GC & SC) is a little wider than before
  • 308053 12/19/09 01:54 apar End_of_Parity_Run_14763__CODA_Config=Pvdis
  • 308052 12/19/09 01:51 auto_snap Left Arm HV snapshot
  • 308051 12/19/09 01:51 auto_snap Happex Charge Asym snapshot
  • 308050 12/19/09 01:51 auto_snap Hall A Tools snapshot
  • 308049 12/19/09 01:51 auto_snap Compton Polarimeter snapshot
  • 308048 12/19/09 01:51 auto_snap Beam Position and Current snapshot
  • 308047 12/19/09 01:51 auto_snap Accelerator Energies snapshot
  • 308046 12/19/09 01:45 Al Tobias Target Screens
  • 308045 12/19/09 01:31 Al Tobias Compton VNC Viewer
  • 308044 12/19/09 00:59 apar Start_Parity_Run_14763__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308039 12/19/09 00:57 apar End_of_Parity_Run_14762__CODA_Config=Pvdis
  • 308038 12/19/09 00:10 Kai hall A tool and beam chart
  • 308036 12/19/09 00:01 apar Start_Parity_Run_14762__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308035 12/19/09 00:00 B. Quinn Swing shift summary
  • Friday

  • 308030 12/18/09 23:59 apar End_of_Parity_Run_14761__CODA_Config=Pvdis
  • 308029 12/18/09 23:04 apar Start_Parity_Run_14761__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308024 12/18/09 23:02 apar End_of_Parity_Run_14760__CODA_Config=Pvdis
  • 308023 12/18/09 22:35 Deng Left/Right HRS DAQ crashed
  • 308019 12/18/09 22:07 apar Start_Parity_Run_14760__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308016 12/18/09 22:06 apar End_of_Parity_Run_14759__CODA_Config=Pvdis
  • 308015 12/18/09 21:51 auto_snap Left Arm HV snapshot
  • 308014 12/18/09 21:51 auto_snap Happex Charge Asym snapshot
  • 308013 12/18/09 21:51 auto_snap Hall A Tools snapshot
  • 308012 12/18/09 21:51 auto_snap Compton Polarimeter snapshot
  • 308011 12/18/09 21:51 auto_snap Beam Position and Current snapshot
  • 308010 12/18/09 21:51 auto_snap Accelerator Energies snapshot
  • 308009 12/18/09 21:10 apar Start_Parity_Run_14759__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 308004 12/18/09 21:08 apar End_of_Parity_Run_14758__CODA_Config=Pvdis
  • 308003 12/18/09 21:08 Deng/Ramesh Right Q1 P0 reduced to 3.10083 GeV/c
  • 308002 12/18/09 20:44 B. Quinn Hall A tools screen
  • 308001 12/18/09 20:39 R. Subedi Run Coordinator Report for Dec 17th Day -18th Swing shift, 2009
  • 308000 12/18/09 20:24 apar Start_Parity_Run_14758__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307995 12/18/09 20:22 apar End_of_Parity_Run_14757__CODA_Config=Pvdis
  • 307994 12/18/09 19:56 aniol target screens
  • 307993 12/18/09 19:27 apar Start_Parity_Run_14757__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307988 12/18/09 19:25 apar End_of_Parity_Run_14756__CODA_Config=Pvdis
  • 307987 12/18/09 19:01 Deng Lumi HV adjusted for 4 pass
  • 307986 12/18/09 18:38 Deng Beam Modulation (dithering) is on now for production
  • 307985 12/18/09 18:33 Deng Q^2 and W plots
  • 307982 12/18/09 18:30 apar Start_Parity_Run_14756__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307979 12/18/09 18:21 Deng spot ++ check, LD2, 2uA, 6x6
  • 307974 12/18/09 18:14 Deng spot++, LD2, 2uA, 4x4mm^2
  • 307971 12/18/09 18:02 Deng spot++ for BeO target, raster size 4x4
  • 307968 12/18/09 17:50 auto_snap Left Arm HV snapshot
  • 307967 12/18/09 17:50 auto_snap Happex Charge Asym snapshot
  • 307966 12/18/09 17:50 auto_snap Hall A Tools snapshot
  • 307965 12/18/09 17:50 auto_snap Compton Polarimeter snapshot
  • 307964 12/18/09 17:50 auto_snap Beam Position and Current snapshot
  • 307963 12/18/09 17:50 auto_snap Accelerator Energies snapshot
  • 307962 12/18/09 17:33 R. Subedi Access finished, right Q1 is coming back ON
  • 307961 12/18/09 17:32 M Friend Compton webcam/apc reboot
  • 307960 12/18/09 17:32 Deng VDC on, right Q1 came back on
  • 307959 12/18/09 17:20 dalton Doing a QWP scan
  • 307958 12/18/09 17:18 dalton Compton laser alignment tweak
  • 307957 12/18/09 17:10 B. Quinn Hall access
  • 307956 12/18/09 17:08 O. Glamazdin Moller results
  • 307955 12/18/09 16:58 R. Subedi Moller done, escorted controlled access to reset right HRS Q1
  • 307954 12/18/09 16:56 dalton Compton laser locking
  • 307953 12/18/09 16:45 Deng Beam modulation script dead, reboot roc23
  • 307951 12/18/09 16:01 W. Deconinck Day shift summary
  • 307950 12/18/09 15:38 apar End_of_Parity_Run_14747__CODA_Config=Pvdis
  • 307949 12/18/09 14:42 M Friend Compton machine reboot
  • 307948 12/18/09 14:04 W. Deconinck Beam is back, continuing Moller setup
  • 307947 12/18/09 13:50 auto_snap Left Arm HV snapshot
  • 307946 12/18/09 13:50 auto_snap Happex Charge Asym snapshot
  • 307945 12/18/09 13:50 auto_snap Hall A Tools snapshot
  • 307944 12/18/09 13:50 auto_snap Compton Polarimeter snapshot
  • 307943 12/18/09 13:50 auto_snap Beam Position and Current snapshot
  • 307942 12/18/09 13:50 auto_snap Accelerator Energies snapshot
  • 307941 12/18/09 13:11 W. Deconinck MCC down for half an hour
  • 307940 12/18/09 12:48 D.Wang target moved to empty position
  • 307939 12/18/09 12:44 M Friend Compton tune
  • 307938 12/18/09 12:41 M Friend Compton epics reboot
  • 307936 12/18/09 12:39 W. Deconinck Starting Moller measurement procedure
  • 307935 12/18/09 12:39 apar End_of_Parity_Run_14755__CODA_Config=Pvdis
  • 307934 12/18/09 12:29 W. Deconinck R Q1 will be reset after Moller measurement
  • 307932 12/18/09 12:18 W. Deconinck Right Q1 problem (comm error)
  • 307931 12/18/09 12:18 apar Start_Parity_Run_14755__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=Right Q1 is not working,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307928 12/18/09 12:14 apar End_of_Parity_Run_14754__CODA_Config=Pvdis
  • 307927 12/18/09 12:13 D.Wang target status and strip chats
  • 307926 12/18/09 11:41 apar Start_Parity_Run_14754__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307921 12/18/09 11:39 apar End_of_Parity_Run_14753__CODA_Config=Pvdis
  • 307916 12/18/09 10:57 D.Wang golden run updated, 26344/5847
  • 307915 12/18/09 10:44 apar Start_Parity_Run_14753__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307912 12/18/09 10:42 apar End_of_Parity_Run_14752__CODA_Config=Pvdis
  • 307911 12/18/09 10:14 apar Start_Parity_Run_14752__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307910 12/18/09 10:13 W. Deconinck MCC starts some invasive injector tuning (est: 15 min)
  • 307909 12/18/09 10:11 apar End_of_Parity_Run_14751__CODA_Config=Pvdis
  • 307908 12/18/09 10:06 D.Wang Electron rates and asymmetries for Al dummy run
  • 307907 12/18/09 09:54 apar Start_Parity_Run_14751__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=reset ROC 23,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307906 12/18/09 09:52 apar End_of_Parity_Run_14750__CODA_Config=Pvdis
  • 307905 12/18/09 09:50 auto_snap Left Arm HV snapshot
  • 307904 12/18/09 09:50 auto_snap Happex Charge Asym snapshot
  • 307903 12/18/09 09:50 auto_snap Hall A Tools snapshot
  • 307902 12/18/09 09:50 auto_snap Compton Polarimeter snapshot
  • 307901 12/18/09 09:49 auto_snap Beam Position and Current snapshot
  • 307900 12/18/09 09:49 auto_snap Accelerator Energies snapshot
  • 307899 12/18/09 09:16 D.Wang Reset ROC 23 and did epics reboot
  • 307898 12/18/09 09:04 W. Deconinck EPICS rebooted
  • 307897 12/18/09 09:04 W. Deconinck Moller measurement postponed
  • 307896 12/18/09 08:58 W. Deconinck Moller measurement procedure started
  • 307895 12/18/09 08:57 apar Start_Parity_Run_14750__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=reset ROC 23,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307890 12/18/09 08:39 apar End_of_Parity_Run_14748__CODA_Config=Pvdis
  • 307887 12/18/09 08:28 apar Start_Parity_Run_14748__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307881 12/18/09 07:59 ZAFAR AHMED Owl shift summary
  • 307880 12/18/09 07:31 Al Tobias Target Screens during Production
  • 307879 12/18/09 07:27 apar Start_Parity_Run_14747__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307874 12/18/09 07:24 apar End_of_Parity_Run_14746__CODA_Config=Pvdis
  • 307873 12/18/09 06:50 ZAFAR AHMED Low N2 level in left arm
  • 307872 12/18/09 06:45 ZAFAR AHMED Hall A tools screen.
  • 307871 12/18/09 06:34 Al Tobias no global electron rate on Al dummy
  • 307870 12/18/09 06:29 apar Start_Parity_Run_14746__CODA_Config=Pvdis_Run_type=Production,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307865 12/18/09 06:27 apar End_of_Parity_Run_14745__CODA_Config=Pvdis
  • 307864 12/18/09 06:19 Apar Rates High? L/R T1 Rates = 4/6x10^5 at 100uA, PS1 prescales 230/360
  • 307863 12/18/09 05:49 auto_snap Left Arm HV snapshot
  • 307862 12/18/09 05:49 auto_snap Happex Charge Asym snapshot
  • 307861 12/18/09 05:49 auto_snap Hall A Tools snapshot
  • 307860 12/18/09 05:49 auto_snap Compton Polarimeter snapshot
  • 307859 12/18/09 05:49 auto_snap Beam Position and Current snapshot
  • 307858 12/18/09 05:49 auto_snap Accelerator Energies snapshot
  • 307857 12/18/09 05:45 A. Camsonne Vertical scan started for Compton
  • 307853 12/18/09 05:21 Al Tobias Rates at 100uA w/ 4x4mm raster on LD2
  • 307851 12/18/09 05:12 Al Tobias Rates again at 2uA 4x4mm on LD2
  • 307850 12/18/09 05:12 J. Rubin Spot++ run 26338 (L)
  • 307849 12/18/09 05:08 Al Tobias Move to LD2 Target to start Production
  • 307847 12/18/09 04:57 apar End_of_Parity_Run_14744__CODA_Config=Pvdis
  • 307844 12/18/09 04:56 J. Rubin Wrong plot -- see log #307840
  • 307843 12/18/09 04:55 J. Rubin Wrong plot -- see log #307840
  • 307842 12/18/09 04:55 J. Rubin Ignore please. I need HALOG training.
  • 307841 12/18/09 04:49 J. Rubin Wrong plot -- see log #307840
  • 307840 12/18/09 04:43 J. Rubin y vs. phi plots for runs 5839 (R) and 5838 (R)
  • 307837 12/18/09 04:14 apar Start_Parity_Run_14744__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=4.87,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=4868,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=in,rightarm_p=3100,rightarm_theta=12.9,target_fanspeed=60 Hz,target_type=20cm Al Dummy
  • 307836 12/18/09 04:04 J. Rubin Run 5840 (R) Plots -- y vs phi
  • 307835 12/18/09 03:58 J. Rubin Run 26336 (L) Plots -- y vs. phi and sieve
  • 307834 12/18/09 03:57 Kai see log #307831
  • 307833 12/18/09 03:55 J. Rubin Run 5839 (R) Plots -- y vs phi
  • 307832 12/18/09 03:55 J. Rubin Error in prior post. Right sieve *was not* in. (Left was)
  • 307831 12/18/09 03:53 Kai update new kinematics parameter and golden rootfile
  • 307830 12/18/09 03:51 J. Rubin Run 26335 (L) Plots -- y vs. phi and sieve
  • 307829 12/18/09 03:47 J. Rubin Run 5838 (R) Plots -- y vs phi
  • 307828 12/18/09 03:47 Al Tobias LHRS Sieve slits are OUT
  • 307827 12/18/09 03:44 Al Tobias Moved to Aluminum Dummy Target
  • 307826 12/18/09 03:42 J. Rubin Run 26334 (L) Plots -- y vs. phi and sieve
  • 307823 12/18/09 03:31 Al Tobias Rates w/ 5uA 4x4 raster on LD2
  • 307822 12/18/09 03:24 Al Tobias Moved to LD2 Target
  • 307817 12/18/09 02:59 R. Michaels different LG gains
  • 307816 12/18/09 02:58 Al Tobias Target Screens
  • 307815 12/18/09 02:36 ZAFAR AHMED Right arm optics run
  • 307812 12/18/09 02:16 R. Subedi Sieve pattern
  • 307809 12/18/09 02:01 R. Subedi y_tg ves ph_tg (Raster ON/OFF)
  • 307808 12/18/09 01:49 auto_snap Left Arm HV snapshot
  • 307807 12/18/09 01:49 auto_snap Happex Charge Asym snapshot
  • 307806 12/18/09 01:49 auto_snap Hall A Tools snapshot
  • 307805 12/18/09 01:49 auto_snap Compton Polarimeter snapshot
  • 307804 12/18/09 01:49 auto_snap Beam Position and Current snapshot
  • 307803 12/18/09 01:49 auto_snap Accelerator Energies snapshot
  • 307802 12/18/09 01:38 R. Subedi Sieve IN
  • 307797 12/18/09 01:27 Al Tobias Rates w/ 2uA 4x4mm raster on OPTICS
  • 307792 12/18/09 01:15 Al Tobias Rates w/ 2uA no raster on OPTICS target
  • 307789 12/18/09 01:09 Al Tobias We are at Optics Target
  • 307786 12/18/09 01:05 Al Tobias We are on Aluminum Dummy Target
  • 307783 12/18/09 00:52 Kai PID on the right arm
  • 307782 12/18/09 00:47 Kai PID performance on left arm
  • 307781 12/18/09 00:44 J. Rubin Q^2 Analysis at New Kinematics
  • 307780 12/18/09 00:42 R. Subedi Long recovery procedure finished
  • 307773 12/18/09 00:27 Al Tobias spot++ w/ 2uA 4x4mm raster on LD2
  • 307772 12/18/09 00:26 Al Tobias T1 rates for 2uA 4x4mm raster on LD2
  • 307771 12/18/09 00:23 Al Tobias spot++ w/ 2uA 2.5x2.5mm raster on LD2
  • 307764 12/18/09 00:10 LeRose Shift Summary
  • 307762 12/18/09 00:07 Al Tobias T1 rates at 2uA and 2.5x2.5mm raster on LD2
  • Thursday

  • 307761 12/17/09 23:52 jhlee spot++
  • 307758 12/17/09 23:39 apar End_of_Parity_Run_14743__CODA_Config=Pvdis
  • 307756 12/17/09 23:27 LeRose MCC working on high beam current
  • 307755 12/17/09 23:09 apar Start_Parity_Run_14743__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307750 12/17/09 22:59 apar End_of_Parity_Run_14742__CODA_Config=Pvdis
  • 307749 12/17/09 22:09 jhlee CODA crashed, restarted them
  • 307748 12/17/09 22:04 apar Start_Parity_Run_14742__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307741 12/17/09 21:55 apar Start_Parity_Run_14741__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307736 12/17/09 21:53 apar End_of_Parity_Run_14740__CODA_Config=Pvdis
  • 307735 12/17/09 21:49 auto_snap Left Arm HV snapshot
  • 307734 12/17/09 21:49 auto_snap Happex Charge Asym snapshot
  • 307733 12/17/09 21:48 auto_snap Hall A Tools snapshot
  • 307732 12/17/09 21:48 auto_snap Compton Polarimeter snapshot
  • 307731 12/17/09 21:48 auto_snap Beam Position and Current snapshot
  • 307730 12/17/09 21:48 auto_snap Accelerator Energies snapshot
  • 307729 12/17/09 21:37 A. Camsonne Compton rates during pathlength adjustement
  • 307728 12/17/09 21:17 R. Subedi Beam Status
  • 307727 12/17/09 20:57 apar Start_Parity_Run_14740__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307722 12/17/09 20:55 apar End_of_Parity_Run_14739__CODA_Config=Pvdis
  • 307721 12/17/09 20:00 apar Start_Parity_Run_14739__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307716 12/17/09 19:57 apar End_of_Parity_Run_14738__CODA_Config=Pvdis
  • 307715 12/17/09 19:01 apar Start_Parity_Run_14738__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307710 12/17/09 18:59 apar End_of_Parity_Run_14737__CODA_Config=Pvdis
  • 307709 12/17/09 18:03 apar Start_Parity_Run_14737__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307704 12/17/09 18:01 apar End_of_Parity_Run_14736__CODA_Config=Pvdis
  • 307703 12/17/09 18:00 aniol target screens
  • 307702 12/17/09 17:48 auto_snap Left Arm HV snapshot
  • 307701 12/17/09 17:48 auto_snap Happex Charge Asym snapshot
  • 307700 12/17/09 17:48 auto_snap Hall A Tools snapshot
  • 307699 12/17/09 17:48 auto_snap Compton Polarimeter snapshot
  • 307698 12/17/09 17:48 auto_snap Beam Position and Current snapshot
  • 307697 12/17/09 17:48 auto_snap Accelerator Energies snapshot
  • 307696 12/17/09 17:23 LeRose Left Arm Collimator Open
  • 307695 12/17/09 17:05 apar Start_Parity_Run_14736__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307690 12/17/09 17:03 apar End_of_Parity_Run_14735__CODA_Config=Pvdis
  • 307689 12/17/09 16:41 Jack (Techs) Shield House doors
  • 307688 12/17/09 16:19 LeRose Sieve in On Left Arm
  • 307687 12/17/09 16:07 apar Start_Parity_Run_14735__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307681 12/17/09 16:04 apar End_of_Parity_Run_14734__CODA_Config=Pvdis
  • 307680 12/17/09 16:04 W. Deconinck Day shift summary
  • 307675 12/17/09 15:18 D.Wang/X.Den Access work
  • 307674 12/17/09 15:11 W. Deconinck Tools screen capture
  • 307673 12/17/09 15:09 mcanan Target Screens
  • 307672 12/17/09 15:09 apar Start_Parity_Run_14734__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307671 12/17/09 15:07 apar End_of_Parity_Run_14733__CODA_Config=Pvdis
  • 307670 12/17/09 14:57 mcanan Cryo Pressure
  • 307669 12/17/09 14:52 W. Deconinck Ramped up L Q2 to 3.66 GeV/c as per new run plan
  • 307668 12/17/09 14:21 W. Deconinck R HRS detector hut door being worked on
  • 307667 12/17/09 14:12 apar Start_Parity_Run_14733__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307664 12/17/09 14:10 apar End_of_Parity_Run_14732__CODA_Config=Pvdis
  • 307663 12/17/09 13:48 auto_snap Left Arm HV snapshot
  • 307662 12/17/09 13:48 auto_snap Happex Charge Asym snapshot
  • 307661 12/17/09 13:48 auto_snap Hall A Tools snapshot
  • 307660 12/17/09 13:48 auto_snap Compton Polarimeter snapshot
  • 307659 12/17/09 13:48 auto_snap Beam Position and Current snapshot
  • 307658 12/17/09 13:48 auto_snap Accelerator Energies snapshot
  • 307657 12/17/09 13:33 W. Deconinck Sweeping crew can't close R HRS door (techs)
  • 307656 12/17/09 13:29 W. Deconinck Spectrometer magnet ramp up (Q2L trip overtemp)
  • 307655 12/17/09 13:15 apar Start_Parity_Run_14732__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307654 12/17/09 13:13 apar End_of_Parity_Run_14731__CODA_Config=Pvdis
  • 307653 12/17/09 12:35 pking Change in Injector Qweak ADC diagnostic signals
  • 307652 12/17/09 12:33 Jack (Techs) Left Q1 control board reset
  • 307651 12/17/09 12:25 W. Deconinck Sweep in progress
  • 307650 12/17/09 12:18 apar Start_Parity_Run_14731__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307649 12/17/09 12:16 apar End_of_Parity_Run_14730__CODA_Config=Pvdis
  • 307648 12/17/09 12:08 R. Michaels re: network
  • 307647 12/17/09 12:05 apar Start_Parity_Run_14730__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307646 12/17/09 12:03 apar End_of_Parity_Run_14729__CODA_Config=Pvdis
  • 307645 12/17/09 11:56 fansler/ewin Checklist
  • 307644 12/17/09 11:51 W. Deconinck No network to outside world in counting house
  • 307643 12/17/09 11:47 W. Deconinck PVDIS CODA restarted
  • 307642 12/17/09 11:47 apar Start_Parity_Run_14729__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307641 12/17/09 11:43 W. Deconinck Power cycled hatsv9
  • 307640 12/17/09 11:15 W. Deconinck Spectrometers L and R moved to 12.9 degree
  • 307639 12/17/09 10:42 mcanan haD2_P_Fill_Target_R pressure fluctuates
  • 307638 12/17/09 10:14 mcanan haD2_P_Fill_Target_R pressure decreases again
  • 307637 12/17/09 09:49 mcanan haD2_P_Fill_Target_R pressure restoring
  • 307636 12/17/09 09:47 auto_snap Left Arm HV snapshot
  • 307635 12/17/09 09:47 auto_snap Happex Charge Asym snapshot
  • 307634 12/17/09 09:47 auto_snap Hall A Tools snapshot
  • 307633 12/17/09 09:47 auto_snap Compton Polarimeter snapshot
  • 307632 12/17/09 09:47 auto_snap Beam Position and Current snapshot
  • 307631 12/17/09 09:47 auto_snap Accelerator Energies snapshot
  • 307630 12/17/09 09:22 apar Start_Parity_Run_14728__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307629 12/17/09 09:20 apar End_of_Parity_Run_14727__CODA_Config=Pvdis
  • 307628 12/17/09 09:04 mcanan haD2_P_Fill_Target_R pressure
  • 307627 12/17/09 08:24 apar Start_Parity_Run_14727__CODA_Config=Pvdis_Run_type=Pedestal,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307626 12/17/09 08:22 apar End_of_Parity_Run_14726__CODA_Config=Pvdis
  • 307625 12/17/09 08:12 R. Michaels disk cleanup
  • 307624 12/17/09 08:11 Al Tobias Owl Shift Summary
  • 307622 12/17/09 07:27 apar Start_Parity_Run_14726__CODA_Config=Pvdis_Run_type=Pedestal,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307621 12/17/09 07:10 kamyers target moved to empty
  • 307618 12/17/09 07:02 apar End_of_Parity_Run_14725__CODA_Config=Pvdis
  • 307617 12/17/09 07:02 Kai ask for beam off, ready for access
  • 307616 12/17/09 06:51 apar Start_Parity_Run_14725__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307611 12/17/09 06:50 apar End_of_Parity_Run_14724__CODA_Config=Pvdis
  • 307610 12/17/09 06:41 Al Tobias Compton DAQ: ROC6 not connecting on download
  • 307609 12/17/09 05:54 apar Start_Parity_Run_14724__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307605 12/17/09 05:52 apar End_of_Parity_Run_14723__CODA_Config=Pvdis
  • 307604 12/17/09 05:52 Al Tobias Hall A Tools Snapshot
  • 307603 12/17/09 05:47 auto_snap Left Arm HV snapshot
  • 307602 12/17/09 05:47 auto_snap Happex Charge Asym snapshot
  • 307601 12/17/09 05:47 auto_snap Hall A Tools snapshot
  • 307600 12/17/09 05:47 auto_snap Compton Polarimeter snapshot
  • 307599 12/17/09 05:47 auto_snap Beam Position and Current snapshot
  • 307598 12/17/09 05:47 auto_snap Accelerator Energies snapshot
  • 307597 12/17/09 04:59 Al Tobias Compton DAQ: coda reboot
  • 307596 12/17/09 04:57 apar Start_Parity_Run_14723__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307591 12/17/09 04:55 apar End_of_Parity_Run_14722__CODA_Config=Pvdis
  • 307590 12/17/09 04:39 Kai xterms of TS1, ROC23 were closed now, but still cannot do "telnet" connection
  • 307589 12/17/09 04:31 Kai TS1, ROC23 hang up, and cannot do reboot. Data is fine.
  • 307587 12/17/09 04:00 apar Start_Parity_Run_14722__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307585 12/17/09 03:56 kamyers target screens
  • 307582 12/17/09 03:49 apar End_of_Parity_Run_14721__CODA_Config=Pvdis
  • 307581 12/17/09 02:54 apar Start_Parity_Run_14721__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307578 12/17/09 02:53 apar End_of_Parity_Run_14720__CODA_Config=Pvdis
  • 307575 12/17/09 02:08 J. Rubin spot++
  • 307574 12/17/09 02:02 Al Tobias Beam is back - Hall A Tools snapshot
  • 307573 12/17/09 01:56 apar Start_Parity_Run_14720__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307568 12/17/09 01:51 apar End_of_Parity_Run_14719__CODA_Config=Pvdis
  • 307567 12/17/09 01:47 auto_snap Left Arm HV snapshot
  • 307566 12/17/09 01:47 auto_snap Happex Charge Asym snapshot
  • 307565 12/17/09 01:47 auto_snap Hall A Tools snapshot
  • 307564 12/17/09 01:47 auto_snap Compton Polarimeter snapshot
  • 307563 12/17/09 01:46 auto_snap Beam Position and Current snapshot
  • 307562 12/17/09 01:46 auto_snap Accelerator Energies snapshot
  • 307561 12/17/09 01:45 Al Tobias Beam OFF for awhile
  • 307560 12/17/09 01:17 apar Start_Parity_Run_14719__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307555 12/17/09 01:15 apar End_of_Parity_Run_14718__CODA_Config=Pvdis
  • 307554 12/17/09 00:19 apar Start_Parity_Run_14718__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307549 12/17/09 00:17 apar End_of_Parity_Run_14717__CODA_Config=Pvdis
  • 307547 12/17/09 00:00 ellie Swing Shift Summary
  • Wednesday

  • 307546 12/16/09 23:21 apar Start_Parity_Run_14717__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307541 12/16/09 23:19 apar End_of_Parity_Run_14716__CODA_Config=Pvdis
  • 307540 12/16/09 23:15 ellie Compton ROC5 and ROC6 stopped responding. Restart CODA, start new run #21635
  • 307539 12/16/09 22:24 apar Start_Parity_Run_14716__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307534 12/16/09 22:22 apar End_of_Parity_Run_14715__CODA_Config=Pvdis
  • 307533 12/16/09 21:46 auto_snap Left Arm HV snapshot
  • 307532 12/16/09 21:46 auto_snap Happex Charge Asym snapshot
  • 307531 12/16/09 21:46 auto_snap Hall A Tools snapshot
  • 307530 12/16/09 21:46 auto_snap Compton Polarimeter snapshot
  • 307529 12/16/09 21:46 auto_snap Beam Position and Current snapshot
  • 307528 12/16/09 21:46 auto_snap Accelerator Energies snapshot
  • 307527 12/16/09 21:42 M Friend Compton orbit lock
  • 307526 12/16/09 21:39 ellie Hall A Tools Screen Grab
  • 307525 12/16/09 21:37 apar Start_Parity_Run_14715__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307524 12/16/09 21:33 M Friend Compton cavity tune
  • 307523 12/16/09 21:33 apar Start_Parity_Run_14714__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307520 12/16/09 21:27 ZAFAR AHMED Parity CODA crashed and restarted
  • 307519 12/16/09 21:22 A. Camsonne iocha22 rebooted again
  • 307516 12/16/09 21:19 apar End_of_Parity_Run_14713__CODA_Config=Pvdis
  • 307515 12/16/09 20:52 ZAFAR AHMED IOCHA22 Rebooted
  • 307514 12/16/09 20:50 apar Start_Parity_Run_14713__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307511 12/16/09 20:43 ellie Hall A IOC Systems screen grab with white problem
  • 307508 12/16/09 20:38 apar End_of_Parity_Run_14712__CODA_Config=Pvdis
  • 307507 12/16/09 20:38 ellie LeCroy HV -- everything turned white!
  • 307506 12/16/09 20:36 aniol target screens
  • 307505 12/16/09 19:59 adaq Compton ROC0 in disconnected state. Restarting Compton CODA...
  • 307504 12/16/09 19:52 apar Start_Parity_Run_14712__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307501 12/16/09 19:40 brads (late entry) Moller VME crate plugged into hareboot26, socket 3
  • 307498 12/16/09 19:38 apar End_of_Parity_Run_14711__CODA_Config=Pvdis
  • 307497 12/16/09 18:53 apar Start_Parity_Run_14711__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307494 12/16/09 18:50 ellie Run 26300 should read PRODUCTION
  • 307493 12/16/09 18:48 ellie Run 5805 should read PRODUCTION
  • 307490 12/16/09 18:39 apar End_of_Parity_Run_14710__CODA_Config=Pvdis
  • 307489 12/16/09 18:38 A. Camsonne Compton ROC6 disconnected state - restarting CODA
  • 307488 12/16/09 17:56 ellie 110uA beam trippy, so it is now set to 105uA
  • 307487 12/16/09 17:54 apar Start_Parity_Run_14710__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307484 12/16/09 17:49 ellie End Compton run. Failed. Rebooted. Started new Compton Run # 21630
  • 307483 12/16/09 17:46 auto_snap Left Arm HV snapshot
  • 307482 12/16/09 17:46 auto_snap Happex Charge Asym snapshot
  • 307481 12/16/09 17:46 auto_snap Hall A Tools snapshot
  • 307480 12/16/09 17:46 auto_snap Compton Polarimeter snapshot
  • 307479 12/16/09 17:46 auto_snap Beam Position and Current snapshot
  • 307478 12/16/09 17:46 auto_snap Accelerator Energies snapshot
  • 307477 12/16/09 17:43 ellie Beginning Production Data Taking
  • 307476 12/16/09 17:42 ellie Moller finished, start Beam Recovery Procedure
  • 307475 12/16/09 17:32 ZAFAR AHMED Q^2 Measurement
  • 307474 12/16/09 17:32 aniol target screens
  • 307469 12/16/09 17:21 ellie Spot++ Check on LD2 target, Run # 26298
  • 307466 12/16/09 16:46 Kai turn the VDC on, and ready to recover the beam
  • 307465 12/16/09 16:37 pking Run coordinator daily summary
  • 307464 12/16/09 16:36 O. Glamazdin Moller results
  • 307463 12/16/09 16:29 ellie LeCroy HV Trip on Card 11, Channel 1
  • 307461 12/16/09 16:00 W. Deconinck Day shift summary
  • 307460 12/16/09 15:25 W. Deconinck Tools screen capture (with pic)
  • 307459 12/16/09 15:25 W. Deconinck Tools screen capture
  • 307458 12/16/09 14:58 mcanan Target Screens
  • 307457 12/16/09 14:47 R. Michaels blowing top out of Coulomb meter !
  • 307456 12/16/09 13:46 auto_snap Left Arm HV snapshot
  • 307455 12/16/09 13:45 auto_snap Happex Charge Asym snapshot
  • 307454 12/16/09 13:45 auto_snap Hall A Tools snapshot
  • 307453 12/16/09 13:45 auto_snap Compton Polarimeter snapshot
  • 307452 12/16/09 13:45 auto_snap Beam Position and Current snapshot
  • 307451 12/16/09 13:45 auto_snap Accelerator Energies snapshot
  • 307450 12/16/09 13:35 W. Deconinck Moller tuning still ongoing
  • 307449 12/16/09 12:07 W. Deconinck Back to beam permit, but still problems in accelerator
  • 307448 12/16/09 11:41 W. Deconinck Another controlled access requested for new Moller crate reboot
  • 307447 12/16/09 11:33 mcanan Target Motion
  • 307446 12/16/09 11:30 mcanan Scattering Chamber Vacuum Alarm
  • 307445 12/16/09 11:25 W. Deconinck Going back to beam permit
  • 307444 12/16/09 11:06 mcanan Scattering Chamber Vacuum Alarm
  • 307443 12/16/09 10:55 W. Deconinck Controlled access w/ escort requested for Moller crate reboot
  • 307442 12/16/09 10:40 W. Deconinck Starting moller measurement
  • 307441 12/16/09 10:39 apar End_of_Parity_Run_14709__CODA_Config=Pvdis
  • 307438 12/16/09 10:19 W. Deconinck Compton CODA reboot
  • 307435 12/16/09 10:00 apar Start_Parity_Run_14709__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307431 12/16/09 09:48 apar End_of_Parity_Run_14708__CODA_Config=Pvdis
  • 307430 12/16/09 09:45 auto_snap Left Arm HV snapshot
  • 307429 12/16/09 09:45 auto_snap Happex Charge Asym snapshot
  • 307428 12/16/09 09:45 auto_snap Hall A Tools snapshot
  • 307427 12/16/09 09:45 auto_snap Compton Polarimeter snapshot
  • 307426 12/16/09 09:45 auto_snap Beam Position and Current snapshot
  • 307425 12/16/09 09:45 auto_snap Accelerator Energies snapshot
  • 307424 12/16/09 09:02 apar Start_Parity_Run_14708__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307419 12/16/09 08:50 apar End_of_Parity_Run_14707__CODA_Config=Pvdis
  • 307418 12/16/09 08:30 mcanan scattering chamber vacuum
  • 307417 12/16/09 08:27 W. Deconinck Compton HV trip
  • 307413 12/16/09 07:58 flay 12/16/09 Owl Shift Summary
  • 307412 12/16/09 07:57 apar Start_Parity_Run_14707__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307409 12/16/09 07:52 apar End_of_Parity_Run_14706__CODA_Config=Pvdis
  • 307408 12/16/09 07:04 apar Start_Parity_Run_14706__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307405 12/16/09 06:54 J. Rubin Spot++ after beam recovery
  • 307400 12/16/09 06:36 flay Beam is back
  • 307399 12/16/09 06:30 apar End_of_Parity_Run_14705__CODA_Config=Pvdis
  • 307397 12/16/09 06:10 apar Start_Parity_Run_14705__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307392 12/16/09 05:57 apar End_of_Parity_Run_14704__CODA_Config=Pvdis
  • 307391 12/16/09 05:54 kamyers target screens
  • 307390 12/16/09 05:45 auto_snap Left Arm HV snapshot
  • 307389 12/16/09 05:45 auto_snap Happex Charge Asym snapshot
  • 307388 12/16/09 05:45 auto_snap Hall A Tools snapshot
  • 307387 12/16/09 05:45 auto_snap Compton Polarimeter snapshot
  • 307386 12/16/09 05:45 auto_snap Beam Position and Current snapshot
  • 307385 12/16/09 05:44 auto_snap Accelerator Energies snapshot
  • 307382 12/16/09 05:02 apar Start_Parity_Run_14704__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307381 12/16/09 03:53 apar End_of_Parity_Run_14703__CODA_Config=Pvdis
  • 307377 12/16/09 02:59 apar Start_Parity_Run_14703__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307374 12/16/09 02:55 apar End_of_Parity_Run_14702__CODA_Config=Pvdis
  • 307373 12/16/09 02:47 flay Beam Down; MCC taking a look at Cryo Modules
  • 307372 12/16/09 02:14 D.Wang Dithering working again
  • 307371 12/16/09 02:09 apar Start_Parity_Run_14702__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307366 12/16/09 01:55 apar End_of_Parity_Run_14701__CODA_Config=Pvdis
  • 307365 12/16/09 01:44 auto_snap Left Arm HV snapshot
  • 307364 12/16/09 01:44 auto_snap Happex Charge Asym snapshot
  • 307363 12/16/09 01:44 auto_snap Hall A Tools snapshot
  • 307362 12/16/09 01:44 auto_snap Compton Polarimeter snapshot
  • 307361 12/16/09 01:44 auto_snap Beam Position and Current snapshot
  • 307360 12/16/09 01:44 auto_snap Accelerator Energies snapshot
  • 307359 12/16/09 01:33 kamyers Scattering chamber vacuum oscillating again
  • 307358 12/16/09 01:32 flay Hall A Tools
  • 307357 12/16/09 01:31 flay Dithering Issue to Be Solved Later
  • 307356 12/16/09 01:10 apar Start_Parity_Run_14701__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307355 12/16/09 00:54 apar End_of_Parity_Run_14700__CODA_Config=Pvdis
  • 307352 12/16/09 00:42 apar Start_Parity_Run_14700__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307349 12/16/09 00:30 apar End_of_Parity_Run_14699__CODA_Config=Pvdis
  • 307348 12/16/09 00:24 apar Start_Parity_Run_14699__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307345 12/16/09 00:19 D.Wang Dithering not working after epics reboot
  • 307342 12/16/09 00:11 apar End_of_Parity_Run_14698__CODA_Config=Pvdis
  • 307341 12/16/09 00:09 D.Wang epics reboot
  • Tuesday

  • 307339 12/15/09 23:59 A. Deur Swing shift summary
  • 307338 12/15/09 23:58 aniol target screens
  • 307337 12/15/09 23:26 apar Start_Parity_Run_14698__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307332 12/15/09 23:13 apar End_of_Parity_Run_14697__CODA_Config=Pvdis
  • 307331 12/15/09 22:32 Compton posi A. Deur
  • 307330 12/15/09 22:28 apar Start_Parity_Run_14697__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307325 12/15/09 22:15 apar End_of_Parity_Run_14696__CODA_Config=Pvdis
  • 307324 12/15/09 22:14 aniol target vacuum and beamtrips
  • 307323 12/15/09 22:07 aniol target vacuum very good
  • 307322 12/15/09 21:44 auto_snap Left Arm HV snapshot
  • 307321 12/15/09 21:44 auto_snap Happex Charge Asym snapshot
  • 307320 12/15/09 21:44 auto_snap Hall A Tools snapshot
  • 307319 12/15/09 21:44 auto_snap Compton Polarimeter snapshot
  • 307318 12/15/09 21:44 auto_snap Beam Position and Current snapshot
  • 307317 12/15/09 21:44 auto_snap Accelerator Energies snapshot
  • 307316 12/15/09 21:30 apar Start_Parity_Run_14696__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307311 12/15/09 21:17 apar End_of_Parity_Run_14695__CODA_Config=Pvdis
  • 307310 12/15/09 20:32 apar Start_Parity_Run_14695__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307305 12/15/09 20:19 apar End_of_Parity_Run_14694__CODA_Config=Pvdis
  • 307304 12/15/09 20:04 jhlee a bit more look
  • 307303 12/15/09 20:00 jhlee run again through run 14683 - 14688
  • 307302 12/15/09 19:34 apar Start_Parity_Run_14694__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307299 12/15/09 19:27 pking Run coordinator daily summary
  • 307298 12/15/09 19:24 pking ROOT files missing for Parity runs 14683-14688
  • 307295 12/15/09 19:21 apar End_of_Parity_Run_14693__CODA_Config=Pvdis
  • 307294 12/15/09 18:55 M Friend Compton orbit lock change
  • 307293 12/15/09 18:36 apar Start_Parity_Run_14693__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307288 12/15/09 18:23 apar End_of_Parity_Run_14692__CODA_Config=Pvdis
  • 307287 12/15/09 18:16 R. Michaels re: rsync
  • 307286 12/15/09 17:44 auto_snap Left Arm HV snapshot
  • 307285 12/15/09 17:44 auto_snap Happex Charge Asym snapshot
  • 307284 12/15/09 17:44 auto_snap Hall A Tools snapshot
  • 307283 12/15/09 17:43 auto_snap Compton Polarimeter snapshot
  • 307282 12/15/09 17:43 auto_snap Beam Position and Current snapshot
  • 307281 12/15/09 17:43 auto_snap Accelerator Energies snapshot
  • 307280 12/15/09 17:41 jhlee This run was ended by the user, because EB1 and ER1 didn't react
  • 307277 12/15/09 17:38 apar Start_Parity_Run_14692__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307276 12/15/09 17:35 apar Start_Parity_Run_14691__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307275 12/15/09 17:24 apar End_of_Parity_Run_14691__CODA_Config=Pvdis
  • 307274 12/15/09 17:22 brads re: question about the backup processes
  • 307271 12/15/09 17:11 aniol target - vac. and beam trips
  • 307270 12/15/09 17:08 jhlee question about the backup processes
  • 307269 12/15/09 16:35 A. Deur bpm/bcm strip chart
  • 307268 12/15/09 16:32 apar Start_Parity_Run_14690__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307263 12/15/09 16:18 apar End_of_Parity_Run_14689__CODA_Config=Pvdis
  • 307262 12/15/09 16:18 apar End_of_Parity_Run_14689__CODA_Config=Pvdis
  • 307261 12/15/09 16:13 aniol target screens
  • 307260 12/15/09 16:13 A. Deur Hall A tools
  • 307259 12/15/09 16:11 A. Deur routine spot++
  • 307257 12/15/09 16:02 Kai ROC3 crash
  • 307255 12/15/09 16:00 W. Deconinck Day shift summary
  • 307254 12/15/09 15:33 apar Start_Parity_Run_14689__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307253 12/15/09 15:33 Holmstrom Target Strip Charts
  • 307250 12/15/09 15:30 W. Deconinck Compton HV trip
  • 307249 12/15/09 15:28 Holmstrom Target Screen
  • 307248 12/15/09 15:27 W. Deconinck Tools screen capture
  • 307247 12/15/09 15:21 W. Deconinck Back to production
  • 307246 12/15/09 15:13 W. Deconinck BSY power supply certification
  • 307245 12/15/09 15:13 apar End_of_Parity_Run_14688__CODA_Config=Pvdis
  • 307242 12/15/09 15:01 apar Start_Parity_Run_14688__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307239 12/15/09 14:58 W. Deconinck Compton HV trip
  • 307235 12/15/09 14:43 apar Start_Parity_Run_14687__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307231 12/15/09 14:33 apar End_of_Parity_Run_14686__CODA_Config=Pvdis
  • 307230 12/15/09 14:25 Holmstrom Scatt Chamber Vac.
  • 307229 12/15/09 14:03 Holmstrom Scattering Vacuum
  • 307228 12/15/09 13:48 apar Start_Parity_Run_14686__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307225 12/15/09 13:43 auto_snap Left Arm HV snapshot
  • 307224 12/15/09 13:43 auto_snap Happex Charge Asym snapshot
  • 307223 12/15/09 13:43 auto_snap Hall A Tools snapshot
  • 307222 12/15/09 13:43 auto_snap Compton Polarimeter snapshot
  • 307221 12/15/09 13:43 auto_snap Beam Position and Current snapshot
  • 307220 12/15/09 13:43 auto_snap Accelerator Energies snapshot
  • 307217 12/15/09 13:35 apar End_of_Parity_Run_14685__CODA_Config=Pvdis
  • 307216 12/15/09 12:48 R. Subedi PID efficiency good
  • 307213 12/15/09 12:40 apar Start_Parity_Run_14685__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307210 12/15/09 12:37 apar End_of_Parity_Run_14684__CODA_Config=Pvdis
  • 307209 12/15/09 12:23 W. Deconinck Compton HV trip
  • 307208 12/15/09 11:52 apar Start_Parity_Run_14684__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307203 12/15/09 11:39 apar End_of_Parity_Run_14683__CODA_Config=Pvdis
  • 307202 12/15/09 10:53 apar Start_Parity_Run_14683__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307199 12/15/09 10:52 Deng Q^2 and W plots
  • 307198 12/15/09 10:48 Holmstrom Scatter Chamber Vac
  • 307197 12/15/09 10:40 W. Deconinck Back to production mode at 105 uA
  • 307196 12/15/09 10:30 W. Deconinck Ramping down VDCs after Q2 measurement
  • 307188 12/15/09 10:13 W. Deconinck Back to production running
  • 307187 12/15/09 10:13 Holmstrom Fan Speed
  • 307186 12/15/09 10:05 Holmstrom Fan Speed
  • 307185 12/15/09 10:01 W. Deconinck Changed from IHWP OUT to IN
  • 307184 12/15/09 10:00 apar Halfwave plate IN, Slug 24
  • 307183 12/15/09 09:45 jpchen target temperature readback
  • 307182 12/15/09 09:43 auto_snap Left Arm HV snapshot
  • 307181 12/15/09 09:43 auto_snap Happex Charge Asym snapshot
  • 307180 12/15/09 09:43 auto_snap Hall A Tools snapshot
  • 307179 12/15/09 09:43 auto_snap Compton Polarimeter snapshot
  • 307178 12/15/09 09:43 auto_snap Beam Position and Current snapshot
  • 307177 12/15/09 09:43 auto_snap Accelerator Energies snapshot
  • 307176 12/15/09 09:41 W. Deconinck Cryo has restored our flows
  • 307175 12/15/09 09:27 W. Deconinck Cryo has to reduce the cryoflow to our target
  • 307174 12/15/09 09:10 Holmstrom Target IOC problem
  • 307173 12/15/09 09:10 W. Deconinck Controlled access for target expert
  • 307170 12/15/09 09:04 apar End_of_Parity_Run_14682__CODA_Config=Pvdis
  • 307169 12/15/09 08:57 W. Deconinck Compton ROC5 crashed
  • 307168 12/15/09 08:51 apar Start_Parity_Run_14682__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307163 12/15/09 08:39 apar End_of_Parity_Run_14681__CODA_Config=Pvdis
  • 307162 12/15/09 08:07 W. Deconinck Shift transfer
  • 307160 12/15/09 08:00 flay 12/15/09 Owl Shift Summary
  • 307159 12/15/09 07:53 apar Start_Parity_Run_14681__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307154 12/15/09 07:41 apar End_of_Parity_Run_14680__CODA_Config=Pvdis
  • 307153 12/15/09 06:56 apar Start_Parity_Run_14680__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307148 12/15/09 06:44 apar End_of_Parity_Run_14679__CODA_Config=Pvdis
  • 307145 12/15/09 05:51 apar Start_Parity_Run_14679__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307144 12/15/09 05:42 auto_snap Left Arm HV snapshot
  • 307143 12/15/09 05:42 auto_snap Happex Charge Asym snapshot
  • 307142 12/15/09 05:42 auto_snap Hall A Tools snapshot
  • 307141 12/15/09 05:42 auto_snap Compton Polarimeter snapshot
  • 307140 12/15/09 05:42 auto_snap Beam Position and Current snapshot
  • 307139 12/15/09 05:42 auto_snap Accelerator Energies snapshot
  • 307138 12/15/09 05:27 apar Start_Parity_Run_14678__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307131 12/15/09 05:15 apar End_of_Parity_Run_14677__CODA_Config=Pvdis
  • 307130 12/15/09 04:50 kamyers target temp not being displayed correctly on graph
  • 307129 12/15/09 04:30 apar Start_Parity_Run_14677__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307124 12/15/09 04:18 apar End_of_Parity_Run_14676__CODA_Config=Pvdis
  • 307123 12/15/09 04:06 kamyers HP heater dropped to zero for ~2 minutes
  • 307122 12/15/09 03:32 apar Start_Parity_Run_14676__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307117 12/15/09 03:20 apar End_of_Parity_Run_14675__CODA_Config=Pvdis
  • 307116 12/15/09 03:15 flay Beamline Issues
  • 307115 12/15/09 03:15 flay Beamline Issues
  • 307114 12/15/09 02:58 kamyers target screens
  • 307113 12/15/09 02:58 flay Hall A Tools
  • 307112 12/15/09 02:35 apar Start_Parity_Run_14675__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307107 12/15/09 02:23 apar End_of_Parity_Run_14674__CODA_Config=Pvdis
  • 307106 12/15/09 01:42 auto_snap Left Arm HV snapshot
  • 307105 12/15/09 01:42 auto_snap Happex Charge Asym snapshot
  • 307104 12/15/09 01:42 auto_snap Hall A Tools snapshot
  • 307103 12/15/09 01:42 auto_snap Compton Polarimeter snapshot
  • 307102 12/15/09 01:42 auto_snap Beam Position and Current snapshot
  • 307101 12/15/09 01:42 auto_snap Accelerator Energies snapshot
  • 307100 12/15/09 01:37 apar Start_Parity_Run_14674__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307095 12/15/09 01:25 apar End_of_Parity_Run_14673__CODA_Config=Pvdis
  • 307094 12/15/09 00:39 J. Rubin Raster
  • 307092 12/15/09 00:39 apar Start_Parity_Run_14673__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307091 12/15/09 00:38 flay Hall A Tools
  • 307087 12/15/09 00:28 apar End_of_Parity_Run_14672__CODA_Config=Pvdis
  • 307085 12/15/09 00:02 S.Sirca/X.De swing shift summary
  • Monday

  • 307084 12/14/09 23:56 Deng spot++ check
  • 307083 12/14/09 23:51 Deng hall A general tools screen shot
  • 307082 12/14/09 23:49 aniol target screens
  • 307081 12/14/09 23:42 apar Start_Parity_Run_14672__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307076 12/14/09 23:30 apar End_of_Parity_Run_14671__CODA_Config=Pvdis
  • 307075 12/14/09 23:29 Han restart DATAMON for L-HRS
  • 307073 12/14/09 23:16 D.Wang reset ROC3
  • 307070 12/14/09 22:35 apar Start_Parity_Run_14671__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307067 12/14/09 22:32 apar End_of_Parity_Run_14670__CODA_Config=Pvdis
  • 307066 12/14/09 22:15 pking Run coordinator daily summary
  • 307065 12/14/09 22:02 X. Zheng How to monitor efficiencies (for shift workers)
  • 307064 12/14/09 21:42 auto_snap Left Arm HV snapshot
  • 307063 12/14/09 21:42 auto_snap Happex Charge Asym snapshot
  • 307062 12/14/09 21:42 auto_snap Hall A Tools snapshot
  • 307061 12/14/09 21:42 auto_snap Compton Polarimeter snapshot
  • 307060 12/14/09 21:41 auto_snap Beam Position and Current snapshot
  • 307059 12/14/09 21:41 auto_snap Accelerator Energies snapshot
  • 307058 12/14/09 21:37 apar Start_Parity_Run_14670__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307053 12/14/09 21:33 apar End_of_Parity_Run_14669__CODA_Config=Pvdis
  • 307052 12/14/09 21:03 D.Wang compton rate lower with beam position -1.05
  • 307049 12/14/09 20:39 S. Sirca HOWTO fix HALOG sequential number
  • 307048 12/14/09 20:38 Deng pvdisL electron rate group 5 and 6 look good now for production
  • 307047 12/14/09 20:37 apar Start_Parity_Run_14669__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307044 12/14/09 20:35 D.Wang Left arm group 5 electron wide efficiency problem seems to be fixed
  • 307043 12/14/09 20:35 Ramesh/Dianc Access Activity
  • 307042 12/14/09 20:34 S. Sirca Test of HALOG number fix
  • 19 12/14/09 20:08 apar End_of_Parity_Run_14668__CODA_Config=Pvdis
  • 17 12/14/09 19:32 aniol target screens
  • 16 12/14/09 19:21 S. Sirca Request controlled access to investigate HRSL inefficiency problem
  • 15 12/14/09 19:14 apar Start_Parity_Run_14668__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=tagger on,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 14 12/14/09 19:12 apar End_of_Parity_Run_14667__CODA_Config=Pvdis
  • 13 12/14/09 19:08 apar Start_Parity_Run_14667__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 9 12/14/09 19:06 apar End_of_Parity_Run_14666__CODA_Config=Pvdis
  • 8 12/14/09 18:45 S. Sirca Beam down to 100uA
  • 5 12/14/09 18:38 apar Start_Parity_Run_14666__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 2 12/14/09 18:34 apar Start_Parity_Run_14665__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307037 12/14/09 18:27 apar End_of_Parity_Run_14664__CODA_Config=Pvdis
  • 307036 12/14/09 17:58 M Friend Comtpon cavity tune
  • 307035 12/14/09 17:41 auto_snap Left Arm HV snapshot
  • 307034 12/14/09 17:41 auto_snap Happex Charge Asym snapshot
  • 307033 12/14/09 17:41 auto_snap Hall A Tools snapshot
  • 307032 12/14/09 17:41 auto_snap Compton Polarimeter snapshot
  • 307031 12/14/09 17:41 auto_snap Beam Position and Current snapshot
  • 307030 12/14/09 17:41 D.Wang reset LHRS trigger supervisor
  • 307029 12/14/09 17:41 auto_snap Accelerator Energies snapshot
  • 307027 12/14/09 17:32 apar Start_Parity_Run_14664__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307024 12/14/09 17:20 D.Wang tagger is on for parity run 14663 and LHRS run 26246
  • 307022 12/14/09 17:18 apar End_of_Parity_Run_14663__CODA_Config=Pvdis
  • 307021 12/14/09 17:17 apar Start_Parity_Run_14663__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=tagger on test,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 307019 12/14/09 17:14 X. Zheng Left gr5 electron efficiency degraded more
  • 307018 12/14/09 17:03 S. Sirca Spot++ for 2uA on LD2, 6x6 MCC raster
  • 307015 12/14/09 16:58 S. Sirca Raw rates with 2uA on LD2, 6x6 MCC raster
  • 307014 12/14/09 16:58 Deng boiling study run summary
  • 307011 12/14/09 16:52 Deng/Dwang Boiling test runs summary
  • 307008 12/14/09 16:41 apar End_of_Parity_Run_14662__CODA_Config=Pvdis
  • 307007 12/14/09 16:33 apar Start_Parity_Run_14662__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 60Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 307002 12/14/09 16:29 apar End_of_Parity_Run_14661__CODA_Config=Pvdis
  • 307001 12/14/09 16:06 apar Start_Parity_Run_14661__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 50Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306996 12/14/09 16:04 apar End_of_Parity_Run_14660__CODA_Config=Pvdis
  • 306994 12/14/09 16:01 W. Deconinck Day shift summary
  • 306993 12/14/09 15:58 apar Start_Parity_Run_14660__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 41Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306990 12/14/09 15:55 apar End_of_Parity_Run_14659__CODA_Config=Pvdis
  • 306987 12/14/09 15:52 W. Deconinck Tools screen capture
  • 306986 12/14/09 15:49 apar Start_Parity_Run_14659__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 41Hz, 4x4, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306981 12/14/09 15:47 apar End_of_Parity_Run_14658__CODA_Config=Pvdis
  • 306980 12/14/09 15:37 apar Start_Parity_Run_14658__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 50Hz, 4x4, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306975 12/14/09 15:35 apar End_of_Parity_Run_14657__CODA_Config=Pvdis
  • 306974 12/14/09 15:34 Sirish Compton
  • 306973 12/14/09 15:23 apar Start_Parity_Run_14657__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 60Hz, 4x4, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306970 12/14/09 15:22 Sirish Compton cavity
  • 306967 12/14/09 15:21 apar End_of_Parity_Run_14656__CODA_Config=Pvdis
  • 306962 12/14/09 15:05 X. Zheng Left arm gr5 EN and EW rates from parity runs
  • 306961 12/14/09 15:05 apar Start_Parity_Run_14656__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 60Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306956 12/14/09 15:03 apar End_of_Parity_Run_14655__CODA_Config=Pvdis
  • 306955 12/14/09 14:57 apar Start_Parity_Run_14655__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 50Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306950 12/14/09 14:53 apar End_of_Parity_Run_14654__CODA_Config=Pvdis
  • 306949 12/14/09 14:50 X. Zheng Timing difference spectrum
  • 306948 12/14/09 14:47 apar Start_Parity_Run_14654__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 115uA, 41Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306942 12/14/09 14:35 apar Start_Parity_Run_14653__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 41Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306941 12/14/09 14:33 apar Start_Parity_Run_14652__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 41Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306935 12/14/09 14:30 mcanan Target Screens
  • 306934 12/14/09 14:24 apar Start_Parity_Run_14651__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 51Hz, 6x6, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306929 12/14/09 14:21 apar End_of_Parity_Run_14650__CODA_Config=Pvdis
  • 306928 12/14/09 14:12 apar Start_Parity_Run_14650__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 51Hz, 4x4, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306923 12/14/09 14:10 apar End_of_Parity_Run_14649__CODA_Config=Pvdis
  • 306922 12/14/09 14:00 apar Start_Parity_Run_14649__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 41Hz, 4x4, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306917 12/14/09 13:56 apar End_of_Parity_Run_14648__CODA_Config=Pvdis
  • 306916 12/14/09 13:50 R. Subedi left hrs hole in wide-efficiency
  • 306915 12/14/09 13:47 apar Start_Parity_Run_14648__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 41Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306910 12/14/09 13:45 apar End_of_Parity_Run_14647__CODA_Config=Pvdis
  • 306909 12/14/09 13:41 auto_snap Left Arm HV snapshot
  • 306908 12/14/09 13:41 auto_snap Happex Charge Asym snapshot
  • 306907 12/14/09 13:41 auto_snap Hall A Tools snapshot
  • 306906 12/14/09 13:41 auto_snap Compton Polarimeter snapshot
  • 306905 12/14/09 13:41 auto_snap Beam Position and Current snapshot
  • 306904 12/14/09 13:41 auto_snap Accelerator Energies snapshot
  • 306903 12/14/09 13:39 apar Start_Parity_Run_14647__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 51Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306898 12/14/09 13:36 apar End_of_Parity_Run_14646__CODA_Config=Pvdis
  • 306897 12/14/09 13:31 apar Start_Parity_Run_14646__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 60Hz, 2x2, LD2 target,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306892 12/14/09 13:21 apar End_of_Parity_Run_14645__CODA_Config=Pvdis
  • 306891 12/14/09 13:15 apar Start_Parity_Run_14645__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 60Hz, 4X4,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306886 12/14/09 13:12 apar End_of_Parity_Run_14644__CODA_Config=Pvdis
  • 306885 12/14/09 13:10 Deng Q^2 measurement and plots
  • 306882 12/14/09 13:03 apar Start_Parity_Run_14644__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Boiling test, 100uA, 60Hz, 6x6,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306881 12/14/09 13:00 Deng spot++ check
  • 306880 12/14/09 12:59 W. Deconinck Back to 100 uA beam current
  • 306879 12/14/09 12:54 W. Deconinck Turning off the VDCs
  • 306876 12/14/09 12:34 Deng/D.wang BCM calibration summary
  • 306875 12/14/09 12:33 Apar spot++
  • 306872 12/14/09 12:27 mcanan Target motion
  • 306871 12/14/09 12:21 W. Deconinck Ramping up VDC HV
  • 306870 12/14/09 12:20 Deng HAPPEX feedback enabled
  • 306869 12/14/09 12:19 W. Deconinck Finished the BCM calibration
  • 306868 12/14/09 12:15 apar End_of_Parity_Run_14643__CODA_Config=Pvdis
  • 306865 12/14/09 12:09 W. Deconinck Target boiling tests: special requests
  • 306864 12/14/09 11:41 apar Start_Parity_Run_14643__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=bcm calibration,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306859 12/14/09 11:39 apar End_of_Parity_Run_14642__CODA_Config=Pvdis
  • 306858 12/14/09 10:54 apar Start_Parity_Run_14642__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=bcm calibration,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306853 12/14/09 10:51 apar End_of_Parity_Run_14641__CODA_Config=Pvdis
  • 306852 12/14/09 10:19 Deng unser is in pvdis scalers now
  • 306851 12/14/09 10:19 apar Start_Parity_Run_14641__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=bcm calibration,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306846 12/14/09 10:18 apar End_of_Parity_Run_14640__CODA_Config=Pvdis
  • 306845 12/14/09 10:15 Deng HAPPEX feedback disabled for bcm calibration
  • 306844 12/14/09 10:08 apar Start_Parity_Run_14640__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=bcm calibration,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306841 12/14/09 10:05 pking Run plan for today
  • 306840 12/14/09 09:54 mcanan Target Motion
  • 306839 12/14/09 09:53 apar End_of_Parity_Run_14639__CODA_Config=Pvdis
  • 306838 12/14/09 09:49 apar Start_Parity_Run_14639__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=test if pvdis scaler has unser signal,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306837 12/14/09 09:47 apar Start_Parity_Run_14638__CODA_Config=Pvdis_Run_type=Test,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=test if pvdis scaler has unser signal,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306836 12/14/09 09:47 W. Deconinck Starting current scan for BCM linearity test
  • 306835 12/14/09 09:46 D.Wang Unser signal plugged into V2F
  • 306832 12/14/09 09:42 apar End_of_Parity_Run_14637__CODA_Config=Pvdis
  • 306831 12/14/09 09:41 auto_snap Left Arm HV snapshot
  • 306830 12/14/09 09:40 auto_snap Happex Charge Asym snapshot
  • 306829 12/14/09 09:40 auto_snap Hall A Tools snapshot
  • 306828 12/14/09 09:40 auto_snap Compton Polarimeter snapshot
  • 306827 12/14/09 09:40 auto_snap Beam Position and Current snapshot
  • 306826 12/14/09 09:40 auto_snap Accelerator Energies snapshot
  • 306825 12/14/09 09:33 apar Start_Parity_Run_14637__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306820 12/14/09 09:32 apar End_of_Parity_Run_14636__CODA_Config=Pvdis
  • 306819 12/14/09 09:06 W. Deconinck Tools screen capture
  • 306818 12/14/09 08:51 W. Deconinck Beam current scan scheduled for 10 am
  • 306817 12/14/09 08:37 apar Start_Parity_Run_14636__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306812 12/14/09 08:35 apar End_of_Parity_Run_14635__CODA_Config=Pvdis
  • 306811 12/14/09 08:23 W. Deconinck Asked for 5 uA more current
  • 306810 12/14/09 08:08 R. Michaels R-HRS DAQ revived
  • 306807 12/14/09 08:00 ZAFAR AHMED Owl shift summary
  • 306806 12/14/09 07:39 apar Start_Parity_Run_14635__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306803 12/14/09 07:37 apar End_of_Parity_Run_14634__CODA_Config=Pvdis
  • 306802 12/14/09 06:41 apar Start_Parity_Run_14634__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306799 12/14/09 06:39 apar End_of_Parity_Run_14633__CODA_Config=Pvdis
  • 306798 12/14/09 05:53 kamyers target screens
  • 306796 12/14/09 05:43 apar Start_Parity_Run_14633__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306794 12/14/09 05:42 apar End_of_Parity_Run_14632__CODA_Config=Pvdis
  • 306793 12/14/09 05:40 auto_snap Left Arm HV snapshot
  • 306792 12/14/09 05:40 auto_snap Happex Charge Asym snapshot
  • 306791 12/14/09 05:40 auto_snap Hall A Tools snapshot
  • 306790 12/14/09 05:40 auto_snap Compton Polarimeter snapshot
  • 306789 12/14/09 05:40 auto_snap Beam Position and Current snapshot
  • 306788 12/14/09 05:40 auto_snap Accelerator Energies snapshot
  • 306787 12/14/09 04:54 ZAFAR AHMED ROC 2 is not rebooting
  • 306786 12/14/09 04:46 apar Start_Parity_Run_14632__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306783 12/14/09 04:21 apar End_of_Parity_Run_14631__CODA_Config=Pvdis
  • 306782 12/14/09 03:26 apar Start_Parity_Run_14631__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306777 12/14/09 03:24 apar End_of_Parity_Run_14630__CODA_Config=Pvdis
  • 306776 12/14/09 02:52 R. Michaels glitches noted in CH/HRS synch
  • 306775 12/14/09 02:28 apar Start_Parity_Run_14630__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306770 12/14/09 02:26 apar End_of_Parity_Run_14629__CODA_Config=Pvdis
  • 306769 12/14/09 01:40 auto_snap Left Arm HV snapshot
  • 306768 12/14/09 01:40 auto_snap Happex Charge Asym snapshot
  • 306767 12/14/09 01:40 auto_snap Hall A Tools snapshot
  • 306766 12/14/09 01:40 auto_snap Compton Polarimeter snapshot
  • 306765 12/14/09 01:40 auto_snap Beam Position and Current snapshot
  • 306764 12/14/09 01:39 auto_snap Accelerator Energies snapshot
  • 306763 12/14/09 01:31 apar Start_Parity_Run_14629__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306758 12/14/09 01:29 ZAFAR AHMED Compton Error
  • 306757 12/14/09 01:29 apar End_of_Parity_Run_14628__CODA_Config=Pvdis
  • 306756 12/14/09 01:11 J. Rubin Raster Size
  • 306755 12/14/09 00:43 J. Rubin PVDIS DAQ Run-Start Info Possibly Mislabeled
  • 306754 12/14/09 00:34 apar Start_Parity_Run_14628__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306749 12/14/09 00:31 apar End_of_Parity_Run_14627__CODA_Config=Pvdis
  • 306748 12/14/09 00:26 dalton Compton problems continue
  • Sunday

  • 306746 12/13/09 23:58 S. Sirca Swing shift SUMMARY
  • 306745 12/13/09 23:54 miham Change in Right Arm S2 plots
  • 306744 12/13/09 23:36 apar Start_Parity_Run_14627__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306739 12/13/09 23:33 apar End_of_Parity_Run_14626__CODA_Config=Pvdis
  • 306738 12/13/09 22:37 apar Start_Parity_Run_14626__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306733 12/13/09 22:35 apar End_of_Parity_Run_14625__CODA_Config=Pvdis
  • 306732 12/13/09 22:33 M Friend Compton problems
  • 306731 12/13/09 21:39 auto_snap Left Arm HV snapshot
  • 306730 12/13/09 21:39 auto_snap Happex Charge Asym snapshot
  • 306729 12/13/09 21:39 auto_snap Hall A Tools snapshot
  • 306728 12/13/09 21:39 auto_snap Compton Polarimeter snapshot
  • 306727 12/13/09 21:39 auto_snap Beam Position and Current snapshot
  • 306726 12/13/09 21:39 auto_snap Accelerator Energies snapshot
  • 306723 12/13/09 21:39 apar Start_Parity_Run_14625__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306720 12/13/09 21:31 apar Start_Parity_Run_14624__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306715 12/13/09 21:27 apar End_of_Parity_Run_14623__CODA_Config=Pvdis
  • 306714 12/13/09 20:31 apar Start_Parity_Run_14623__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306711 12/13/09 20:28 apar Halfwave plate OUT, Slug 23
  • 306708 12/13/09 20:26 apar End_of_Parity_Run_14622__CODA_Config=Pvdis
  • 306707 12/13/09 19:30 apar Start_Parity_Run_14622__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306702 12/13/09 19:28 apar End_of_Parity_Run_14621__CODA_Config=Pvdis
  • 306701 12/13/09 18:46 Abdurahim Target screens
  • 306700 12/13/09 18:32 apar Start_Parity_Run_14621__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306696 12/13/09 18:30 apar End_of_Parity_Run_14620__CODA_Config=Pvdis
  • 306695 12/13/09 17:52 pking Run coordinator daily summary
  • 306694 12/13/09 17:39 auto_snap Left Arm HV snapshot
  • 306693 12/13/09 17:39 auto_snap Happex Charge Asym snapshot
  • 306692 12/13/09 17:39 auto_snap Hall A Tools snapshot
  • 306691 12/13/09 17:39 auto_snap Compton Polarimeter snapshot
  • 306690 12/13/09 17:39 auto_snap Beam Position and Current snapshot
  • 306689 12/13/09 17:39 auto_snap Accelerator Energies snapshot
  • 306688 12/13/09 17:34 apar Start_Parity_Run_14620__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306683 12/13/09 17:31 apar End_of_Parity_Run_14619__CODA_Config=Pvdis
  • 306682 12/13/09 17:09 M Friend Compton laser tune and vertical scan
  • 306679 12/13/09 16:35 apar Start_Parity_Run_14619__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306676 12/13/09 16:33 apar End_of_Parity_Run_14618__CODA_Config=Pvdis
  • 306670 12/13/09 15:38 apar Start_Parity_Run_14618__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306669 12/13/09 15:36 apar End_of_Parity_Run_14617__CODA_Config=Pvdis
  • 306668 12/13/09 15:34 kbrider target screens
  • 306667 12/13/09 15:19 pking Left arm slow controls crate reset
  • 306664 12/13/09 14:41 apar Start_Parity_Run_14617__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306661 12/13/09 14:39 apar End_of_Parity_Run_14616__CODA_Config=Pvdis
  • 306660 12/13/09 13:48 kbrider compton is having trouble locking
  • 306657 12/13/09 13:44 apar Start_Parity_Run_14616__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306654 12/13/09 13:42 apar End_of_Parity_Run_14615__CODA_Config=Pvdis
  • 306653 12/13/09 13:39 auto_snap Left Arm HV snapshot
  • 306652 12/13/09 13:39 auto_snap Happex Charge Asym snapshot
  • 306651 12/13/09 13:38 auto_snap Hall A Tools snapshot
  • 306650 12/13/09 13:38 auto_snap Compton Polarimeter snapshot
  • 306649 12/13/09 13:38 auto_snap Beam Position and Current snapshot
  • 306648 12/13/09 13:38 auto_snap Accelerator Energies snapshot
  • 306647 12/13/09 13:07 kbrider target screens
  • 306643 12/13/09 12:46 apar Start_Parity_Run_14615__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306642 12/13/09 12:44 apar End_of_Parity_Run_14614__CODA_Config=Pvdis
  • 306639 12/13/09 11:50 apar Start_Parity_Run_14614__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306636 12/13/09 11:47 apar End_of_Parity_Run_14613__CODA_Config=Pvdis
  • 306633 12/13/09 10:53 apar Start_Parity_Run_14613__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306630 12/13/09 10:50 apar End_of_Parity_Run_14612__CODA_Config=Pvdis
  • 306627 12/13/09 09:55 apar Start_Parity_Run_14612__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306624 12/13/09 09:52 apar End_of_Parity_Run_14611__CODA_Config=Pvdis
  • 306623 12/13/09 09:51 R. Subedi iocha4 dead, no worry now
  • 306622 12/13/09 09:38 auto_snap Left Arm HV snapshot
  • 306621 12/13/09 09:38 auto_snap Happex Charge Asym snapshot
  • 306620 12/13/09 09:38 auto_snap Hall A Tools snapshot
  • 306619 12/13/09 09:38 auto_snap Compton Polarimeter snapshot
  • 306618 12/13/09 09:38 auto_snap Beam Position and Current snapshot
  • 306617 12/13/09 09:38 auto_snap Accelerator Energies snapshot
  • 306614 12/13/09 08:57 apar Start_Parity_Run_14611__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306613 12/13/09 08:52 apar Start_Parity_Run_14610__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306608 12/13/09 08:44 apar Start_Parity_Run_14609__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306605 12/13/09 08:42 apar End_of_Parity_Run_14608__CODA_Config=Pvdis
  • 306603 12/13/09 08:00 Deng owl shift summary
  • 306600 12/13/09 07:47 apar Start_Parity_Run_14608__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306597 12/13/09 07:44 apar End_of_Parity_Run_14607__CODA_Config=Pvdis
  • 306594 12/13/09 06:49 apar Start_Parity_Run_14607__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306591 12/13/09 06:47 apar End_of_Parity_Run_14606__CODA_Config=Pvdis
  • 306590 12/13/09 06:45 meziani Target charts
  • 306589 12/13/09 06:42 meziani target screen
  • 306588 12/13/09 06:13 Deng Hall A tools screen shot
  • 306587 12/13/09 06:09 J. Rubin Raster Size
  • 306586 12/13/09 05:51 apar Start_Parity_Run_14606__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306581 12/13/09 05:49 apar End_of_Parity_Run_14605__CODA_Config=Pvdis
  • 306580 12/13/09 05:38 auto_snap Left Arm HV snapshot
  • 306579 12/13/09 05:38 auto_snap Happex Charge Asym snapshot
  • 306578 12/13/09 05:38 auto_snap Hall A Tools snapshot
  • 306577 12/13/09 05:38 auto_snap Compton Polarimeter snapshot
  • 306576 12/13/09 05:38 auto_snap Beam Position and Current snapshot
  • 306575 12/13/09 05:38 auto_snap Accelerator Energies snapshot
  • 306574 12/13/09 04:53 apar Start_Parity_Run_14605__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306569 12/13/09 04:51 apar End_of_Parity_Run_14604__CODA_Config=Pvdis
  • 306568 12/13/09 03:55 apar Start_Parity_Run_14604__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306563 12/13/09 03:53 apar End_of_Parity_Run_14603__CODA_Config=Pvdis
  • 306562 12/13/09 03:47 Deng Q2/W plots for Dec12
  • 306561 12/13/09 02:57 apar Start_Parity_Run_14603__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306556 12/13/09 02:54 apar End_of_Parity_Run_14602__CODA_Config=Pvdis
  • 306555 12/13/09 02:30 A. Camsonne Compton HV changed to 1650 V and VNC restarted
  • 306554 12/13/09 01:58 apar Start_Parity_Run_14602__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306549 12/13/09 01:56 apar End_of_Parity_Run_14601__CODA_Config=Pvdis
  • 306548 12/13/09 01:44 Deng screen shot of HRS scaler rates for production data
  • 306547 12/13/09 01:37 auto_snap Left Arm HV snapshot
  • 306546 12/13/09 01:37 auto_snap Happex Charge Asym snapshot
  • 306545 12/13/09 01:37 auto_snap Hall A Tools snapshot
  • 306544 12/13/09 01:37 auto_snap Compton Polarimeter snapshot
  • 306543 12/13/09 01:37 auto_snap Beam Position and Current snapshot
  • 306542 12/13/09 01:37 auto_snap Accelerator Energies snapshot
  • 306541 12/13/09 01:33 Deng Beam strip chart
  • 306540 12/13/09 01:08 D.Wang right arm preshower 34 keeps degrading
  • 306537 12/13/09 01:00 apar Start_Parity_Run_14601__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306535 12/13/09 00:58 D.Wang This problem for group 4 started after we changed the HV for preshower 34
  • 306534 12/13/09 00:58 apar End_of_Parity_Run_14600__CODA_Config=Pvdis
  • 306533 12/13/09 00:16 D.Wang Rebooted ROC 1
  • 306529 12/13/09 00:06 miham SWING Shift Summary (12/12/09)
  • 306527 12/13/09 00:02 apar Start_Parity_Run_14600__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • Saturday

  • 306525 12/12/09 23:46 apar End_of_Parity_Run_14599__CODA_Config=Pvdis
  • 306524 12/12/09 23:45 aniol target screens
  • 306523 12/12/09 23:09 apar Start_Parity_Run_14599__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306518 12/12/09 23:07 apar End_of_Parity_Run_14598__CODA_Config=Pvdis
  • 306517 12/12/09 22:15 miham Spot++ check for Production run #26182
  • 306516 12/12/09 22:11 apar Start_Parity_Run_14598__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306511 12/12/09 22:09 apar End_of_Parity_Run_14597__CODA_Config=Pvdis
  • 306510 12/12/09 21:37 auto_snap Left Arm HV snapshot
  • 306509 12/12/09 21:37 auto_snap Happex Charge Asym snapshot
  • 306508 12/12/09 21:37 auto_snap Hall A Tools snapshot
  • 306507 12/12/09 21:37 auto_snap Compton Polarimeter snapshot
  • 306506 12/12/09 21:37 auto_snap Beam Position and Current snapshot
  • 306505 12/12/09 21:37 auto_snap Accelerator Energies snapshot
  • 306504 12/12/09 21:13 apar Start_Parity_Run_14597__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306499 12/12/09 21:11 apar End_of_Parity_Run_14596__CODA_Config=Pvdis
  • 306498 12/12/09 21:09 miham Compton HV trip
  • 306497 12/12/09 20:18 miham Roc 25 Problems during run #14595
  • 306496 12/12/09 20:16 R. Subedi Re: Strange kink in R-arm fastbus TDC (Gr 3-4)
  • 306495 12/12/09 20:15 apar Start_Parity_Run_14596__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306490 12/12/09 19:48 apar Start_Parity_Run_14595__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306486 12/12/09 19:46 apar End_of_Parity_Run_14594__CODA_Config=Pvdis
  • 306484 12/12/09 19:22 apar Start_Parity_Run_14594__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306483 12/12/09 19:20 apar End_of_Parity_Run_14593__CODA_Config=Pvdis
  • 306482 12/12/09 19:07 Abdurahim Strange kink in R-arm fastbus TDC (Gr 3-4)
  • 306481 12/12/09 18:55 pking Status update: back to production running.
  • 306480 12/12/09 18:33 aniol target screens
  • 306479 12/12/09 18:29 miham Spot++ check for First Production run #26178
  • 306478 12/12/09 18:29 Abdurahim Q^2 measurement result from right HRS
  • 306477 12/12/09 18:26 apar Start_Parity_Run_14593__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306476 12/12/09 18:25 miham Beam current set to 105uA, Starting Production
  • 306473 12/12/09 18:21 Abdurahim Q^2 measurement result fro LHRS
  • 306472 12/12/09 18:15 miham VDC Chambers are OFF for Production running
  • 306471 12/12/09 18:13 miham Results of the HARP scan
  • 306468 12/12/09 18:06 miham Spot++ check for run #26177
  • 306465 12/12/09 17:58 Aniol loop 1, 20cm LD2 in beam
  • 306464 12/12/09 17:57 miham VDC Chambers are on for Low Current measurement
  • 306463 12/12/09 17:56 A. Camsonne Compton test spare CPU
  • 306462 12/12/09 17:37 auto_snap Left Arm HV snapshot
  • 306461 12/12/09 17:37 auto_snap Happex Charge Asym snapshot
  • 306460 12/12/09 17:37 auto_snap Hall A Tools snapshot
  • 306459 12/12/09 17:36 auto_snap Compton Polarimeter snapshot
  • 306458 12/12/09 17:36 auto_snap Beam Position and Current snapshot
  • 306457 12/12/09 17:36 auto_snap Accelerator Energies snapshot
  • 306456 12/12/09 16:46 pking Run coordinator daily summary
  • 306455 12/12/09 16:01 S. Sirca Day shift SUMMARY
  • 306452 12/12/09 16:00 apar End_of_Parity_Run_14592__CODA_Config=Pvdis
  • 306450 12/12/09 15:58 kbrider target screens
  • 306449 12/12/09 15:05 apar Start_Parity_Run_14592__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306444 12/12/09 15:03 apar End_of_Parity_Run_14591__CODA_Config=Pvdis
  • 306443 12/12/09 14:07 apar Start_Parity_Run_14591__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306440 12/12/09 13:36 auto_snap Left Arm HV snapshot
  • 306439 12/12/09 13:36 auto_snap Happex Charge Asym snapshot
  • 306438 12/12/09 13:36 auto_snap Hall A Tools snapshot
  • 306437 12/12/09 13:36 auto_snap Compton Polarimeter snapshot
  • 306436 12/12/09 13:36 auto_snap Beam Position and Current snapshot
  • 306435 12/12/09 13:36 auto_snap Accelerator Energies snapshot
  • 306432 12/12/09 13:21 apar End_of_Parity_Run_14590__CODA_Config=Pvdis
  • 306431 12/12/09 12:25 apar Start_Parity_Run_14590__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306428 12/12/09 09:36 auto_snap Left Arm HV snapshot
  • 306427 12/12/09 09:36 auto_snap Happex Charge Asym snapshot
  • 306426 12/12/09 09:36 auto_snap Hall A Tools snapshot
  • 306425 12/12/09 09:36 auto_snap Compton Polarimeter snapshot
  • 306424 12/12/09 09:36 auto_snap Beam Position and Current snapshot
  • 306423 12/12/09 09:36 auto_snap Accelerator Energies snapshot
  • 306422 12/12/09 09:09 kbrider target screens
  • 306421 12/12/09 08:10 kbrider Target move
  • 306418 12/12/09 08:04 apar End_of_Parity_Run_14589__CODA_Config=Pvdis
  • 306417 12/12/09 08:03 V. Sulkosky Owl Shift Summary
  • 306415 12/12/09 07:43 apar Start_Parity_Run_14589__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306410 12/12/09 07:41 apar End_of_Parity_Run_14588__CODA_Config=Pvdis
  • 306409 12/12/09 06:45 apar Start_Parity_Run_14588__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306404 12/12/09 06:44 apar End_of_Parity_Run_14587__CODA_Config=Pvdis
  • 306403 12/12/09 06:43 Meziani Target Interface Snapshot
  • 306402 12/12/09 06:42 meziani Target Charts
  • 306401 12/12/09 06:32 V. Sulkosky Compton Laser Off
  • 306400 12/12/09 05:59 Deng spot++ check for production runs
  • 306399 12/12/09 05:51 V. Sulkosky Beam Trips
  • 306398 12/12/09 05:48 apar Start_Parity_Run_14587__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306393 12/12/09 05:47 apar End_of_Parity_Run_14586__CODA_Config=Pvdis
  • 306392 12/12/09 05:36 auto_snap Left Arm HV snapshot
  • 306391 12/12/09 05:35 auto_snap Happex Charge Asym snapshot
  • 306390 12/12/09 05:35 auto_snap Hall A Tools snapshot
  • 306389 12/12/09 05:35 auto_snap Compton Polarimeter snapshot
  • 306388 12/12/09 05:35 auto_snap Beam Position and Current snapshot
  • 306387 12/12/09 05:35 auto_snap Accelerator Energies snapshot
  • 306384 12/12/09 04:51 apar Start_Parity_Run_14586__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306381 12/12/09 04:50 apar End_of_Parity_Run_14585__CODA_Config=Pvdis
  • 306380 12/12/09 03:54 apar Start_Parity_Run_14585__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306375 12/12/09 03:53 D.Wang compton gain changed
  • 306374 12/12/09 03:53 apar End_of_Parity_Run_14584__CODA_Config=Pvdis
  • 306373 12/12/09 03:17 V. Sulkosky Compton Edge?
  • 306372 12/12/09 03:05 V. Sulkosky BPM Stripchart
  • 306371 12/12/09 02:57 apar Start_Parity_Run_14584__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306366 12/12/09 02:56 apar End_of_Parity_Run_14583__CODA_Config=Pvdis
  • 306365 12/12/09 02:23 V. Sulkosky Hall A Tools
  • 306364 12/12/09 02:05 Deng HWP IN, Slug 22
  • 306363 12/12/09 02:02 V. Sulkosky Set RWP to 1200
  • 306362 12/12/09 02:00 apar Start_Parity_Run_14583__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306359 12/12/09 01:53 apar Halfwave plate IN, Slug 22
  • 306356 12/12/09 01:51 apar End_of_Parity_Run_14582__CODA_Config=Pvdis
  • 306355 12/12/09 01:35 auto_snap Left Arm HV snapshot
  • 306354 12/12/09 01:35 auto_snap Happex Charge Asym snapshot
  • 306353 12/12/09 01:35 auto_snap Hall A Tools snapshot
  • 306352 12/12/09 01:35 auto_snap Compton Polarimeter snapshot
  • 306351 12/12/09 01:35 auto_snap Beam Position and Current snapshot
  • 306350 12/12/09 01:35 auto_snap Accelerator Energies snapshot
  • 306349 12/12/09 01:03 Deng Dithering was off for compton epics reboot
  • 306348 12/12/09 01:01 D.Wang compton epics reboot
  • 306345 12/12/09 00:55 apar Start_Parity_Run_14582__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306342 12/12/09 00:52 apar End_of_Parity_Run_14581__CODA_Config=Pvdis
  • 306341 12/12/09 00:47 apar Start_Parity_Run_14581__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306338 12/12/09 00:43 apar End_of_Parity_Run_14580__CODA_Config=Pvdis
  • 306336 12/12/09 00:30 V. Sulkosky Compton EPICS Not Reporting
  • 306335 12/12/09 00:12 apar Start_Parity_Run_14580__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306334 12/12/09 00:10 apar End_of_Parity_Run_14579__CODA_Config=Pvdis
  • 306329 12/12/09 00:03 miham SWING Shift Summary (12/11/09)
  • Friday

  • 306327 12/11/09 23:54 miham Spot++ check for production run #26163
  • 306326 12/11/09 23:12 apar Start_Parity_Run_14579__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306325 12/11/09 23:06 apar Start_Parity_Run_14578__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306321 12/11/09 23:03 apar End_of_Parity_Run_14577__CODA_Config=Pvdis
  • 306320 12/11/09 22:21 A. Camsonne Standard Compton DAQ back up
  • 306319 12/11/09 22:08 apar Start_Parity_Run_14577__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306314 12/11/09 22:05 apar End_of_Parity_Run_14576__CODA_Config=Pvdis
  • 306313 12/11/09 21:35 auto_snap Left Arm HV snapshot
  • 306312 12/11/09 21:35 auto_snap Happex Charge Asym snapshot
  • 306311 12/11/09 21:35 auto_snap Hall A Tools snapshot
  • 306310 12/11/09 21:35 auto_snap Compton Polarimeter snapshot
  • 306309 12/11/09 21:34 auto_snap Beam Position and Current snapshot
  • 306308 12/11/09 21:34 auto_snap Accelerator Energies snapshot
  • 306307 12/11/09 21:32 ZAFAR AHMED Q^2
  • 306306 12/11/09 21:30 Armstrong Target screen shoots
  • 306305 12/11/09 21:09 apar Start_Parity_Run_14576__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306302 12/11/09 21:04 miham VDC Chambers are OFF.
  • 306297 12/11/09 20:27 miham VDC Chambers are on for Q2 measurement
  • 306294 12/11/09 20:18 apar End_of_Parity_Run_14575__CODA_Config=Pvdis
  • 306293 12/11/09 19:23 apar Start_Parity_Run_14575__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306288 12/11/09 19:20 apar End_of_Parity_Run_14574__CODA_Config=Pvdis
  • 306287 12/11/09 18:31 miham Beam is a bit better now
  • 306286 12/11/09 18:24 apar Start_Parity_Run_14574__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306281 12/11/09 18:22 apar End_of_Parity_Run_14573__CODA_Config=Pvdis
  • 306280 12/11/09 17:41 D.Wang/R.Sub HV adjust for right arm preshower 34
  • 306279 12/11/09 17:34 auto_snap Left Arm HV snapshot
  • 306278 12/11/09 17:34 auto_snap Happex Charge Asym snapshot
  • 306277 12/11/09 17:34 auto_snap Hall A Tools snapshot
  • 306276 12/11/09 17:34 auto_snap Compton Polarimeter snapshot
  • 306275 12/11/09 17:34 auto_snap Beam Position and Current snapshot
  • 306274 12/11/09 17:34 auto_snap Accelerator Energies snapshot
  • 306273 12/11/09 17:26 apar Start_Parity_Run_14573__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306270 12/11/09 17:25 miham Beam with many trips
  • 306267 12/11/09 17:21 apar End_of_Parity_Run_14572__CODA_Config=Pvdis
  • 306254 12/11/09 16:26 apar Start_Parity_Run_14572__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306251 12/11/09 16:20 Armstrong Target fluctuations OK
  • 306250 12/11/09 16:20 Armstrong Target fluctuations OK
  • 306247 12/11/09 16:06 apar End_of_Parity_Run_14571__CODA_Config=Pvdis
  • 306246 12/11/09 16:01 S. Sirca Day shift SUMMARY
  • 306244 12/11/09 15:58 A. Camsonne Compton issue with Lecroy scaler initialization
  • 306243 12/11/09 15:21 W. Deconinck Compton coda reboot
  • 306242 12/11/09 15:13 apar Start_Parity_Run_14571__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306237 12/11/09 15:10 apar End_of_Parity_Run_14570__CODA_Config=Pvdis
  • 306236 12/11/09 14:23 W. Deconinck Target tools screen capture
  • 306235 12/11/09 14:14 apar Start_Parity_Run_14570__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306230 12/11/09 14:11 apar End_of_Parity_Run_14569__CODA_Config=Pvdis
  • 306229 12/11/09 13:51 A. Camsonne Compton rates ok
  • 306228 12/11/09 13:48 R. Michaels planned MSS power outage 7:00 am
  • 306227 12/11/09 13:34 auto_snap Left Arm HV snapshot
  • 306226 12/11/09 13:34 auto_snap Happex Charge Asym snapshot
  • 306225 12/11/09 13:34 auto_snap Hall A Tools snapshot
  • 306224 12/11/09 13:34 auto_snap Compton Polarimeter snapshot
  • 306223 12/11/09 13:34 auto_snap Beam Position and Current snapshot
  • 306222 12/11/09 13:34 auto_snap Accelerator Energies snapshot
  • 306219 12/11/09 13:16 apar Start_Parity_Run_14569__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306216 12/11/09 13:13 apar End_of_Parity_Run_14568__CODA_Config=Pvdis
  • 306215 12/11/09 13:10 A. Camsonne Checking with MCC for Compton photon rates
  • 306214 12/11/09 12:46 W. Deconinck I meant heater power
  • 306213 12/11/09 12:30 W. Deconinck More flucts on target temp than usual
  • 306212 12/11/09 12:26 W. Deconinck Compton: high background rates
  • 306211 12/11/09 12:20 S. Sirca Spot++ with 105uA on LD2, 6x6 MCC raster
  • 306210 12/11/09 12:19 S. Sirca Raw rates with 105uA on LD2, 6x6 MCC raster
  • 306206 12/11/09 12:13 S. Sirca Spot++ with 2uA on LD2, 6x6 MCC raster
  • 306204 12/11/09 12:12 S. Sirca Raw rates with 2uA on LD2, 6x6 MCC raster
  • 306200 12/11/09 12:01 S. Sirca Harp scan results
  • 306199 12/11/09 12:00 W. Deconinck Moving target to liquid deuterium
  • 306198 12/11/09 11:38 pking Run coordinator daily summary
  • 306197 12/11/09 11:36 W. Deconinck Moving target to BeO position
  • 306196 12/11/09 10:17 W. Deconinck Compton coda reboot
  • 306195 12/11/09 09:34 auto_snap Left Arm HV snapshot
  • 306194 12/11/09 09:34 auto_snap Happex Charge Asym snapshot
  • 306193 12/11/09 09:33 auto_snap Hall A Tools snapshot
  • 306192 12/11/09 09:33 auto_snap Compton Polarimeter snapshot
  • 306191 12/11/09 09:33 auto_snap Beam Position and Current snapshot
  • 306190 12/11/09 09:33 auto_snap Accelerator Energies snapshot
  • 306189 12/11/09 08:52 S. Sirca Beam OFF, valve problem in Souoth Linac, access needed
  • 306188 12/11/09 08:49 pking Brief RC status report
  • 306185 12/11/09 08:46 apar End_of_Parity_Run_14567__CODA_Config=Pvdis
  • 306184 12/11/09 08:11 W. Deconinck Compton coda reboot
  • 306183 12/11/09 08:08 V. Sulkosky Owl Shift Summary
  • 306180 12/11/09 07:52 apar Start_Parity_Run_14567__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306176 12/11/09 07:50 apar End_of_Parity_Run_14566__CODA_Config=Pvdis
  • 306175 12/11/09 07:49 V. Sulkosky Beam Trips Worse
  • 306174 12/11/09 07:41 Deng spot++ check for production run
  • 306173 12/11/09 07:38 Deng Q2 and W plots on single 12C target
  • 306172 12/11/09 07:27 Reimer Target Screen Shots
  • 306171 12/11/09 07:14 A. Camsonne Compton DAQ up again
  • 306170 12/11/09 06:54 apar Start_Parity_Run_14566__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306165 12/11/09 06:52 apar End_of_Parity_Run_14565__CODA_Config=Pvdis
  • 306164 12/11/09 06:49 V. Sulkosky PVDIS Run 14564 DAQ Problem
  • 306163 12/11/09 06:48 A. Camsonne Compton standard DAQ down
  • 306162 12/11/09 06:08 V. Sulkosky PVDIS DAQ Crash
  • 306161 12/11/09 05:57 apar Start_Parity_Run_14565__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306156 12/11/09 05:39 apar Start_Parity_Run_14564__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306153 12/11/09 05:38 V. Sulkosky BPM Stripchart
  • 306150 12/11/09 05:37 apar End_of_Parity_Run_14563__CODA_Config=Pvdis
  • 306149 12/11/09 05:33 auto_snap Left Arm HV snapshot
  • 306148 12/11/09 05:33 auto_snap Happex Charge Asym snapshot
  • 306147 12/11/09 05:33 auto_snap Hall A Tools snapshot
  • 306146 12/11/09 05:33 auto_snap Compton Polarimeter snapshot
  • 306145 12/11/09 05:33 auto_snap Beam Position and Current snapshot
  • 306144 12/11/09 05:33 auto_snap Accelerator Energies snapshot
  • 306143 12/11/09 04:42 apar Start_Parity_Run_14563__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306138 12/11/09 04:40 apar End_of_Parity_Run_14562__CODA_Config=Pvdis
  • 306137 12/11/09 04:14 M Friend Compton beam position tune
  • 306136 12/11/09 04:10 V. Sulkosky Hall A Tools Screen
  • 306135 12/11/09 04:10 V. Sulkosky BCM Chart
  • 306134 12/11/09 03:51 M Friend Compton cavity tune
  • 306133 12/11/09 03:49 V. Sulkosky Back to Production
  • 306132 12/11/09 03:47 Deng Beam modulation and feedback on
  • 306131 12/11/09 03:45 apar Start_Parity_Run_14562__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306128 12/11/09 03:36 Deng spot++ check, mcc 6x6, LD2 target
  • 306125 12/11/09 03:31 Deng spot++ on LD2 target, 4x4 mcc unit
  • 306122 12/11/09 03:23 V. Sulkosky Final Harp Scans
  • 306121 12/11/09 03:19 Reimer Target lD2
  • 306120 12/11/09 03:18 Deng spot ++ on BeO target
  • 306117 12/11/09 03:06 Reimer Target BeO
  • 306116 12/11/09 02:55 V. Sulkosky Initial Harp Scans
  • 306115 12/11/09 02:47 Deng remove golden run settings for Q2/W measurement and others
  • 306114 12/11/09 02:38 V. Sulkosky Beam is Back!
  • 306111 12/11/09 02:33 apar End_of_Parity_Run_14561__CODA_Config=Pvdis
  • 306110 12/11/09 02:20 apar Start_Parity_Run_14561__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306105 12/11/09 02:16 apar End_of_Parity_Run_14560__CODA_Config=Pvdis
  • 306104 12/11/09 02:13 Deng Q2 and W plots on Al-Dummy, double peak on Q2 plot
  • 306103 12/11/09 02:07 M Friend Compton photon collimator
  • 306102 12/11/09 01:50 V. Sulkosky Controlled Access
  • 306101 12/11/09 01:40 Deng Q2 and W plots on multifoil 12C target
  • 306100 12/11/09 01:35 V. Sulkosky Beam Update
  • 306099 12/11/09 01:33 auto_snap Left Arm HV snapshot
  • 306098 12/11/09 01:33 auto_snap Happex Charge Asym snapshot
  • 306097 12/11/09 01:33 auto_snap Hall A Tools snapshot
  • 306096 12/11/09 01:33 auto_snap Compton Polarimeter snapshot
  • 306095 12/11/09 01:33 auto_snap Beam Position and Current snapshot
  • 306094 12/11/09 01:33 auto_snap Accelerator Energies snapshot
  • 306093 12/11/09 01:20 apar Start_Parity_Run_14560__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306088 12/11/09 01:17 apar End_of_Parity_Run_14559__CODA_Config=Pvdis
  • 306086 12/11/09 00:22 apar Start_Parity_Run_14559__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=Empty
  • 306083 12/11/09 00:03 miham SWING Shift Summary (12/10/09)
  • Thursday

  • 306081 12/10/09 23:57 jpchen target to Empty
  • 306080 12/10/09 23:52 jpchen target charts and screen
  • 306079 12/10/09 22:41 miham A Serious cryo problem
  • 306077 12/10/09 22:01 miham Cryo modules problem
  • 306074 12/10/09 21:57 apar End_of_Parity_Run_14558__CODA_Config=Pvdis
  • 306073 12/10/09 21:52 apar Start_Parity_Run_14558__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=dummy run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm Al Dummy
  • 306068 12/10/09 21:51 apar End_of_Parity_Run_14557__CODA_Config=Pvdis
  • 306067 12/10/09 21:32 auto_snap Left Arm HV snapshot
  • 306066 12/10/09 21:32 auto_snap Happex Charge Asym snapshot
  • 306065 12/10/09 21:32 auto_snap Hall A Tools snapshot
  • 306064 12/10/09 21:32 auto_snap Compton Polarimeter snapshot
  • 306063 12/10/09 21:32 auto_snap Beam Position and Current snapshot
  • 306062 12/10/09 21:32 auto_snap Accelerator Energies snapshot
  • 306061 12/10/09 20:55 apar Start_Parity_Run_14557__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=dummy run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm Al Dummy
  • 306056 12/10/09 20:54 apar End_of_Parity_Run_14556__CODA_Config=Pvdis
  • 306055 12/10/09 20:21 miham Spot++ check for Al-dummy run
  • 306049 12/10/09 19:58 apar Start_Parity_Run_14556__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=dummy run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm Al Dummy
  • 306047 12/10/09 19:57 miham Beam current during Optics runs
  • 306046 12/10/09 19:54 pking Run coordinator daily summary
  • 306045 12/10/09 19:52 miham VDC Chambers are OFF
  • 306042 12/10/09 19:38 miham Spot++ check for Al-Dummy runs with (6x6MCC)raster
  • 306041 12/10/09 19:35 jpchen moved to dummy target
  • 306037 12/10/09 19:26 miham Spot++ check for Single C target (6x6MCC)
  • 306030 12/10/09 19:03 jpchen Moved to single 12C target
  • 306027 12/10/09 18:58 miham Spot++ check for run with Multi. 12C target with (6x6MCC) raster
  • 306022 12/10/09 18:43 miham Scaler Snap Shot for the multifoil 12C runs
  • 306021 12/10/09 18:41 miham VDC Chambers are ON!
  • 306018 12/10/09 18:32 jpchen to optics target
  • 306015 12/10/09 18:24 apar End_of_Parity_Run_14555__CODA_Config=Pvdis
  • 306014 12/10/09 17:32 auto_snap Left Arm HV snapshot
  • 306013 12/10/09 17:32 auto_snap Happex Charge Asym snapshot
  • 306012 12/10/09 17:32 auto_snap Hall A Tools snapshot
  • 306011 12/10/09 17:32 auto_snap Compton Polarimeter snapshot
  • 306010 12/10/09 17:32 auto_snap Beam Position and Current snapshot
  • 306009 12/10/09 17:32 auto_snap Accelerator Energies snapshot
  • 306008 12/10/09 17:29 apar Start_Parity_Run_14555__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 306003 12/10/09 17:27 apar End_of_Parity_Run_14554__CODA_Config=Pvdis
  • 306002 12/10/09 16:36 miham Spot++ check on LD2 target with 105uA beam (6x6MCC)
  • 306001 12/10/09 16:32 apar Start_Parity_Run_14554__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305997 12/10/09 16:27 miham Spot++ check on LD2 target (6x6MCC) raster
  • 305996 12/10/09 16:25 jpchen Target moved to LD2
  • 305991 12/10/09 15:59 S. Sirca Day shift SUMMARY
  • 305990 12/10/09 15:38 O. Glamazdin Moller results
  • 305989 12/10/09 14:53 S. Sirca Alarm handler silenced 1hr, remained silent forever
  • 305988 12/10/09 13:32 auto_snap Left Arm HV snapshot
  • 305987 12/10/09 13:32 auto_snap Happex Charge Asym snapshot
  • 305986 12/10/09 13:32 auto_snap Hall A Tools snapshot
  • 305985 12/10/09 13:32 auto_snap Compton Polarimeter snapshot
  • 305984 12/10/09 13:31 auto_snap Beam Position and Current snapshot
  • 305983 12/10/09 13:31 auto_snap Accelerator Energies snapshot
  • 305982 12/10/09 13:10 W. Deconinck Target tools snapshot
  • 305981 12/10/09 09:31 auto_snap Left Arm HV snapshot
  • 305980 12/10/09 09:31 auto_snap Happex Charge Asym snapshot
  • 305979 12/10/09 09:31 auto_snap Hall A Tools snapshot
  • 305978 12/10/09 09:31 auto_snap Compton Polarimeter snapshot
  • 305977 12/10/09 09:31 auto_snap Beam Position and Current snapshot
  • 305976 12/10/09 09:31 auto_snap Accelerator Energies snapshot
  • 305975 12/10/09 09:12 W. Deconinck Moving target to empty position for Moller measurement
  • 305972 12/10/09 09:10 apar End_of_Parity_Run_14553__CODA_Config=Pvdis
  • 305971 12/10/09 08:56 apar Start_Parity_Run_14553__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305966 12/10/09 08:54 apar End_of_Parity_Run_14552__CODA_Config=Pvdis
  • 305965 12/10/09 08:23 D.Wang counting house sync signal was not so good
  • 305964 12/10/09 08:12 D.Wang Reset TS11 and ROC 3
  • 305962 12/10/09 08:08 M Friend Owl Shift Summary
  • 305961 12/10/09 08:05 M Friend Compton finger scans
  • 305960 12/10/09 07:59 apar Start_Parity_Run_14552__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305953 12/10/09 07:11 apar End_of_Parity_Run_14550__CODA_Config=Pvdis
  • 305952 12/10/09 06:57 apar Start_Parity_Run_14550__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305950 12/10/09 06:52 apar End_of_Parity_Run_14549__CODA_Config=Pvdis
  • 305947 12/10/09 05:57 apar Start_Parity_Run_14549__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305944 12/10/09 05:52 apar End_of_Parity_Run_14548__CODA_Config=Pvdis
  • 305943 12/10/09 05:48 V BELLINI Beam: x at target
  • 305942 12/10/09 05:42 V BELLINI Beam x at target
  • 305941 12/10/09 05:36 V BELLINI Anomalous Beam Position Run
  • 305940 12/10/09 05:31 auto_snap Left Arm HV snapshot
  • 305939 12/10/09 05:31 auto_snap Happex Charge Asym snapshot
  • 305938 12/10/09 05:31 auto_snap Hall A Tools snapshot
  • 305937 12/10/09 05:31 auto_snap Compton Polarimeter snapshot
  • 305936 12/10/09 05:31 auto_snap Beam Position and Current snapshot
  • 305935 12/10/09 05:31 auto_snap Accelerator Energies snapshot
  • 305934 12/10/09 04:57 apar Start_Parity_Run_14548__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305929 12/10/09 04:53 apar End_of_Parity_Run_14547__CODA_Config=Pvdis
  • 305928 12/10/09 04:43 M Friend Hall A tools and bcm snapshot
  • 305925 12/10/09 03:57 apar Start_Parity_Run_14547__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305922 12/10/09 03:54 apar End_of_Parity_Run_14546__CODA_Config=Pvdis
  • 305921 12/10/09 03:00 V BELLINI CODA Restarted and ROC's rebooted
  • 305919 12/10/09 02:58 apar Start_Parity_Run_14546__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305915 12/10/09 02:37 V BELLINI CODA Restart for HAPPEX DAQ
  • 305912 12/10/09 02:28 apar Start_Parity_Run_14545__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305911 12/10/09 02:26 apar Start_Parity_Run_14544__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305910 12/10/09 02:26 apar End_of_Parity_Run_14544__CODA_Config=Pvdis
  • 305905 12/10/09 01:31 auto_snap Left Arm HV snapshot
  • 305904 12/10/09 01:31 auto_snap Happex Charge Asym snapshot
  • 305903 12/10/09 01:30 auto_snap Hall A Tools snapshot
  • 305902 12/10/09 01:30 auto_snap Compton Polarimeter snapshot
  • 305901 12/10/09 01:30 auto_snap Beam Position and Current snapshot
  • 305900 12/10/09 01:30 auto_snap Accelerator Energies snapshot
  • 305899 12/10/09 01:29 apar Start_Parity_Run_14543__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305896 12/10/09 01:26 apar End_of_Parity_Run_14542__CODA_Config=Pvdis
  • 305895 12/10/09 01:14 apar End_of_Parity_Run_14541__CODA_Config=Pvdis
  • 305894 12/10/09 00:56 M Friend Compton vertical scan
  • 305893 12/10/09 00:30 apar Start_Parity_Run_14542__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305888 12/10/09 00:15 M Friend Compton laser tune
  • 305887 12/10/09 00:05 miham SWING Shift Summary (12/10/09)
  • 305885 12/10/09 00:00 miham Beam current lowered for 5uA
  • Wednesday

  • 305884 12/09/09 23:33 Ramesh/miham Q3 off and ON for a test data
  • 305883 12/09/09 23:28 apar Start_Parity_Run_14541__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305882 12/09/09 23:28 ZAFAR AHMED Target screen and strip charts
  • 305877 12/09/09 23:25 apar End_of_Parity_Run_14540__CODA_Config=Pvdis
  • 305876 12/09/09 23:23 ellie Run 2625 was Production, NOT Q2
  • 305875 12/09/09 23:23 apar Start_Parity_Run_14540__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 ramping up during run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305870 12/09/09 23:19 apar End_of_Parity_Run_14539__CODA_Config=Pvdis
  • 305869 12/09/09 22:28 apar Start_Parity_Run_14539__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305866 12/09/09 22:22 miham Beam curret chart during beam recovery procedure
  • 305865 12/09/09 21:56 miham VDC are OFF!
  • 305864 12/09/09 21:52 ellie Q2 and W Plots
  • 305858 12/09/09 21:32 miham Spot++ check for LD2 with (6x6MCC)
  • 305856 12/09/09 21:30 auto_snap Left Arm HV snapshot
  • 305855 12/09/09 21:30 auto_snap Happex Charge Asym snapshot
  • 305854 12/09/09 21:30 auto_snap Hall A Tools snapshot
  • 305853 12/09/09 21:30 auto_snap Compton Polarimeter snapshot
  • 305852 12/09/09 21:30 auto_snap Beam Position and Current snapshot
  • 305851 12/09/09 21:30 auto_snap Accelerator Energies snapshot
  • 305848 12/09/09 21:25 miham Spot++ check with LD2 target (4x4MCC)
  • 305845 12/09/09 21:19 miham VDC are ON!
  • 305844 12/09/09 21:09 miham Spot++ check with BeO Target (4x4MCC)
  • 305841 12/09/09 20:30 M Friend Compton check
  • 305840 12/09/09 20:26 apar End_of_Parity_Run_14538__CODA_Config=Pvdis
  • 305839 12/09/09 19:57 apar Start_Parity_Run_14538__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305838 12/09/09 19:56 apar End_of_Parity_Run_14537__CODA_Config=Pvdis
  • 305837 12/09/09 19:36 miham Still no beam
  • 305836 12/09/09 19:08 apar Start_Parity_Run_14537__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305835 12/09/09 18:59 apar End_of_Parity_Run_14536__CODA_Config=Pvdis
  • 305834 12/09/09 18:50 pking Run plan for tonight
  • 305833 12/09/09 18:29 R. Subedi Power cycling hatsv10
  • 305832 12/09/09 18:03 apar Start_Parity_Run_14536__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305831 12/09/09 17:30 auto_snap Left Arm HV snapshot
  • 305830 12/09/09 17:30 auto_snap Happex Charge Asym snapshot
  • 305829 12/09/09 17:30 auto_snap Hall A Tools snapshot
  • 305828 12/09/09 17:30 auto_snap Compton Polarimeter snapshot
  • 305827 12/09/09 17:30 auto_snap Beam Position and Current snapshot
  • 305826 12/09/09 17:30 auto_snap Accelerator Energies snapshot
  • 305825 12/09/09 17:01 apar Start_Parity_Run_14535__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305824 12/09/09 16:57 apar End_of_Parity_Run_14534__CODA_Config=Pvdis
  • 305823 12/09/09 16:16 R. Michaels Recent Q2
  • 305822 12/09/09 16:07 pking Summary of the Mott measurement of vertical polarization
  • 305821 12/09/09 16:02 apar Start_Parity_Run_14534__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=cosmic Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305819 12/09/09 15:59 S. Sirca Day shift SUMMARY
  • 305818 12/09/09 15:51 R. Michaels electron effic and pion rej (R-HRS)
  • 305817 12/09/09 15:45 santa JP and Santa
  • 305816 12/09/09 15:34 R. Michaels electron effic and pion rej.
  • 305815 12/09/09 14:54 pking At the next opportunity for access, reboot hatsv10
  • 305814 12/09/09 14:39 X. Zheng Right arm rate with Q3 off
  • 305813 12/09/09 14:27 W. Deconinck Portserver hatsv10 not reacting (but still pingable)
  • 305812 12/09/09 14:19 W. Deconinck ROC26 and ROC28 not accessible through port server
  • 305811 12/09/09 14:19 apar Start_Parity_Run_14532__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305810 12/09/09 14:18 apar End_of_Parity_Run_14532__CODA_Config=Pvdis
  • 305807 12/09/09 14:07 S. Sirca Beam OFF for Mott measurement
  • 305806 12/09/09 14:04 S. Sirca Compton HV can not be set
  • 305805 12/09/09 13:58 apar Start_Parity_Run_14531__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305800 12/09/09 13:52 apar End_of_Parity_Run_14530__CODA_Config=Pvdis
  • 305797 12/09/09 13:34 apar Start_Parity_Run_14530__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production Right-Q3 off,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305796 12/09/09 13:29 auto_snap Left Arm HV snapshot
  • 305795 12/09/09 13:29 auto_snap Happex Charge Asym snapshot
  • 305794 12/09/09 13:29 auto_snap Hall A Tools snapshot
  • 305793 12/09/09 13:29 auto_snap Compton Polarimeter snapshot
  • 305792 12/09/09 13:29 auto_snap Beam Position and Current snapshot
  • 305791 12/09/09 13:29 auto_snap Accelerator Energies snapshot
  • 305788 12/09/09 13:14 R. Subedi Spot++ during Q2 measurement
  • 305785 12/09/09 13:09 S. Sirca Turning HRSR Q3 OFF
  • 305782 12/09/09 13:02 S. Sirca Raw rates with 2uA on LD2, 6x6 MCC raster
  • 305779 12/09/09 13:00 pking We will do the "RHRS Q3 off" run after finishing the recovery
  • 305778 12/09/09 12:56 S. Sirca Spot++ check with 2uA on LD2, 4x4 MCC raster
  • 305775 12/09/09 12:52 S. Sirca Raw rates with 2uA on LD2, 4x4 MCC raster
  • 305774 12/09/09 12:47 S. Sirca R_TOP VDCs drawing large currents with no beam
  • 305773 12/09/09 12:37 W. Deconinck Moving to liquid deuterium target position
  • 305772 12/09/09 12:36 S. Sirca Spot++ check with 10uA on BeO, 4x4 MCC raster
  • 305770 12/09/09 12:35 S. Sirca Raw rates with 10uA on BeO, 4x4 MCC raster
  • 305768 12/09/09 12:25 W. Deconinck Moving target to BeO position
  • 305767 12/09/09 12:21 W. Deconinck Moved target to empty position
  • 305766 12/09/09 11:20 brads Integrated charge plot for PVDIS
  • 305765 12/09/09 11:06 brads Run coordinator daily summary
  • 305764 12/09/09 10:58 brads Summary of Compton tuning on Tuesday.
  • 305763 12/09/09 10:56 brads Update: opportunistic Mott after spot move is cancelled (Hall B is back up)
  • 305762 12/09/09 09:29 auto_snap Left Arm HV snapshot
  • 305761 12/09/09 09:29 auto_snap Happex Charge Asym snapshot
  • 305760 12/09/09 09:29 auto_snap Hall A Tools snapshot
  • 305759 12/09/09 09:29 auto_snap Compton Polarimeter snapshot
  • 305758 12/09/09 09:29 auto_snap Beam Position and Current snapshot
  • 305757 12/09/09 09:29 auto_snap Accelerator Energies snapshot
  • 305756 12/09/09 09:26 S. Sirca Flipper output for change to slug 21
  • 305755 12/09/09 09:23 apar Halfwave plate OUT, Slug 21
  • 305753 12/09/09 09:11 brads MCC will do spot move now while SL repair in progress
  • 305752 12/09/09 08:52 S. Sirca Loop 3 sensors are disconnected
  • 305751 12/09/09 08:15 W. Deconinck Target loop 3 cryotemps not reporting
  • 305749 12/09/09 08:00 D.Wang Owl shift summary
  • 305743 12/09/09 07:47 apar End_of_Parity_Run_14529__CODA_Config=Pvdis
  • 305740 12/09/09 07:00 apar Start_Parity_Run_14529__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305737 12/09/09 06:57 apar End_of_Parity_Run_14528__CODA_Config=Pvdis
  • 305734 12/09/09 06:01 apar Start_Parity_Run_14528__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305731 12/09/09 05:56 apar End_of_Parity_Run_14527__CODA_Config=Pvdis
  • 305730 12/09/09 05:29 auto_snap Left Arm HV snapshot
  • 305729 12/09/09 05:29 auto_snap Happex Charge Asym snapshot
  • 305728 12/09/09 05:29 auto_snap Hall A Tools snapshot
  • 305727 12/09/09 05:29 auto_snap Compton Polarimeter snapshot
  • 305726 12/09/09 05:28 auto_snap Beam Position and Current snapshot
  • 305725 12/09/09 05:28 auto_snap Accelerator Energies snapshot
  • 305722 12/09/09 05:01 apar Start_Parity_Run_14527__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305719 12/09/09 04:58 apar End_of_Parity_Run_14526__CODA_Config=Pvdis
  • 305716 12/09/09 04:02 apar Start_Parity_Run_14526__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305713 12/09/09 03:58 apar End_of_Parity_Run_14525__CODA_Config=Pvdis
  • 305712 12/09/09 03:20 D.Wang compton laser power decreased
  • 305709 12/09/09 03:02 apar Start_Parity_Run_14525__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305706 12/09/09 02:58 apar End_of_Parity_Run_14524__CODA_Config=Pvdis
  • 305705 12/09/09 02:13 apar End_of_Parity_Run_14522__CODA_Config=Pvdis
  • 305704 12/09/09 02:12 apar End_of_Parity_Run_14523__CODA_Config=Pvdis
  • 305701 12/09/09 02:02 apar Start_Parity_Run_14524__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305698 12/09/09 01:28 auto_snap Left Arm HV snapshot
  • 305697 12/09/09 01:28 auto_snap Happex Charge Asym snapshot
  • 305696 12/09/09 01:28 auto_snap Hall A Tools snapshot
  • 305695 12/09/09 01:28 auto_snap Compton Polarimeter snapshot
  • 305694 12/09/09 01:28 auto_snap Beam Position and Current snapshot
  • 305693 12/09/09 01:28 auto_snap Accelerator Energies snapshot
  • 305692 12/09/09 01:13 D.Wang compton run number set to21500
  • 305689 12/09/09 01:01 apar Start_Parity_Run_14523__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305686 12/09/09 00:16 reimer Target Screen Shots
  • 305684 12/09/09 00:00 S. Sirca Swing shift SUMMARY
  • Tuesday

  • 305681 12/08/09 23:59 apar Start_Parity_Run_14522__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305680 12/08/09 23:49 miham Q2 and W plots for runs #26088/#5609
  • 305677 12/08/09 23:44 Armstrong Target screens
  • 305676 12/08/09 23:31 S. Sirca Spot++ with 2uA on LD2, 6x6 MCC raster
  • 305673 12/08/09 23:25 S. Sirca Spot++ with 2uA on LD2, 4x4 MCC raster
  • 305670 12/08/09 23:12 Armstrong move to lD2 target
  • 305669 12/08/09 23:10 S. Sirca Spot++ with 10uA on BeO, 4x4 MCC raster
  • 305666 12/08/09 23:01 S. Sirca Results of harp scans
  • 305665 12/08/09 22:58 Armstrong move to BeO target
  • 305664 12/08/09 22:43 Armstrong Which harps to use...
  • 305663 12/08/09 22:34 Glamazdin/Sa Moller results
  • 305662 12/08/09 22:02 brads Backing out of Moller setup -- will have to try again in the morning...
  • 305661 12/08/09 21:28 auto_snap Left Arm HV snapshot
  • 305660 12/08/09 21:28 auto_snap Happex Charge Asym snapshot
  • 305658 12/08/09 21:28 auto_snap Hall A Tools snapshot
  • 305657 12/08/09 21:28 auto_snap Compton Polarimeter snapshot
  • 305656 12/08/09 21:28 auto_snap Beam Position and Current snapshot
  • 305655 12/08/09 21:28 auto_snap Accelerator Energies snapshot
  • 305654 12/08/09 18:49 brads Moller setup started
  • 305653 12/08/09 18:48 brads iocha22 seems to be recovered, but comm is quite slow -- working for now
  • 305652 12/08/09 18:41 M Friend Compton photon collimator moved
  • 305651 12/08/09 17:46 brads iocha22 is accelerator crate, software on call looking at it
  • 305650 12/08/09 17:38 pking Dangling cable in INJ crates was the Qweak trigger cable
  • 305649 12/08/09 17:28 auto_snap Left Arm HV snapshot
  • 305648 12/08/09 17:28 auto_snap Happex Charge Asym snapshot
  • 305647 12/08/09 17:27 auto_snap Hall A Tools snapshot
  • 305646 12/08/09 17:27 auto_snap Compton Polarimeter snapshot
  • 305645 12/08/09 17:27 auto_snap Beam Position and Current snapshot
  • 305644 12/08/09 17:27 auto_snap Accelerator Energies snapshot
  • 305643 12/08/09 17:24 A. Camsonne iocha22 still down - iocha11 back up
  • 305642 12/08/09 17:22 S. Sirca Trouble with IOCHA22
  • 305641 12/08/09 17:18 apar End_of_Parity_Run_14521__CODA_Config=Pvdis
  • 305640 12/08/09 17:17 apar Start_Parity_Run_14521__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305639 12/08/09 16:23 pking Replaced batteries for INJ null asymmetry test
  • 305637 12/08/09 15:58 Kowalski PVDIS
  • 305636 12/08/09 15:39 kowalski PVDIS
  • 305635 12/08/09 15:36 A. Camsonne Compton electron detector moved to position 10 mm
  • 305634 12/08/09 15:35 apar Start_Parity_Run_14520__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305633 12/08/09 15:34 W. Deconinck Target tools screen capture
  • 305632 12/08/09 15:29 apar End_of_Parity_Run_14520__CODA_Config=Pvdis
  • 305631 12/08/09 15:26 brads Compton tuning will end at 4pm
  • 305630 12/08/09 15:13 apar Start_Parity_Run_14519__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305629 12/08/09 15:12 A. Camsonne ETROC power cycled
  • 305628 12/08/09 15:10 apar End_of_Parity_Run_14519__CODA_Config=Pvdis
  • 305627 12/08/09 15:06 apar Start_Parity_Run_14518__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305626 12/08/09 15:04 W. Deconinck Battery swap in injector
  • 305625 12/08/09 15:03 apar Start_Parity_Run_14517__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305624 12/08/09 14:59 apar End_of_Parity_Run_14518__CODA_Config=Pvdis
  • 305623 12/08/09 14:57 W. Deconinck Happex DAQ reboot/kcoda/startcoda
  • 305622 12/08/09 14:55 apar Start_Parity_Run_14516__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305621 12/08/09 14:51 apar Start_Parity_Run_14515__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305620 12/08/09 14:43 apar Start_Parity_Run_14514__CODA_Config=Pvdis_Run_type=Other,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=DAQ test run,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305619 12/08/09 13:59 W. Deconinck Target in empty position
  • 305618 12/08/09 13:59 A. Camsonne hareboot22, iocha22 and iocha11 down - will have to take an access to check the switch
  • 305617 12/08/09 13:48 A. Camsonne Hall A tools before right arm IOC reboot
  • 305616 12/08/09 13:46 A. Camsonne Asked to move to empty target to try new optics tune
  • 305615 12/08/09 13:40 A. Camsonne ETROC power cycled
  • 305614 12/08/09 13:28 A. Camsonne Compton electron detector currents 5 uA position 20 mm
  • 305613 12/08/09 13:27 auto_snap Left Arm HV snapshot
  • 305612 12/08/09 13:27 auto_snap Happex Charge Asym snapshot
  • 305611 12/08/09 13:27 auto_snap Hall A Tools snapshot
  • 305610 12/08/09 13:27 auto_snap Compton Polarimeter snapshot
  • 305609 12/08/09 13:27 auto_snap Beam Position and Current snapshot
  • 305608 12/08/09 13:27 auto_snap Accelerator Energies snapshot
  • 305607 12/08/09 13:19 A. Camsonne Compton electron detector current 5 uA position 25 mm
  • 305606 12/08/09 13:13 A. Camsonne Compton electron detector currents position 25 mm 10 uA
  • 305605 12/08/09 13:10 A. Camsonne Moving Compton electron detector to position 25 mm
  • 305604 12/08/09 13:06 A. Camsonne Compton run singles - position 30 mm 10 uA Run 21465
  • 305603 12/08/09 12:59 A. Camsonne Compton electron detector current at 30 mm 10 uA
  • 305602 12/08/09 12:54 A. Camsonne Compton ETROC power cycled
  • 305599 12/08/09 12:48 apar End_of_Parity_Run_14513__CODA_Config=Pvdis
  • 305598 12/08/09 12:46 A. Camsonne iocha22 diagnostic
  • 305597 12/08/09 12:23 A. Camsonne iocha22 cannot boot
  • 305596 12/08/09 12:11 A. Camsonne iocha22 down
  • 305595 12/08/09 12:10 A. Camsonne Compton electron detector cuurent at position 30 mm 110 uA
  • 305594 12/08/09 11:53 apar Start_Parity_Run_14513__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305591 12/08/09 11:36 W. Deconinck Moving target to deuterium
  • 305590 12/08/09 10:22 R. Subedi Q2 measurement and plots
  • 305589 12/08/09 09:58 brads Run coordinator daily summary
  • 305588 12/08/09 09:48 A. Camsonne Compton tune starts
  • 305587 12/08/09 09:45 W. Deconinck Moving target to empty position
  • 305582 12/08/09 09:32 A. Camsonne Compton EPICS rebooted
  • 305579 12/08/09 09:27 auto_snap Left Arm HV snapshot
  • 305578 12/08/09 09:27 auto_snap Happex Charge Asym snapshot
  • 305577 12/08/09 09:27 auto_snap Hall A Tools snapshot
  • 305576 12/08/09 09:27 auto_snap Compton Polarimeter snapshot
  • 305575 12/08/09 09:27 auto_snap Beam Position and Current snapshot
  • 305574 12/08/09 09:27 auto_snap Accelerator Energies snapshot
  • 305573 12/08/09 09:13 A. Camsonne Compton electron detector slow control down
  • 305568 12/08/09 08:39 apar End_of_Parity_Run_14512__CODA_Config=Pvdis
  • 305564 12/08/09 07:48 D.Wang Owl shift summary
  • 305563 12/08/09 07:33 apar Start_Parity_Run_14512__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305558 12/08/09 07:30 apar End_of_Parity_Run_14511__CODA_Config=Pvdis
  • 305555 12/08/09 06:35 apar Start_Parity_Run_14511__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305550 12/08/09 06:32 apar End_of_Parity_Run_14510__CODA_Config=Pvdis
  • 305549 12/08/09 06:11 reimer Target screens
  • 305548 12/08/09 06:00 D.Wang error in end of run logging for RHRS
  • 305547 12/08/09 05:36 apar Start_Parity_Run_14510__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305542 12/08/09 05:34 apar End_of_Parity_Run_14509__CODA_Config=Pvdis
  • 305541 12/08/09 05:26 auto_snap Left Arm HV snapshot
  • 305540 12/08/09 05:26 auto_snap Happex Charge Asym snapshot
  • 305539 12/08/09 05:26 auto_snap Hall A Tools snapshot
  • 305538 12/08/09 05:26 auto_snap Compton Polarimeter snapshot
  • 305537 12/08/09 05:26 auto_snap Beam Position and Current snapshot
  • 305536 12/08/09 05:26 auto_snap Accelerator Energies snapshot
  • 305535 12/08/09 04:38 apar Start_Parity_Run_14509__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305530 12/08/09 04:36 apar End_of_Parity_Run_14508__CODA_Config=Pvdis
  • 305529 12/08/09 03:40 apar Start_Parity_Run_14508__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305524 12/08/09 03:38 apar End_of_Parity_Run_14507__CODA_Config=Pvdis
  • 305523 12/08/09 03:12 D.Wang Dithering was off, turned back on now
  • 305521 12/08/09 02:42 apar Start_Parity_Run_14507__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305516 12/08/09 02:40 apar End_of_Parity_Run_14506__CODA_Config=Pvdis
  • 305515 12/08/09 01:44 apar Start_Parity_Run_14506__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305510 12/08/09 01:42 apar End_of_Parity_Run_14505__CODA_Config=Pvdis
  • 305509 12/08/09 01:26 auto_snap Left Arm HV snapshot
  • 305508 12/08/09 01:26 auto_snap Happex Charge Asym snapshot
  • 305507 12/08/09 01:26 auto_snap Hall A Tools snapshot
  • 305506 12/08/09 01:26 auto_snap Compton Polarimeter snapshot
  • 305505 12/08/09 01:26 auto_snap Beam Position and Current snapshot
  • 305504 12/08/09 01:26 auto_snap Accelerator Energies snapshot
  • 305503 12/08/09 01:08 D.Wang compton photon detector base current gets better
  • 305502 12/08/09 00:57 D.Wang Lumi HV adjusted
  • 305501 12/08/09 00:46 apar Start_Parity_Run_14505__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305496 12/08/09 00:44 apar End_of_Parity_Run_14504__CODA_Config=Pvdis
  • 305495 12/08/09 00:02 S. Sirca Swing shift SUMMARY
  • Monday

  • 305493 12/07/09 23:48 apar Start_Parity_Run_14504__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305488 12/07/09 23:46 apar End_of_Parity_Run_14503__CODA_Config=Pvdis
  • 305487 12/07/09 23:40 ZAFAR AHMED Target
  • 305486 12/07/09 23:35 S. Sirca spot++ check
  • 305485 12/07/09 23:34 S. Sirca Hall A general tools & beam stripchart
  • 305484 12/07/09 23:11 S. Sirca Running at 115uA?
  • 305481 12/07/09 22:50 apar Start_Parity_Run_14503__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305478 12/07/09 22:47 apar End_of_Parity_Run_14502__CODA_Config=Pvdis
  • 305477 12/07/09 21:52 apar Start_Parity_Run_14502__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305475 12/07/09 21:43 apar Start_Parity_Run_14501__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305472 12/07/09 21:36 apar Start_Parity_Run_14500__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305467 12/07/09 21:34 apar End_of_Parity_Run_14499__CODA_Config=Pvdis
  • 305466 12/07/09 21:26 auto_snap Left Arm HV snapshot
  • 305465 12/07/09 21:26 auto_snap Happex Charge Asym snapshot
  • 305464 12/07/09 21:26 auto_snap Hall A Tools snapshot
  • 305463 12/07/09 21:25 auto_snap Compton Polarimeter snapshot
  • 305462 12/07/09 21:25 auto_snap Beam Position and Current snapshot
  • 305461 12/07/09 21:25 auto_snap Accelerator Energies snapshot
  • 305460 12/07/09 20:38 apar Start_Parity_Run_14499__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305455 12/07/09 20:36 apar End_of_Parity_Run_14498__CODA_Config=Pvdis
  • 305454 12/07/09 19:40 apar Start_Parity_Run_14498__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305449 12/07/09 19:38 apar End_of_Parity_Run_14497__CODA_Config=Pvdis
  • 305448 12/07/09 18:42 apar Start_Parity_Run_14497__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305443 12/07/09 18:40 apar End_of_Parity_Run_14496__CODA_Config=Pvdis
  • 305442 12/07/09 17:45 apar Start_Parity_Run_14496__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305439 12/07/09 17:42 apar End_of_Parity_Run_14495__CODA_Config=Pvdis
  • 305436 12/07/09 17:41 R. Subedi List of runs during deadtime measurement with a 10.26 kHz tagger
  • 305435 12/07/09 17:36 R. Subedi Beam chart during the deadtime test
  • 305434 12/07/09 17:25 auto_snap Left Arm HV snapshot
  • 305433 12/07/09 17:25 auto_snap Happex Charge Asym snapshot
  • 305432 12/07/09 17:25 auto_snap Hall A Tools snapshot
  • 305431 12/07/09 17:25 auto_snap Compton Polarimeter snapshot
  • 305430 12/07/09 17:25 auto_snap Beam Position and Current snapshot
  • 305429 12/07/09 17:25 auto_snap Accelerator Energies snapshot
  • 305428 12/07/09 17:07 apar Start_Parity_Run_14495__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 20-108 uA (ramping current) (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305425 12/07/09 17:06 R. Subedi Tagger (10.26 kHz) is disabled
  • 305421 12/07/09 16:54 R. Subedi Right scaler @ 20 uA with a 10.26 kHz tagger
  • 305420 12/07/09 16:54 R. Subedi Left scaler @ 20 uA with a 10.26 kHz tagger
  • 305419 12/07/09 16:53 apar Start_Parity_Run_14494__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 20 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305416 12/07/09 16:51 apar End_of_Parity_Run_14493__CODA_Config=Pvdis
  • 305413 12/07/09 16:39 R. Subedi Right scaler @ 40 uA with a 10.26 kHz tagger
  • 305412 12/07/09 16:38 R. Subedi Left scaler @ 40 uA with a 10.26 kHz tagger
  • 305411 12/07/09 16:37 apar Start_Parity_Run_14493__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 40 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305406 12/07/09 16:35 apar End_of_Parity_Run_14492__CODA_Config=Pvdis
  • 305405 12/07/09 16:33 M Friend Compton photon detector
  • 305404 12/07/09 16:26 R. Subedi Right scaler @ 60 uA with a 10.26 kHz tagger
  • 305403 12/07/09 16:25 R. Subedi Left scaler @ 60 uA with a 10.26 kHz tagger
  • 305402 12/07/09 16:20 apar Start_Parity_Run_14492__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 60 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305399 12/07/09 16:19 M Friend Compton photon detector
  • 305396 12/07/09 16:17 apar End_of_Parity_Run_14491__CODA_Config=Pvdis
  • 305395 12/07/09 16:10 brads Run coordinator daily summary
  • 305393 12/07/09 16:06 R. Subedi Right scaler @80 uA with 10.26 kHz tagger
  • 305392 12/07/09 16:06 R. Subedi Left scaler @80 uA with 10.26 kHz tagger
  • 305391 12/07/09 16:03 Kowalski day shift summary
  • 305390 12/07/09 16:00 apar Start_Parity_Run_14491__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 80 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305387 12/07/09 15:57 apar End_of_Parity_Run_14490__CODA_Config=Pvdis
  • 305384 12/07/09 15:47 R. Subedi Right scaler @ 108 uA
  • 305383 12/07/09 15:46 R. Subedi Left scaler @ 108 uA
  • 305382 12/07/09 15:43 apar Start_Parity_Run_14490__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 108 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305379 12/07/09 15:42 apar End_of_Parity_Run_14489__CODA_Config=Pvdis
  • 305376 12/07/09 15:39 apar Start_Parity_Run_14489__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 108 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305373 12/07/09 15:38 apar End_of_Parity_Run_14488__CODA_Config=Pvdis
  • 305370 12/07/09 15:33 R. Subedi 10.26 kHz tagger enabled
  • 305369 12/07/09 15:32 apar Start_Parity_Run_14488__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Deadtime measurement at 108 uA (with 10.26 kHz tagger),dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305367 12/07/09 15:30 W. Deconinck Target tools screen capture
  • 305363 12/07/09 15:26 apar End_of_Parity_Run_14487__CODA_Config=Pvdis
  • 305362 12/07/09 15:04 Lamiaa Dithering is working again starting from run# 14486
  • 305361 12/07/09 14:51 Kowalski PVDIS
  • 305360 12/07/09 14:46 R. Michaels dithering working again
  • 305359 12/07/09 14:44 R. Michaels re: ET instabilities
  • 305358 12/07/09 14:41 Lamiaa Dithering is working fine for run# 14487
  • 305357 12/07/09 14:30 apar Start_Parity_Run_14487__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305356 12/07/09 14:28 apar End_of_Parity_Run_14486__CODA_Config=Pvdis
  • 305355 12/07/09 14:20 apar Start_Parity_Run_14486__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305352 12/07/09 14:11 A. Camsonne IOC remote power switch power cycled
  • 305351 12/07/09 13:43 R. Michaels runbird and panguin fixed, dithering still not
  • 305350 12/07/09 13:27 apar End_of_Parity_Run_14485__CODA_Config=Pvdis
  • 305349 12/07/09 13:25 apar Start_Parity_Run_14485__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305345 12/07/09 13:25 auto_snap Accelerator Energies snapshot
  • 305344 12/07/09 13:21 apar End_of_Parity_Run_14484__CODA_Config=Pvdis
  • 305343 12/07/09 13:20 Lamiaa Faulty Entry
  • 305342 12/07/09 13:18 apar Start_Parity_Run_14484__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305341 12/07/09 13:15 apar End_of_Parity_Run_14483__CODA_Config=Pvdis
  • 305340 12/07/09 13:05 Lamiaa Beam modulation (Dithering) is ON for run # 14483
  • 305339 12/07/09 12:48 apar Start_Parity_Run_14483__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305336 12/07/09 12:43 apar End_of_Parity_Run_14482__CODA_Config=Pvdis
  • 305333 12/07/09 12:40 Lamiaa Beam modulation is stuck @ OFF state
  • 305332 12/07/09 12:02 D.Wang The failure of runbird and panguin is probably due to epics reboot
  • 305331 12/07/09 11:47 apar Start_Parity_Run_14482__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305330 12/07/09 11:45 apar End_of_Parity_Run_14481__CODA_Config=Pvdis
  • 305326 12/07/09 11:16 Lamiaa Runbird is frozen!
  • 305325 12/07/09 10:50 apar Start_Parity_Run_14481__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305322 12/07/09 10:45 A. Camsonne Compton remote controlled power strip down
  • 305321 12/07/09 10:44 apar End_of_Parity_Run_14480__CODA_Config=Pvdis
  • 305318 12/07/09 10:09 Lamiaa Spot++ Check
  • 305317 12/07/09 09:52 Lamiaa Q^2 & W distribution of Q2scan run
  • 305316 12/07/09 09:52 A. Camsonne Compton EPICS reboot
  • 305315 12/07/09 09:52 apar Start_Parity_Run_14480__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305312 12/07/09 09:41 apar End_of_Parity_Run_14479__CODA_Config=Pvdis
  • 305309 12/07/09 09:35 apar Start_Parity_Run_14479__CODA_Config=Pvdis_Run_type=Q2scan,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305308 12/07/09 09:32 apar End_of_Parity_Run_14478__CODA_Config=Pvdis
  • 305307 12/07/09 09:32 W. Deconinck Spot++
  • 305298 12/07/09 09:25 apar Start_Parity_Run_14478__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305297 12/07/09 09:25 auto_snap Left Arm HV snapshot
  • 305296 12/07/09 09:25 auto_snap Happex Charge Asym snapshot
  • 305295 12/07/09 09:25 auto_snap Hall A Tools snapshot
  • 305294 12/07/09 09:25 auto_snap Compton Polarimeter snapshot
  • 305293 12/07/09 09:24 auto_snap Beam Position and Current snapshot
  • 305292 12/07/09 09:24 auto_snap Accelerator Energies snapshot
  • 305289 12/07/09 09:08 Deng Dithering is dead for runs 14475-14477
  • 305288 12/07/09 08:57 Deng Beam modulation (dithering) dead, reboot roc 23 and it came back
  • 305287 12/07/09 08:52 apar End_of_Parity_Run_14477__CODA_Config=Pvdis
  • 305284 12/07/09 08:24 apar Start_Parity_Run_14477__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305283 12/07/09 08:22 apar End_of_Parity_Run_14476__CODA_Config=Pvdis
  • 305278 12/07/09 08:09 Al Tobias Owl Shift Summary
  • 305276 12/07/09 07:26 apar Start_Parity_Run_14476__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305275 12/07/09 07:24 Al Tobias Compton Laser ON or OFF?
  • 305274 12/07/09 07:24 apar End_of_Parity_Run_14475__CODA_Config=Pvdis
  • 305269 12/07/09 07:16 Al Tobias Compton Cavity Power & Photon Rates Flat Line
  • 305268 12/07/09 06:34 Deng screen shot of HRS scaler rates in production mode (108uA)
  • 305267 12/07/09 06:31 Deng screen shot of happex feedback monitor
  • 305266 12/07/09 06:29 Al Tobias Alarm Handler: LeCroy HV Beamline Slot 11 Channel 7 tripped off - now OK
  • 305265 12/07/09 06:29 Deng half wave plate IN now, slug number is 20
  • 305264 12/07/09 06:28 Al Tobias Compton DAQ ROC2 not responding - Reboot
  • 305263 12/07/09 06:26 apar Start_Parity_Run_14475__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305260 12/07/09 06:24 apar Halfwave plate IN, Slug 20
  • 305258 12/07/09 06:22 apar End_of_Parity_Run_14474__CODA_Config=Pvdis
  • 305257 12/07/09 06:10 V BELLINI Target Stripcharts
  • 305256 12/07/09 05:26 apar Start_Parity_Run_14474__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305251 12/07/09 05:24 auto_snap Left Arm HV snapshot
  • 305250 12/07/09 05:24 auto_snap Happex Charge Asym snapshot
  • 305248 12/07/09 05:24 apar End_of_Parity_Run_14473__CODA_Config=Pvdis
  • 305249 12/07/09 05:24 auto_snap Hall A Tools snapshot
  • 305247 12/07/09 05:24 auto_snap Compton Polarimeter snapshot
  • 305246 12/07/09 05:24 auto_snap Beam Position and Current snapshot
  • 305245 12/07/09 05:24 auto_snap Accelerator Energies snapshot
  • 305244 12/07/09 05:24 Al Tobias Hall A General Tools & BCM Snapshots
  • 305243 12/07/09 04:58 Al Tobias Spot++ on Current run 14473/26044/5568
  • 305242 12/07/09 04:29 apar Start_Parity_Run_14473__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305237 12/07/09 04:26 apar End_of_Parity_Run_14472__CODA_Config=Pvdis
  • 305236 12/07/09 04:22 Al Tobias Compton DAQ ROC1 not responding for about a minute #21434 - Reboot
  • 305235 12/07/09 03:56 Al Tobias Compton ROC2 not responding for 50mins run 21433
  • 305234 12/07/09 03:31 apar Start_Parity_Run_14472__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305229 12/07/09 03:29 apar End_of_Parity_Run_14471__CODA_Config=Pvdis
  • 305228 12/07/09 02:35 Al Tobias Reboot Happex DAQ w/ run 14470
  • 305227 12/07/09 02:33 apar Start_Parity_Run_14471__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305224 12/07/09 02:27 apar Start_Parity_Run_14470__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305221 12/07/09 02:24 apar End_of_Parity_Run_14469__CODA_Config=Pvdis
  • 305220 12/07/09 02:10 Al Tobias Compton DAQ Reboot
  • 305219 12/07/09 01:52 V BELLINI Target stripcharts
  • 305218 12/07/09 01:29 apar Start_Parity_Run_14469__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305213 12/07/09 01:26 apar End_of_Parity_Run_14468__CODA_Config=Pvdis
  • 305212 12/07/09 01:24 auto_snap Left Arm HV snapshot
  • 305211 12/07/09 01:24 auto_snap Happex Charge Asym snapshot
  • 305210 12/07/09 01:24 auto_snap Hall A Tools snapshot
  • 305209 12/07/09 01:24 auto_snap Compton Polarimeter snapshot
  • 305208 12/07/09 01:24 auto_snap Beam Position and Current snapshot
  • 305207 12/07/09 01:24 auto_snap Accelerator Energies snapshot
  • 305206 12/07/09 00:48 Al Tobias Compton DAQ Reboot: Run 21431 ROC5 was not responding
  • 305205 12/07/09 00:31 apar Start_Parity_Run_14468__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305200 12/07/09 00:28 apar End_of_Parity_Run_14467__CODA_Config=Pvdis
  • 305198 12/07/09 00:00 S. Sirca Swing shift SUMMARY
  • Sunday

  • 305197 12/06/09 23:55 Abdurahim Target screens
  • 305196 12/06/09 23:32 apar Start_Parity_Run_14467__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305191 12/06/09 23:30 apar End_of_Parity_Run_14466__CODA_Config=Pvdis
  • 305188 12/06/09 22:34 apar Start_Parity_Run_14466__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305185 12/06/09 22:32 Kai slug18 : global electron L=R, global pion L=R(within error bar)
  • 305184 12/06/09 22:24 apar End_of_Parity_Run_14465__CODA_Config=Pvdis
  • 305183 12/06/09 21:29 S. Sirca Hall A general tools
  • 305182 12/06/09 21:28 apar Start_Parity_Run_14465__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305177 12/06/09 21:26 apar End_of_Parity_Run_14464__CODA_Config=Pvdis
  • 305176 12/06/09 21:26 Abdurahim Target screens
  • 305175 12/06/09 21:24 auto_snap Left Arm HV snapshot
  • 305174 12/06/09 21:24 auto_snap Happex Charge Asym snapshot
  • 305173 12/06/09 21:23 auto_snap Hall A Tools snapshot
  • 305172 12/06/09 21:23 auto_snap Compton Polarimeter snapshot
  • 305171 12/06/09 21:23 auto_snap Beam Position and Current snapshot
  • 305170 12/06/09 21:23 auto_snap Accelerator Energies snapshot
  • 305169 12/06/09 21:01 Kai Please check the Lecroy offset of TS gr6 on the right arm when available
  • 305168 12/06/09 20:31 apar Start_Parity_Run_14464__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305164 12/06/09 20:28 apar End_of_Parity_Run_14463__CODA_Config=Pvdis
  • 305163 12/06/09 19:32 apar Start_Parity_Run_14463__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305158 12/06/09 19:30 apar End_of_Parity_Run_14462__CODA_Config=Pvdis
  • 305157 12/06/09 18:35 apar Start_Parity_Run_14462__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305154 12/06/09 18:25 miham Q2 and W plots for runs #26033/#5557
  • 305153 12/06/09 18:21 S. Sirca Spot++ during low-current run 26033
  • 305146 12/06/09 18:04 apar End_of_Parity_Run_14461__CODA_Config=Pvdis
  • 305145 12/06/09 17:58 Abdurahim Target Screens
  • 305144 12/06/09 17:54 A. Camsonne Compton evenSelection modified
  • 305143 12/06/09 17:41 R. Subedi Golden run for RHRS changed to 5555
  • 305142 12/06/09 17:23 auto_snap Left Arm HV snapshot
  • 305141 12/06/09 17:23 auto_snap Happex Charge Asym snapshot
  • 305140 12/06/09 17:23 auto_snap Hall A Tools snapshot
  • 305139 12/06/09 17:23 auto_snap Compton Polarimeter snapshot
  • 305138 12/06/09 17:23 auto_snap Beam Position and Current snapshot
  • 305137 12/06/09 17:23 auto_snap Accelerator Energies snapshot
  • 305136 12/06/09 17:08 apar Start_Parity_Run_14461__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305131 12/06/09 17:06 apar End_of_Parity_Run_14460__CODA_Config=Pvdis
  • 305130 12/06/09 16:27 miham Trigger T8 missing for HRS-R run
  • 305127 12/06/09 16:10 apar Start_Parity_Run_14460__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305124 12/06/09 16:07 apar End_of_Parity_Run_14459__CODA_Config=Pvdis
  • 305122 12/06/09 15:59 Kowalski PVDIS
  • 305121 12/06/09 15:51 Kowalski PVDIS
  • 305118 12/06/09 15:11 apar Start_Parity_Run_14459__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305115 12/06/09 15:10 apar End_of_Parity_Run_14458__CODA_Config=Pvdis
  • 305114 12/06/09 14:16 jpchen tech: alarm on a valve controller
  • 305113 12/06/09 14:14 apar Start_Parity_Run_14458__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305110 12/06/09 14:10 jpchen D2 tank/loop pressure raised by ~ 2psi
  • 305109 12/06/09 14:06 brads Run coordinator daily summary
  • 305108 12/06/09 14:06 jpchen Temperature readback comm problem fixed
  • 305107 12/06/09 13:25 kbrider target temp control not working
  • 305106 12/06/09 13:23 auto_snap Left Arm HV snapshot
  • 305105 12/06/09 13:23 auto_snap Happex Charge Asym snapshot
  • 305104 12/06/09 13:23 auto_snap Hall A Tools snapshot
  • 305103 12/06/09 13:23 auto_snap Compton Polarimeter snapshot
  • 305102 12/06/09 13:23 auto_snap Beam Position and Current snapshot
  • 305101 12/06/09 13:22 auto_snap Accelerator Energies snapshot
  • 305100 12/06/09 13:14 Kai expert check report
  • 305097 12/06/09 13:11 apar End_of_Parity_Run_14457__CODA_Config=Pvdis
  • 305096 12/06/09 13:07 Kai bring CH synch back after reboot TS1
  • 305095 12/06/09 13:06 apar Start_Parity_Run_14457__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305090 12/06/09 13:01 apar End_of_Parity_Run_14455__CODA_Config=Pvdis
  • 305089 12/06/09 12:47 R. Subedi Re: target temp sensor not responding
  • 305088 12/06/09 12:42 Kai possible glitch in CH synchronization scaler
  • 305087 12/06/09 12:41 Kai J.P. comes in to fix the target problem
  • 305086 12/06/09 12:08 Kai nearly all of the leadglass blocks get back
  • 305085 12/06/09 12:05 apar Start_Parity_Run_14455__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305084 12/06/09 12:04 kbrider target temp sensor not responding
  • 305078 12/06/09 11:32 R. Subedi Re: expert shift check: is it normal for HRS online gui
  • 305077 12/06/09 11:08 Kai dithering back
  • 305076 12/06/09 11:06 A. Camsonne Compton VNC
  • 305075 12/06/09 11:00 apar Start_Parity_Run_14454__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305070 12/06/09 10:53 apar End_of_Parity_Run_14452__CODA_Config=Pvdis
  • 305069 12/06/09 10:45 A. Camsonne Dithering Kill switch stuck to on
  • 305068 12/06/09 10:36 A. Camsonne Compton IOC rebooted
  • 305067 12/06/09 10:14 A. Camsonne Compton EPICS IOC crashed
  • 305066 12/06/09 10:00 apar Start_Parity_Run_14452__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305065 12/06/09 09:58 apar End_of_Parity_Run_14451__CODA_Config=Pvdis
  • 305060 12/06/09 09:22 auto_snap Left Arm HV snapshot
  • 305059 12/06/09 09:22 auto_snap Happex Charge Asym snapshot
  • 305058 12/06/09 09:22 auto_snap Hall A Tools snapshot
  • 305057 12/06/09 09:22 auto_snap Compton Polarimeter snapshot
  • 305056 12/06/09 09:22 auto_snap Beam Position and Current snapshot
  • 305055 12/06/09 09:22 auto_snap Accelerator Energies snapshot
  • 305054 12/06/09 09:01 apar Start_Parity_Run_14451__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305049 12/06/09 08:59 apar End_of_Parity_Run_14450__CODA_Config=Pvdis
  • 305048 12/06/09 08:46 Lamiaa Spot++ Check
  • 305047 12/06/09 08:41 kbrider target screens
  • 305045 12/06/09 08:21 Al Tobias Owl Shift Summary
  • 305044 12/06/09 08:14 Al Tobias Alarm Handler went off: Lecroy High Voltage, Slot 11, channel 1
  • 305043 12/06/09 08:09 Deng updated golden runs for more files
  • 305040 12/06/09 08:03 apar Start_Parity_Run_14450__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305037 12/06/09 08:01 apar End_of_Parity_Run_14449__CODA_Config=Pvdis
  • 305036 12/06/09 07:05 apar Start_Parity_Run_14449__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305031 12/06/09 07:02 apar End_of_Parity_Run_14448__CODA_Config=Pvdis
  • 305030 12/06/09 06:16 V Bellini Target stripcharts
  • 305029 12/06/09 06:15 Al Tobias Hall A General Tools Screen Capture
  • 305028 12/06/09 06:07 apar Start_Parity_Run_14448__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305023 12/06/09 06:04 apar End_of_Parity_Run_14447__CODA_Config=Pvdis
  • 305022 12/06/09 05:47 Deng expert shift check: is it normal for HRS online gui
  • 305021 12/06/09 05:37 A. Tobias More beam trips past 25min at 108uA
  • 305020 12/06/09 05:22 auto_snap Left Arm HV snapshot
  • 305019 12/06/09 05:22 auto_snap Happex Charge Asym snapshot
  • 305018 12/06/09 05:22 auto_snap Hall A Tools snapshot
  • 305017 12/06/09 05:22 auto_snap Compton Polarimeter snapshot
  • 305016 12/06/09 05:22 auto_snap Beam Position and Current snapshot
  • 305015 12/06/09 05:22 auto_snap Accelerator Energies snapshot
  • 305014 12/06/09 05:21 Al Tobias Happex & Compton DAQs rebooted in past hour
  • 305013 12/06/09 05:08 apar Start_Parity_Run_14447__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305008 12/06/09 05:05 apar End_of_Parity_Run_14446__CODA_Config=Pvdis
  • 305003 12/06/09 04:10 apar Start_Parity_Run_14446__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 305002 12/06/09 03:59 apar Start_Parity_Run_14445__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304997 12/06/09 03:56 apar End_of_Parity_Run_14444__CODA_Config=Pvdis
  • 304996 12/06/09 03:55 Al Tobias spot++ check run# 26016
  • 304995 12/06/09 03:54 Deng Global electron/pion rate during production
  • 304994 12/06/09 03:02 Al Tobias spot++ at 2uA and 4x4mm raster
  • 304993 12/06/09 03:01 apar Start_Parity_Run_14444__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304990 12/06/09 02:56 Al Tobias T1 rates at 2uA 4x4mm raster
  • 304985 12/06/09 02:41 V BELLINI Moved back to LD2 target
  • 304984 12/06/09 01:50 Al Tobias Lower Pressure due to Environment Temperature Lower
  • 304983 12/06/09 01:42 Vince BELLIN Moving to empty target
  • 304980 12/06/09 01:37 apar End_of_Parity_Run_14443__CODA_Config=Pvdis
  • 304979 12/06/09 01:35 Al Tobias MCC beam down again: 20+ minutes
  • 304978 12/06/09 01:25 Deng half wave plate changed to be OUT, Slug=19
  • 304977 12/06/09 01:22 Al Tobias Beam back at 108uA & IHWP out
  • 304976 12/06/09 01:22 auto_snap Left Arm HV snapshot
  • 304975 12/06/09 01:22 auto_snap Happex Charge Asym snapshot
  • 304974 12/06/09 01:22 auto_snap Hall A Tools snapshot
  • 304973 12/06/09 01:21 auto_snap Compton Polarimeter snapshot
  • 304972 12/06/09 01:21 auto_snap Beam Position and Current snapshot
  • 304971 12/06/09 01:21 auto_snap Accelerator Energies snapshot
  • 304970 12/06/09 01:21 apar Start_Parity_Run_14443__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304967 12/06/09 01:17 apar Halfwave plate OUT, Slug 19
  • 304964 12/06/09 01:09 apar End_of_Parity_Run_14442__CODA_Config=Pvdis
  • 304963 12/06/09 00:44 Al Tobias MCC needs to reboot an RF zone crate, will be just a few minutes
  • 304962 12/06/09 00:26 Vincenzo BEL Lower target pressure
  • 304961 12/06/09 00:21 Al Tobias MCC trying to deliver 115uA beam: Request back to 108uA.
  • 304960 12/06/09 00:17 apar Start_Parity_Run_14442__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304955 12/06/09 00:14 apar End_of_Parity_Run_14441__CODA_Config=Pvdis
  • 304953 12/06/09 00:00 S. Sirca Swing shift SUMMARY
  • Saturday

  • 304952 12/05/09 23:23 S. Sirca Hall A general tools
  • 304950 12/05/09 23:18 apar Start_Parity_Run_14441__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304948 12/05/09 23:12 apar End_of_Parity_Run_14440__CODA_Config=Pvdis
  • 304946 12/05/09 23:01 Deng golden run updated
  • 304945 12/05/09 22:31 apar End_of_Parity_Run_14439__CODA_Config=Pvdis
  • 304944 12/05/09 22:28 apar Start_Parity_Run_14440__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304939 12/05/09 21:29 apar Start_Parity_Run_14439__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304934 12/05/09 21:27 apar End_of_Parity_Run_14438__CODA_Config=Pvdis
  • 304933 12/05/09 21:21 auto_snap Left Arm HV snapshot
  • 304932 12/05/09 21:21 auto_snap Happex Charge Asym snapshot
  • 304931 12/05/09 21:21 auto_snap Hall A Tools snapshot
  • 304930 12/05/09 21:21 auto_snap Compton Polarimeter snapshot
  • 304929 12/05/09 21:21 auto_snap Beam Position and Current snapshot
  • 304928 12/05/09 21:21 auto_snap Accelerator Energies snapshot
  • 304927 12/05/09 21:11 Abdurahim Target screens
  • 304926 12/05/09 20:31 apar Start_Parity_Run_14438__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304921 12/05/09 20:29 apar End_of_Parity_Run_14437__CODA_Config=Pvdis
  • 304920 12/05/09 20:16 R. Michaels disk cleanup
  • 304917 12/05/09 19:33 apar Start_Parity_Run_14437__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304914 12/05/09 19:30 apar End_of_Parity_Run_14436__CODA_Config=Pvdis
  • 304913 12/05/09 19:12 X. Zheng more on left arm efficiency (threshold/cut change)
  • 304912 12/05/09 18:58 X. Zheng Left arm electron efficiency now
  • 304911 12/05/09 18:55 Abdurahim Target Alarm
  • 304908 12/05/09 18:34 X. Zheng Left gr5 timing after adding a 4ns delay to TS
  • 304907 12/05/09 18:34 apar Start_Parity_Run_14436__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304906 12/05/09 18:29 miham Q2 and W plots for runs #26007/#5531
  • 304905 12/05/09 18:24 X. Zheng PRL2 block [31] in gr6 is now back
  • 304902 12/05/09 18:14 S. Sirca Raw rates with 2uA on 20cm LD2, 6x6 MCC raster
  • 304899 12/05/09 18:09 S. Sirca Spot++ check after controlled access
  • 304898 12/05/09 18:08 Ramesh/Xiaoc Access work
  • 304895 12/05/09 17:21 auto_snap Left Arm HV snapshot
  • 304894 12/05/09 17:21 auto_snap Happex Charge Asym snapshot
  • 304893 12/05/09 17:21 auto_snap Hall A Tools snapshot
  • 304892 12/05/09 17:21 auto_snap Compton Polarimeter snapshot
  • 304891 12/05/09 17:21 auto_snap Beam Position and Current snapshot
  • 304890 12/05/09 17:21 auto_snap Accelerator Energies snapshot
  • 304889 12/05/09 16:42 M Friend Compton diagnostic scintillators moved
  • 304888 12/05/09 16:39 brads Run coordinator daily summary
  • 304886 12/05/09 16:01 A. Deur Day shift summary
  • 304883 12/05/09 15:54 apar End_of_Parity_Run_14435__CODA_Config=Pvdis
  • 304882 12/05/09 15:04 apar Start_Parity_Run_14435__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304877 12/05/09 15:02 apar End_of_Parity_Run_14434__CODA_Config=Pvdis
  • 304876 12/05/09 14:36 kbrider target screen
  • 304875 12/05/09 14:06 apar Start_Parity_Run_14434__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304870 12/05/09 14:04 apar End_of_Parity_Run_14433__CODA_Config=Pvdis
  • 304869 12/05/09 13:20 auto_snap Left Arm HV snapshot
  • 304868 12/05/09 13:20 auto_snap Happex Charge Asym snapshot
  • 304867 12/05/09 13:20 auto_snap Hall A Tools snapshot
  • 304866 12/05/09 13:20 auto_snap Compton Polarimeter snapshot
  • 304865 12/05/09 13:20 auto_snap Beam Position and Current snapshot
  • 304864 12/05/09 13:20 auto_snap Accelerator Energies snapshot
  • 304863 12/05/09 13:09 apar Start_Parity_Run_14433__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304858 12/05/09 13:07 apar End_of_Parity_Run_14432__CODA_Config=Pvdis
  • 304857 12/05/09 12:11 apar Start_Parity_Run_14432__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304856 12/05/09 12:05 apar Start_Parity_Run_14431__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304851 12/05/09 12:03 apar End_of_Parity_Run_14430__CODA_Config=Pvdis
  • 304850 12/05/09 11:40 Kai double peak on group5,6 TDC disappear and some blocks get back from trx:try
  • 304849 12/05/09 11:07 apar Start_Parity_Run_14430__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304844 12/05/09 11:05 apar End_of_Parity_Run_14429__CODA_Config=Pvdis
  • 304843 12/05/09 10:57 Kai software simulation on the ele efficiency using more accurate cut
  • 304842 12/05/09 10:41 R. Subedi PVDIS scaler displays @ 98 uA (all 8 scalers in left and right pvdis setups)
  • 304841 12/05/09 10:26 Kai current PID on right arm
  • 304840 12/05/09 10:19 Kai current PID on the left (real case)
  • 304839 12/05/09 10:10 apar Start_Parity_Run_14429__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304834 12/05/09 10:07 apar End_of_Parity_Run_14428__CODA_Config=Pvdis
  • 304833 12/05/09 09:20 auto_snap Left Arm HV snapshot
  • 304832 12/05/09 09:20 auto_snap Happex Charge Asym snapshot
  • 304831 12/05/09 09:20 auto_snap Hall A Tools snapshot
  • 304830 12/05/09 09:20 auto_snap Compton Polarimeter snapshot
  • 304829 12/05/09 09:20 auto_snap Beam Position and Current snapshot
  • 304828 12/05/09 09:20 auto_snap Accelerator Energies snapshot
  • 304827 12/05/09 09:12 A. Deur BCM/BPM strip chart
  • 304826 12/05/09 09:12 apar Start_Parity_Run_14428__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304821 12/05/09 09:10 apar End_of_Parity_Run_14427__CODA_Config=Pvdis
  • 304820 12/05/09 09:05 A. Deur Hall A tools snapshot
  • 304819 12/05/09 09:04 Lamiaa Spot++ Check
  • 304818 12/05/09 09:01 kbrider target screens
  • 304817 12/05/09 08:13 apar Start_Parity_Run_14427__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304812 12/05/09 08:11 apar End_of_Parity_Run_14426__CODA_Config=Pvdis
  • 304810 12/05/09 08:06 Al Tobias Owl Shift Summary
  • 304809 12/05/09 07:36 Al Tobias spot++ 4x4mm raster w/ 100uA on LD2
  • 304808 12/05/09 07:15 apar Start_Parity_Run_14426__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304803 12/05/09 07:12 apar End_of_Parity_Run_14425__CODA_Config=Pvdis
  • 304802 12/05/09 06:49 Al Tobias Hall A General Tools Snapshot
  • 304801 12/05/09 06:16 apar Start_Parity_Run_14425__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304797 12/05/09 06:10 apar End_of_Parity_Run_14424__CODA_Config=Pvdis
  • 304796 12/05/09 05:50 Vincenzo BEL Target Stripcharts
  • 304795 12/05/09 05:20 auto_snap Left Arm HV snapshot
  • 304794 12/05/09 05:20 auto_snap Happex Charge Asym snapshot
  • 304793 12/05/09 05:20 auto_snap Hall A Tools snapshot
  • 304792 12/05/09 05:19 auto_snap Compton Polarimeter snapshot
  • 304791 12/05/09 05:19 auto_snap Beam Position and Current snapshot
  • 304790 12/05/09 05:19 auto_snap Accelerator Energies snapshot
  • 304789 12/05/09 05:14 apar Start_Parity_Run_14424__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304784 12/05/09 05:11 apar End_of_Parity_Run_14423__CODA_Config=Pvdis
  • 304783 12/05/09 04:21 Al Tobias Compton restart issue - now OK
  • 304782 12/05/09 04:15 apar Start_Parity_Run_14423__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304777 12/05/09 04:13 apar End_of_Parity_Run_14422__CODA_Config=Pvdis
  • 304776 12/05/09 03:31 D.Wang Lumi HV adjusted
  • 304775 12/05/09 03:17 apar Start_Parity_Run_14422__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304770 12/05/09 03:14 apar End_of_Parity_Run_14421__CODA_Config=Pvdis
  • 304769 12/05/09 03:10 Al Tobias Happex Run# 14419 - first 10min Beam Modulation was OFF
  • 304768 12/05/09 03:03 Al Tobias Happex DAQ not counting at start OWL
  • 304767 12/05/09 03:02 D.Wang scaler rate zero is due to threshold
  • 304766 12/05/09 02:46 Al Tobias T3 and T4 scalars Zero - Is this OK?
  • 304765 12/05/09 02:18 apar Start_Parity_Run_14421__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304760 12/05/09 02:15 apar End_of_Parity_Run_14420__CODA_Config=Pvdis
  • 304759 12/05/09 01:20 apar Start_Parity_Run_14420__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304755 12/05/09 01:19 auto_snap Happex Charge Asym snapshot
  • 304754 12/05/09 01:19 auto_snap Hall A Tools snapshot
  • 304753 12/05/09 01:19 auto_snap Compton Polarimeter snapshot
  • 304752 12/05/09 01:19 auto_snap Beam Position and Current snapshot
  • 304751 12/05/09 01:19 auto_snap Accelerator Energies snapshot
  • 304748 12/05/09 01:17 apar End_of_Parity_Run_14419__CODA_Config=Pvdis
  • 304747 12/05/09 00:55 R. Subedi New golden runs
  • 304744 12/05/09 00:21 apar Start_Parity_Run_14419__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304743 12/05/09 00:13 apar End_of_Parity_Run_14418__CODA_Config=Pvdis
  • 304742 12/05/09 00:12 miham Swing Shift Summary (12/04/09)
  • 304741 12/05/09 00:11 apar Start_Parity_Run_14418__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304740 12/05/09 00:07 apar Start_Parity_Run_14417__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304739 12/05/09 00:05 apar Start_Parity_Run_14416__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304735 12/05/09 00:00 apar Start_Parity_Run_14415__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • Friday

  • 304730 12/04/09 23:58 apar End_of_Parity_Run_14414__CODA_Config=Pvdis
  • 304729 12/04/09 23:20 X. Zheng more on left gr2 and gr5 efficiency
  • 304728 12/04/09 23:02 apar Start_Parity_Run_14414__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304723 12/04/09 23:00 apar End_of_Parity_Run_14413__CODA_Config=Pvdis
  • 304722 12/04/09 22:58 miham Many Beam trips - Problem solved
  • 304721 12/04/09 22:46 X. Zheng left gr5 timing after removing the delay
  • 304720 12/04/09 22:46 liye target is stable
  • 304719 12/04/09 22:12 S. Sirca Please update golden run for HRSL
  • 304718 12/04/09 22:10 S. Sirca Raw rates at production running with 100uA on 20cm LD2, 4x4 raster
  • 304717 12/04/09 22:09 liye target is stable
  • 304716 12/04/09 22:08 miham Spot++ check during first production run
  • 304715 12/04/09 22:05 miham Beam current increased to 100uA
  • 304714 12/04/09 22:05 apar Start_Parity_Run_14413__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304711 12/04/09 22:01 miham VDC are turned OFF for production!
  • 304710 12/04/09 21:57 S. Sirca Q2/W plots after beam recovery (runs 25986, 5510)
  • 304707 12/04/09 21:50 miham Spot++ check on LD2 with (6x6)MCC
  • 304704 12/04/09 21:45 S. Sirca Raw rates with 2uA on 20cm LD2, 6x6 raster
  • 304701 12/04/09 21:39 miham Spot++ check with 2uA on 20cm LD2, 4x4 raster
  • 304698 12/04/09 21:36 S. Sirca Raw rates with 2uA on 20cm LD2, 4x4 raster
  • 304697 12/04/09 21:33 liye target->LD2
  • 304696 12/04/09 21:32 miham VDCs are ON!
  • 304695 12/04/09 21:19 miham Spot++ check on BeO (2x2)MCC - no tail with this raster size
  • 304694 12/04/09 21:19 auto_snap Left Arm HV snapshot
  • 304693 12/04/09 21:19 auto_snap Happex Charge Asym snapshot
  • 304692 12/04/09 21:19 auto_snap Hall A Tools snapshot
  • 304691 12/04/09 21:19 auto_snap Compton Polarimeter snapshot
  • 304690 12/04/09 21:19 auto_snap Beam Position and Current snapshot
  • 304689 12/04/09 21:19 auto_snap Accelerator Energies snapshot
  • 304688 12/04/09 21:17 S. Sirca Label of run 25982
  • 304685 12/04/09 21:06 miham Spot++ check on BeO (4x4)MCC
  • 304684 12/04/09 20:59 S. Sirca Tails on raster
  • 304681 12/04/09 20:52 liye target->BeO
  • 304680 12/04/09 20:15 miham BPM values
  • 304679 12/04/09 20:03 Kai/Dianchen add 16ns delay to prl1[30]a,prl1[30]b, prl1[31]a,prl1[31]b
  • 304678 12/04/09 19:45 R. Subedi Left hrs pvdis disr threshold, gr5 TS delay cable
  • 304677 12/04/09 19:04 X. Zheng this is a software bug - please take only T1 events for Q2 (VDC on) measurements unless instructed otherwise
  • 304676 12/04/09 19:01 Deng runlist summary for kinematics when LHRS angle 15 deg and momentum 3.66GeV/c
  • 304675 12/04/09 18:32 X. Zheng right arm gr4 has gone bad
  • 304674 12/04/09 18:24 deng One more plot for pion rate
  • 304673 12/04/09 18:23 Deng More plots for pion rates
  • 304672 12/04/09 18:19 Deng Pion rates for new kinematics, run 14412
  • 304670 12/04/09 18:13 X. Zheng left arm gr5 need to remove delay from TS
  • 304666 12/04/09 18:04 Kai Fig 1 is here.
  • 304665 12/04/09 17:57 Kai electron efficiency on the left arm during the whole experiment
  • 304664 12/04/09 17:53 Deng More plots for LHRS electron rates in New kinematics, run 14412
  • 304663 12/04/09 17:47 deng plots for electron rates in New kinematics, run 14412
  • 304662 12/04/09 17:21 miham Beam problems and Controlled Access
  • 304661 12/04/09 17:19 auto_snap Left Arm HV snapshot
  • 304660 12/04/09 17:18 auto_snap Happex Charge Asym snapshot
  • 304659 12/04/09 17:18 auto_snap Hall A Tools snapshot
  • 304658 12/04/09 17:18 auto_snap Compton Polarimeter snapshot
  • 304657 12/04/09 17:18 auto_snap Beam Position and Current snapshot
  • 304656 12/04/09 17:18 auto_snap Accelerator Energies snapshot
  • 304655 12/04/09 17:17 brads re: Valve won't open in North Linac -- could be > 1 hour
  • 304654 12/04/09 17:11 Deng target moved to Empty
  • 304653 12/04/09 16:51 Deng screen shot of Q2 and W for new kinematics
  • 304652 12/04/09 16:16 brads Valve won't open in North Linac -- expert looking at it
  • 304651 12/04/09 16:12 miham No beam
  • 304650 12/04/09 16:09 apar End_of_Parity_Run_14412__CODA_Config=Pvdis
  • 304647 12/04/09 16:06 M Friend Compton vertical scan
  • 304645 12/04/09 16:00 Deng day shift summary
  • 304644 12/04/09 15:58 Deng screen shot of hall A General tools
  • 304643 12/04/09 15:51 kbrider target screens
  • 304642 12/04/09 15:44 Deng Screen shot of HRS scaler rates for production (100uA)
  • 304641 12/04/09 15:42 apar Start_Parity_Run_14412__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=2.63,leftarm_theta=20,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304638 12/04/09 15:35 hoyoung Q2 scan results
  • 304637 12/04/09 15:34 Deng Beam Recovery Procedure
  • 304636 12/04/09 15:32 Deng VDC is off
  • 304631 12/04/09 15:20 hoyoung spot++ at LD2, 6x6(MCC) raster
  • 304628 12/04/09 15:16 hoyoung spot++ at LD2, 4x4(MCC) raster
  • 304625 12/04/09 15:02 hoyoung spot++ at BeO, 10 uA
  • 304621 12/04/09 14:28 brads Sweep finished, moving to Beam Permit
  • 304620 12/04/09 13:43 Ramesh/Kai Left hrs pvdis discriminator threshold etc.
  • 304619 12/04/09 13:41 Folts checklist
  • 304618 12/04/09 13:19 D.Wang slug number changed to 18 in SLUGNUMBER by hand
  • 304617 12/04/09 13:18 auto_snap Left Arm HV snapshot
  • 304616 12/04/09 13:18 auto_snap Happex Charge Asym snapshot
  • 304615 12/04/09 13:18 auto_snap Hall A Tools snapshot
  • 304614 12/04/09 13:18 auto_snap Compton Polarimeter snapshot
  • 304613 12/04/09 13:18 auto_snap Beam Position and Current snapshot
  • 304612 12/04/09 13:18 auto_snap Accelerator Energies snapshot
  • 304610 12/04/09 12:50 Deng Hall wiil be ready for sweeping in 10 minutes
  • 304609 12/04/09 12:32 deng screen shot of halla tools --LHRS momentun changed to 2.63GeV/c
  • 304608 12/04/09 12:06 brads LHRS dipole reset smoothly and is ramping back up
  • 304607 12/04/09 11:58 brads VDCs are ON
  • 304606 12/04/09 11:52 brads [Tech] LHRS dipole tripped
  • 304605 12/04/09 11:35 Deng LHRS spectrometer move is done within one hour!!!
  • 304604 12/04/09 10:41 deng restricted access now
  • 304603 12/04/09 10:16 brads Run coordinator daily summary
  • 304602 12/04/09 09:54 deng controlled access now
  • 304601 12/04/09 09:48 kbrider Target move to empty
  • 304600 12/04/09 09:44 hoyoung End of 25975 failed, daq restart
  • 304597 12/04/09 09:37 apar End_of_Parity_Run_14411__CODA_Config=Pvdis
  • 304595 12/04/09 09:18 auto_snap Left Arm HV snapshot
  • 304594 12/04/09 09:18 auto_snap Happex Charge Asym snapshot
  • 304593 12/04/09 09:18 auto_snap Hall A Tools snapshot
  • 304592 12/04/09 09:18 auto_snap Compton Polarimeter snapshot
  • 304591 12/04/09 09:18 auto_snap Beam Position and Current snapshot
  • 304590 12/04/09 09:17 auto_snap Accelerator Energies snapshot
  • 304589 12/04/09 09:08 apar Start_Parity_Run_14411__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304584 12/04/09 09:06 apar End_of_Parity_Run_14410__CODA_Config=Pvdis
  • 304583 12/04/09 08:21 kbrider target screens
  • 304582 12/04/09 08:11 apar Start_Parity_Run_14410__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304577 12/04/09 08:09 apar End_of_Parity_Run_14409__CODA_Config=Pvdis
  • 304576 12/04/09 08:02 ZAFAR AHMED Owl shift summary
  • 304574 12/04/09 07:59 R. Michaels spectrometer move at 9:30
  • 304573 12/04/09 07:15 R. Michaels reboot L-HRS DAQ
  • 304572 12/04/09 07:13 apar Start_Parity_Run_14409__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304567 12/04/09 07:09 apar End_of_Parity_Run_14408__CODA_Config=Pvdis
  • 304566 12/04/09 06:43 R. Michaels Coulombmeter : 50.5 C
  • 304563 12/04/09 06:13 apar Start_Parity_Run_14408__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304560 12/04/09 06:11 apar End_of_Parity_Run_14407__CODA_Config=Pvdis
  • 304559 12/04/09 05:17 auto_snap Left Arm HV snapshot
  • 304558 12/04/09 05:17 auto_snap Happex Charge Asym snapshot
  • 304557 12/04/09 05:17 auto_snap Hall A Tools snapshot
  • 304556 12/04/09 05:17 auto_snap Compton Polarimeter snapshot
  • 304555 12/04/09 05:17 auto_snap Beam Position and Current snapshot
  • 304554 12/04/09 05:17 auto_snap Accelerator Energies snapshot
  • 304553 12/04/09 05:15 apar Start_Parity_Run_14407__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304548 12/04/09 05:12 apar End_of_Parity_Run_14406__CODA_Config=Pvdis
  • 304547 12/04/09 05:01 R. Michaels Tools and beam screens
  • 304546 12/04/09 04:59 R. Michaels target screen capture
  • 304545 12/04/09 04:48 R. Michaels spot++
  • 304544 12/04/09 04:23 R. Michaels screen grabs of xscaler
  • 304543 12/04/09 04:16 apar Start_Parity_Run_14406__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304538 12/04/09 04:13 apar End_of_Parity_Run_14405__CODA_Config=Pvdis
  • 304537 12/04/09 04:05 R. Michaels odd rate peaks
  • 304536 12/04/09 04:01 R. Michaels synch glitch
  • 304535 12/04/09 03:17 apar Start_Parity_Run_14405__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304530 12/04/09 03:14 apar End_of_Parity_Run_14404__CODA_Config=Pvdis
  • 304529 12/04/09 02:39 R. Michaels Compton results, pretty clean
  • 304528 12/04/09 02:19 apar Start_Parity_Run_14404__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304523 12/04/09 02:16 apar End_of_Parity_Run_14403__CODA_Config=Pvdis
  • 304522 12/04/09 01:51 R. Michaels Compton reboot. new run 21392.
  • 304521 12/04/09 01:37 R. Michaels cavity mon update
  • 304520 12/04/09 01:20 apar Start_Parity_Run_14403__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304519 12/04/09 01:18 apar End_of_Parity_Run_14402__CODA_Config=Pvdis
  • 304518 12/04/09 01:17 auto_snap Left Arm HV snapshot
  • 304517 12/04/09 01:17 auto_snap Happex Charge Asym snapshot
  • 304516 12/04/09 01:17 auto_snap Hall A Tools snapshot
  • 304515 12/04/09 01:17 auto_snap Compton Polarimeter snapshot
  • 304514 12/04/09 01:17 auto_snap Beam Position and Current snapshot
  • 304513 12/04/09 01:17 auto_snap Accelerator Energies snapshot
  • 304508 12/04/09 01:03 R. Michaels target pressure excursions
  • 304507 12/04/09 00:54 R. Michaels dithering turned on
  • 304506 12/04/09 00:28 deng Lumi HV adjusted
  • 304505 12/04/09 00:23 apar Start_Parity_Run_14402__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304504 12/04/09 00:22 R. Michaels run 14401 no good, reboot CODA
  • 304503 12/04/09 00:13 apar Start_Parity_Run_14401__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304499 12/04/09 00:11 apar End_of_Parity_Run_14400__CODA_Config=Pvdis
  • 304497 12/04/09 00:03 miham Swing shift Summary (12/03/09)
  • Thursday

  • 304495 12/03/09 23:58 miham Spot++ check during run #25965
  • 304494 12/03/09 23:47 deng explanation why we did twice flipper for slug 17, HWP IN
  • 304493 12/03/09 23:28 LeRose Target Screens
  • 304492 12/03/09 23:21 apar Start_Parity_Run_14400__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304489 12/03/09 23:19 Deng screen shot of happex feedback monitor for half wave plate IN
  • 304488 12/03/09 23:13 apar Halfwave plate IN, Slug 17
  • 304487 12/03/09 23:12 apar End_of_Parity_Run_14399__CODA_Config=Pvdis
  • 304484 12/03/09 23:11 sirca/miham Output from flipper script
  • 304483 12/03/09 23:09 A. Camsonne Compton photon detector position
  • 304482 12/03/09 23:07 miham HWP changed to IN
  • 304481 12/03/09 23:06 apar Start_Parity_Run_14399__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304478 12/03/09 23:03 apar Halfwave plate IN, Slug 17
  • 304477 12/03/09 23:02 brads Run plan for OWL
  • 304474 12/03/09 23:00 apar End_of_Parity_Run_14398__CODA_Config=Pvdis
  • 304473 12/03/09 22:57 miham Problems with the beam
  • 304472 12/03/09 22:42 A. Camsonne Compton EPICS rebooted
  • 304471 12/03/09 22:33 miham Problems with Compton
  • 304470 12/03/09 22:25 apar Start_Parity_Run_14398__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304469 12/03/09 22:22 apar End_of_Parity_Run_14397__CODA_Config=Pvdis
  • 304459 12/03/09 21:22 apar End_of_Parity_Run_14396__CODA_Config=Pvdis
  • 304458 12/03/09 21:17 auto_snap Left Arm HV snapshot
  • 304457 12/03/09 21:17 auto_snap Happex Charge Asym snapshot
  • 304456 12/03/09 21:17 auto_snap Hall A Tools snapshot
  • 304455 12/03/09 21:16 auto_snap Compton Polarimeter snapshot
  • 304454 12/03/09 21:16 auto_snap Beam Position and Current snapshot
  • 304453 12/03/09 21:16 auto_snap Accelerator Energies snapshot
  • 304452 12/03/09 20:32 LeRose Target Screens
  • 304451 12/03/09 20:26 apar Start_Parity_Run_14396__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304446 12/03/09 20:24 apar End_of_Parity_Run_14395__CODA_Config=Pvdis
  • 304445 12/03/09 19:51 M Friend Compton cavity power optimized
  • 304444 12/03/09 19:43 miham Spot++ check during run #25960
  • 304443 12/03/09 19:28 apar Start_Parity_Run_14395__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304438 12/03/09 19:26 apar End_of_Parity_Run_14394__CODA_Config=Pvdis
  • 304437 12/03/09 18:32 miham Beam is back
  • 304436 12/03/09 18:31 apar Start_Parity_Run_14394__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304433 12/03/09 18:18 miham No beam for a while
  • 304432 12/03/09 18:14 apar End_of_Parity_Run_14393__CODA_Config=Pvdis
  • 304429 12/03/09 17:57 apar Start_Parity_Run_14393__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304424 12/03/09 17:53 apar End_of_Parity_Run_14392__CODA_Config=Pvdis
  • 304423 12/03/09 17:48 R. Subedi New golden runs, cfg files changed
  • 304422 12/03/09 17:16 auto_snap Left Arm HV snapshot
  • 304421 12/03/09 17:16 auto_snap Happex Charge Asym snapshot
  • 304420 12/03/09 17:16 auto_snap Hall A Tools snapshot
  • 304419 12/03/09 17:16 auto_snap Compton Polarimeter snapshot
  • 304418 12/03/09 17:16 auto_snap Beam Position and Current snapshot
  • 304417 12/03/09 17:16 auto_snap Accelerator Energies snapshot
  • 304416 12/03/09 16:58 apar Start_Parity_Run_14392__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304412 12/03/09 16:55 apar End_of_Parity_Run_14391__CODA_Config=Pvdis
  • 304410 12/03/09 16:44 R. Subedi Electron and pion asymmetry in a new kin
  • 304409 12/03/09 16:04 Holmstrom Day Shift Summary
  • 304407 12/03/09 15:59 apar Start_Parity_Run_14391__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304403 12/03/09 15:56 apar End_of_Parity_Run_14390__CODA_Config=Pvdis
  • 304401 12/03/09 15:46 G. Russo strip charts and target interface
  • 304400 12/03/09 15:39 M Friend Compton rates look amazing!
  • 304399 12/03/09 15:02 apar Start_Parity_Run_14390__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=15,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304394 12/03/09 14:59 apar End_of_Parity_Run_14389__CODA_Config=Pvdis
  • 304393 12/03/09 14:18 hoyoung Q2 data analyzed
  • 304392 12/03/09 14:12 G. Russo strip charts
  • 304390 12/03/09 14:11 A. Camsonne Compton electron detector back to garage position
  • 304386 12/03/09 14:04 apar Start_Parity_Run_14389__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304383 12/03/09 14:00 G. Russo target screen
  • 304382 12/03/09 13:58 R. Subedi VDCs off for production running
  • 304381 12/03/09 13:56 A. Camsonne Odd Compton electron detector background dependence on beam current
  • 304377 12/03/09 13:49 hoyoung spot++ at LD2, 6x6(MCC)
  • 304376 12/03/09 13:45 hoyoung Run # 25952 and 5484 is Q2 data
  • 304371 12/03/09 13:40 hoyoung spot++ for LD2, 4x4(MCC)
  • 304368 12/03/09 13:33 brads Run coordinator daily summary
  • 304367 12/03/09 13:24 hoyoung Spot++ at BeO
  • 304365 12/03/09 13:21 G. Russo target screen
  • 304363 12/03/09 13:19 Holmstrom HARP Scan
  • 304362 12/03/09 13:16 auto_snap Left Arm HV snapshot
  • 304361 12/03/09 13:16 auto_snap Happex Charge Asym snapshot
  • 304360 12/03/09 13:16 auto_snap Hall A Tools snapshot
  • 304359 12/03/09 13:16 auto_snap Compton Polarimeter snapshot
  • 304358 12/03/09 13:16 auto_snap Beam Position and Current snapshot
  • 304357 12/03/09 13:16 auto_snap Accelerator Energies snapshot
  • 304356 12/03/09 13:12 A. Camsonne Compton electron detector threshold scan 2 uA
  • 304355 12/03/09 13:06 A. Camsonne Compton electron detector current position 25 mm 2 uA
  • 304354 12/03/09 13:00 R. Subedi Slug 16 starts, IHWP is still OUT
  • 304353 12/03/09 12:49 apar Halfwave plate OUT, Slug 16
  • 304352 12/03/09 11:59 A. Camsonne Compton electron background threshold scan
  • 304351 12/03/09 11:25 G.Russo Target empty
  • 304350 12/03/09 10:57 Holmstrom Angle Change
  • 304349 12/03/09 10:25 deng summary for new kinematics with LHRS 14deg and momentum 3.66Gev/c
  • 304346 12/03/09 10:04 apar End_of_Parity_Run_14388__CODA_Config=Pvdis
  • 304345 12/03/09 09:52 G .Russo target
  • 304344 12/03/09 09:51 apar Start_Parity_Run_14388__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304339 12/03/09 09:50 apar End_of_Parity_Run_14387__CODA_Config=Pvdis
  • 304338 12/03/09 09:15 auto_snap Left Arm HV snapshot
  • 304337 12/03/09 09:15 auto_snap Happex Charge Asym snapshot
  • 304336 12/03/09 09:15 auto_snap Hall A Tools snapshot
  • 304335 12/03/09 09:15 auto_snap Compton Polarimeter snapshot
  • 304334 12/03/09 09:15 auto_snap Beam Position and Current snapshot
  • 304333 12/03/09 09:15 auto_snap Accelerator Energies snapshot
  • 304332 12/03/09 08:57 apar Start_Parity_Run_14387__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304327 12/03/09 08:52 apar End_of_Parity_Run_14386__CODA_Config=Pvdis
  • 304326 12/03/09 08:49 A. Camsonne Compton tuning
  • 304325 12/03/09 08:43 brads LHRS move to 15 degrees scheduled for 10am
  • 304324 12/03/09 08:31 G, Russo target screen
  • 304323 12/03/09 08:26 A. Camsonne Compton electron detector current at 100 uA 25 mm from beam
  • 304321 12/03/09 08:09 A. Camsonne Compton electron detector moved to position 25 mm
  • 304320 12/03/09 08:03 A. Camsonne Compton electron detector moved to position 20 mm
  • 304319 12/03/09 08:02 silwal owl shift summary
  • 304317 12/03/09 07:56 apar Start_Parity_Run_14386__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304313 12/03/09 07:41 apar End_of_Parity_Run_14385__CODA_Config=Pvdis
  • 304312 12/03/09 07:23 deng expert shift check-- no electron wide hit for LHRS grp 5 gets much larger
  • 304311 12/03/09 06:46 apar Start_Parity_Run_14385__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304306 12/03/09 06:42 apar End_of_Parity_Run_14384__CODA_Config=Pvdis
  • 304305 12/03/09 05:47 apar Start_Parity_Run_14384__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304302 12/03/09 05:45 apar Halfwave plate OUT, Slug 15
  • 304301 12/03/09 05:45 silwal change IHWP from IN to OUT
  • 304298 12/03/09 05:39 apar End_of_Parity_Run_14383__CODA_Config=Pvdis
  • 304297 12/03/09 05:15 auto_snap Left Arm HV snapshot
  • 304296 12/03/09 05:15 auto_snap Happex Charge Asym snapshot
  • 304295 12/03/09 05:15 auto_snap Hall A Tools snapshot
  • 304294 12/03/09 05:15 auto_snap Compton Polarimeter snapshot
  • 304293 12/03/09 05:15 auto_snap Beam Position and Current snapshot
  • 304292 12/03/09 05:15 auto_snap Accelerator Energies snapshot
  • 304291 12/03/09 04:43 apar Start_Parity_Run_14383__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304290 12/03/09 04:38 apar Start_Parity_Run_14382__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304285 12/03/09 04:35 apar End_of_Parity_Run_14381__CODA_Config=Pvdis
  • 304284 12/03/09 04:19 M Friend Compton rates/vertical scan
  • 304283 12/03/09 04:03 A. Giusa grab of target screens
  • 304282 12/03/09 03:49 M Friend Compton rates
  • 304281 12/03/09 03:41 apar Start_Parity_Run_14381__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304277 12/03/09 03:35 apar End_of_Parity_Run_14380__CODA_Config=Pvdis
  • 304276 12/03/09 03:30 M Friend Compton laser alignment
  • 304275 12/03/09 03:08 silwal stable beam @ 90uA
  • 304274 12/03/09 02:39 apar Start_Parity_Run_14380__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304269 12/03/09 02:36 apar End_of_Parity_Run_14379__CODA_Config=Pvdis
  • 304268 12/03/09 02:14 silwal back to lots of beam trips!
  • 304267 12/03/09 02:00 silwal compton CODA rebooted
  • 304266 12/03/09 01:57 silwal compton mess
  • 304265 12/03/09 01:54 silwal beam stable @ 110 uA
  • 304264 12/03/09 01:42 apar Start_Parity_Run_14379__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304261 12/03/09 01:41 silwal beam back @ 100uA
  • 304258 12/03/09 01:39 apar End_of_Parity_Run_14378__CODA_Config=Pvdis
  • 304257 12/03/09 01:15 auto_snap Left Arm HV snapshot
  • 304256 12/03/09 01:15 auto_snap Happex Charge Asym snapshot
  • 304255 12/03/09 01:15 auto_snap Hall A Tools snapshot
  • 304254 12/03/09 01:15 auto_snap Compton Polarimeter snapshot
  • 304253 12/03/09 01:14 auto_snap Beam Position and Current snapshot
  • 304252 12/03/09 01:14 auto_snap Accelerator Energies snapshot
  • 304251 12/03/09 01:11 silwal MCC taking beam away for 15-20 minutes; they are looking at the cause of too many beam trips
  • 304250 12/03/09 01:10 deng Q2 and W plots for new kinematics
  • 304249 12/03/09 00:58 silwal too many beam trips @ 110uA, so asked for 100uA
  • 304248 12/03/09 00:53 deng HRS scaler rates for production
  • 304246 12/03/09 00:45 apar Start_Parity_Run_14378__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production 110uA,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=14.0,parity_feedback=on,parity_ihwp=in,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304244 12/03/09 00:28 ellie/deng Beam Checkout Procedure
  • 304241 12/03/09 00:20 deng screen shot of RHRS/LHRS scalers
  • 304238 12/03/09 00:15 A. Giusa Grab of target screens.
  • 304235 12/03/09 00:14 ellie Swing Shift Summary
  • 304228 12/03/09 00:08 ellie 6x6 MCC raster spot++
  • 304224 12/03/09 00:03 ellie 4x4 MCC raster spot++
  • 304222 12/03/09 00:02 liye target->BeO->LD2
  • Wednesday

  • 304220 12/02/09 23:51 ellie 4x4 MCC spot++
  • 304217 12/02/09 23:36 ellie Back to beam check-out.
  • 304216 12/02/09 23:29 ellie And HV Trip again...
  • 304215 12/02/09 23:28 ellie LeCroy HV Alarm again
  • 304214 12/02/09 23:26 K.Pan/J.Huan change kinematics setting in podd DB file and add invariant mass "W" plot into onlinegui
  • 304213 12/02/09 23:21 ellie Well, not really...
  • 304212 12/02/09 23:16 ellie LeCroy HV tripped, Channel 7. Turned off and back on fixed it
  • 304211 12/02/09 23:09 ellie Beginning beam check-out procedure
  • 304210 12/02/09 22:53 liye target
  • 304209 12/02/09 22:53 ellie Hall A Tools Screen Grab
  • 304208 12/02/09 22:46 O.Glamazdin Moller results
  • 304207 12/02/09 22:27 ellie Spin Dance and Moller complete, MCC preparing for Production
  • 304206 12/02/09 21:39 ellie Spin Dance in progress, going smoothly
  • 304205 12/02/09 21:14 auto_snap Left Arm HV snapshot
  • 304204 12/02/09 21:14 auto_snap Happex Charge Asym snapshot
  • 304203 12/02/09 21:14 auto_snap Hall A Tools snapshot
  • 304202 12/02/09 21:14 auto_snap Compton Polarimeter snapshot
  • 304201 12/02/09 21:14 auto_snap Beam Position and Current snapshot
  • 304200 12/02/09 21:14 auto_snap Accelerator Energies snapshot
  • 304199 12/02/09 20:51 adaq test
  • 304198 12/02/09 17:36 ZAFAR AHMED Cavity Pregains Reduced by 1
  • 304197 12/02/09 17:28 R. Subedi Delay cable added at noon today was 200 ns long.
  • 304196 12/02/09 17:16 brads LHRS maximum P0 is 3.66 GeV/c
  • 304195 12/02/09 17:14 auto_snap Left Arm HV snapshot
  • 304194 12/02/09 17:14 auto_snap Happex Charge Asym snapshot
  • 304193 12/02/09 17:14 auto_snap Hall A Tools snapshot
  • 304192 12/02/09 17:14 auto_snap Compton Polarimeter snapshot
  • 304191 12/02/09 17:14 auto_snap Beam Position and Current snapshot
  • 304190 12/02/09 17:14 auto_snap Accelerator Energies snapshot
  • 304189 12/02/09 16:39 ellie Going back to Controlled Access for Left Q2 problem
  • 304188 12/02/09 16:32 ellie LHRS Q2 turned off, Over Temperature Fault
  • 304187 12/02/09 16:31 ellie Changing LHRS P0 to 4.00
  • 304186 12/02/09 16:27 ellie IHWP set to IN
  • 304185 12/02/09 16:26 apar Halfwave plate IN, Slug 14
  • 304182 12/02/09 16:19 liye target->empty
  • 304178 12/02/09 16:07 D.Wang 200ns delay cable for GC removed
  • 304177 12/02/09 16:07 bogdanw shift summary
  • 304176 12/02/09 16:07 M Friend Compton photon collimator
  • 304175 12/02/09 16:02 brads LHRS moved to 14 degrees
  • 304174 12/02/09 15:57 Moffit Put the runnumber back into the panguin/prompt gui.
  • 304173 12/02/09 15:23 Kai summary of random coincident runs
  • 304172 12/02/09 14:53 apar End_of_Parity_Run_14377__CODA_Config=Pvdis
  • 304169 12/02/09 14:39 apar Start_Parity_Run_14377__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=high current random coincidence run, VDC off, FADC on,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304168 12/02/09 14:39 bogdanw Hall A tools
  • 304167 12/02/09 14:38 apar End_of_Parity_Run_14376__CODA_Config=Pvdis
  • 304164 12/02/09 14:01 X. Zheng about random coincidence rate
  • 304159 12/02/09 13:50 D.Wang global electron rate for radon coinc run
  • 304158 12/02/09 13:48 M Friend Compton photon beam position
  • 304153 12/02/09 13:43 apar Start_Parity_Run_14376__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=high current random coincidence run, VDC off, FADC on,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304150 12/02/09 13:32 bogdanw Hall ATools VDC ON
  • 304149 12/02/09 13:30 bogdanw Tools
  • 304144 12/02/09 13:18 apar End_of_Parity_Run_14375__CODA_Config=Pvdis
  • 304141 12/02/09 13:14 auto_snap Left Arm HV snapshot
  • 304140 12/02/09 13:14 auto_snap Happex Charge Asym snapshot
  • 304139 12/02/09 13:13 auto_snap Hall A Tools snapshot
  • 304138 12/02/09 13:13 auto_snap Compton Polarimeter snapshot
  • 304137 12/02/09 13:13 auto_snap Beam Position and Current snapshot
  • 304136 12/02/09 13:13 auto_snap Accelerator Energies snapshot
  • 304135 12/02/09 13:06 apar Start_Parity_Run_14375__CODA_Config=Pvdis_Run_type=Calibration,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=low current random coincidence run,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304132 12/02/09 13:05 Kai start take data with low current, VDC ON, FADC off
  • 304131 12/02/09 13:00 R. Subedi Tagger disabled
  • 304130 12/02/09 12:57 D.Wang/R.Sub added 200ns delay to GC input of VETO circuit
  • 304129 12/02/09 12:47 M Friend Compton photon beam position
  • 304128 12/02/09 12:39 Ramesh /Dian left scaler 3 and 4 while in Hall
  • 304127 12/02/09 12:38 Rames/Dianch Scaler 1 and 2 (left) while in Hall
  • 304126 12/02/09 12:14 R. Michaels cavities ok at 90 uA after gain adjustment
  • 304125 12/02/09 12:11 R. Subedi Escorted access
  • 304124 12/02/09 12:09 brads Run coordinator daily summary
  • 304123 12/02/09 12:05 R. Subedi Tagger ON
  • 304120 12/02/09 12:02 apar End_of_Parity_Run_14374__CODA_Config=Pvdis
  • 304119 12/02/09 11:23 apar Start_Parity_Run_14374__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304114 12/02/09 11:20 apar End_of_Parity_Run_14373__CODA_Config=Pvdis
  • 304113 12/02/09 11:17 Bob/Zafar Cavity Mon Status
  • 304112 12/02/09 11:15 A. Camsonne Compton photon detector looks better with new collimator
  • 304111 12/02/09 10:24 apar Start_Parity_Run_14373__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304106 12/02/09 10:22 apar End_of_Parity_Run_14372__CODA_Config=Pvdis
  • 304105 12/02/09 09:56 O.Glamazdin Beam condition
  • 304104 12/02/09 09:26 apar Start_Parity_Run_14372__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304099 12/02/09 09:23 apar End_of_Parity_Run_14371__CODA_Config=Pvdis
  • 304098 12/02/09 09:13 auto_snap Left Arm HV snapshot
  • 304097 12/02/09 09:13 auto_snap Happex Charge Asym snapshot
  • 304096 12/02/09 09:13 auto_snap Hall A Tools snapshot
  • 304095 12/02/09 09:13 auto_snap Compton Polarimeter snapshot
  • 304094 12/02/09 09:13 auto_snap Beam Position and Current snapshot
  • 304093 12/02/09 09:13 auto_snap Accelerator Energies snapshot
  • 304092 12/02/09 08:27 apar Start_Parity_Run_14371__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304089 12/02/09 08:25 D.Wang IHWP OUT
  • 304088 12/02/09 08:24 apar Halfwave plate OUT, Slug 13
  • 304085 12/02/09 08:21 apar End_of_Parity_Run_14370__CODA_Config=Pvdis
  • 304084 12/02/09 08:13 D.Wang Hall A Tools
  • 304082 12/02/09 07:56 Holt snapshot of target screens
  • 304081 12/02/09 07:56 D.Wang Owl shift summary
  • 304080 12/02/09 07:26 apar Start_Parity_Run_14370__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304075 12/02/09 07:21 apar End_of_Parity_Run_14369__CODA_Config=Pvdis
  • 304074 12/02/09 06:44 G. Russo target screen
  • 304073 12/02/09 06:25 apar Start_Parity_Run_14369__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304070 12/02/09 06:22 D.Wang Happex Coda crashed
  • 304067 12/02/09 06:10 apar Start_Parity_Run_14368__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304064 12/02/09 06:09 Kai new spot++ result shows clean shape
  • 304061 12/02/09 06:04 apar End_of_Parity_Run_14367__CODA_Config=Pvdis
  • 304060 12/02/09 06:02 G. Russo target screen
  • 304059 12/02/09 05:37 R. Subedi Unclean spot is due to beam trip during first 50k events
  • 304058 12/02/09 05:13 auto_snap Left Arm HV snapshot
  • 304057 12/02/09 05:13 auto_snap Happex Charge Asym snapshot
  • 304056 12/02/09 05:13 auto_snap Hall A Tools snapshot
  • 304055 12/02/09 05:13 auto_snap Compton Polarimeter snapshot
  • 304054 12/02/09 05:13 auto_snap Beam Position and Current snapshot
  • 304053 12/02/09 05:12 auto_snap Accelerator Energies snapshot
  • 304052 12/02/09 05:08 apar Start_Parity_Run_14367__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304047 12/02/09 05:05 apar End_of_Parity_Run_14366__CODA_Config=Pvdis
  • 304046 12/02/09 04:21 G, Russo target screen
  • 304044 12/02/09 04:09 apar Start_Parity_Run_14366__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304040 12/02/09 04:04 apar End_of_Parity_Run_14365__CODA_Config=Pvdis
  • 304039 12/02/09 03:52 Kai I called MCC to confirm if we are in the transverse polarization mode
  • 304038 12/02/09 03:18 Kai global asymmetry in transverse mode
  • 304037 12/02/09 03:17 Kai Global rate in 1st production under transverse mode
  • 304036 12/02/09 03:08 apar Start_Parity_Run_14365__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304031 12/02/09 03:04 apar End_of_Parity_Run_14364__CODA_Config=Pvdis
  • 304030 12/02/09 02:56 Kai unclean plot of spot++ under production
  • 304029 12/02/09 02:35 Kai scaler: Right3,4
  • 304028 12/02/09 02:34 Kai scaler: Right 1,2
  • 304027 12/02/09 02:33 Kai scaler: Left 3,4
  • 304026 12/02/09 02:30 G. Russo target
  • 304025 12/02/09 02:19 Kai scaler in transverse mode: Left 1,2
  • 304024 12/02/09 02:13 G. Russo target screen
  • 304023 12/02/09 02:11 apar Start_Parity_Run_14364__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=production,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 304020 12/02/09 01:46 Kai turn off VDC, and go to 100uA
  • 304019 12/02/09 01:43 Kai 2uA, 6*6 MCC
  • 304016 12/02/09 01:38 Kai spot++, 2uA, 4*4 MCC
  • 304013 12/02/09 01:12 auto_snap Left Arm HV snapshot
  • 304012 12/02/09 01:12 auto_snap Happex Charge Asym snapshot
  • 304011 12/02/09 01:12 auto_snap Hall A Tools snapshot
  • 304010 12/02/09 01:12 auto_snap Compton Polarimeter snapshot
  • 304009 12/02/09 01:12 auto_snap Beam Position and Current snapshot
  • 304008 12/02/09 01:12 auto_snap Accelerator Energies snapshot
  • 304006 12/02/09 00:02 ellie Swing Shift Summary
  • Tuesday

  • 304005 12/01/09 23:48 ellie Beam returned until midnight-ish, returning to beam check-out procedure. Almost made it to production before beam is gone!
  • 304004 12/01/09 23:41 ellie MBSY1C is in Red on Hall A Tools
  • 303997 12/01/09 23:25 hoyoung spot++ with raster 4x4
  • 303994 12/01/09 23:20 hoyoung spot++ for LD2
  • 303991 12/01/09 23:06 brads RF seperator problem
  • 303990 12/01/09 23:05 ellie Beam Checkout Prodecure - Interrupted
  • 303989 12/01/09 23:00 Armstrong Target screenshots
  • 303988 12/01/09 22:51 ellie MCC givith, MCC taketh the beam away.
  • 303987 12/01/09 22:46 hoyoung LD2 spot++
  • 303984 12/01/09 22:42 hoyoung Trigger rate of 2uA at LD2
  • 303983 12/01/09 22:30 Armstrong moving to LD2 target
  • 303982 12/01/09 22:29 hoyoung spot++
  • 303977 12/01/09 22:19 Armstrong taregt to BeO
  • 303976 12/01/09 22:01 ellie MCC is going to start giving us CW beam
  • 303975 12/01/09 21:31 ellie RHRS Dipole giving tiny trips then going back to normal
  • 303974 12/01/09 21:12 auto_snap Left Arm HV snapshot
  • 303973 12/01/09 21:12 auto_snap Happex Charge Asym snapshot
  • 303972 12/01/09 21:12 auto_snap Hall A Tools snapshot
  • 303971 12/01/09 21:12 auto_snap Compton Polarimeter snapshot
  • 303970 12/01/09 21:12 auto_snap Beam Position and Current snapshot
  • 303969 12/01/09 21:12 auto_snap Accelerator Energies snapshot
  • 303968 12/01/09 18:37 D.Wang IHWP IN
  • 303967 12/01/09 18:26 ellie IHWP set to IN
  • 303966 12/01/09 18:24 apar Halfwave plate OUT, Slug 12
  • 303963 12/01/09 17:57 apar End_of_Parity_Run_14363__CODA_Config=Pvdis
  • 303962 12/01/09 17:57 apar Start_Parity_Run_14363__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=NO BEAM,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303957 12/01/09 17:55 apar End_of_Parity_Run_14362__CODA_Config=Pvdis
  • 303956 12/01/09 17:54 ellie MCC Called, they're getting ready so we're going to Beam Permit
  • 303955 12/01/09 17:37 Armstrong Target: move to empty position
  • 303954 12/01/09 17:33 brads Red LHe flow readback for RHRS Q1 is OK
  • 303953 12/01/09 17:12 auto_snap Left Arm HV snapshot
  • 303952 12/01/09 17:12 auto_snap Happex Charge Asym snapshot
  • 303951 12/01/09 17:11 auto_snap Hall A Tools snapshot
  • 303950 12/01/09 17:11 auto_snap Compton Polarimeter snapshot
  • 303949 12/01/09 17:11 auto_snap Beam Position and Current snapshot
  • 303948 12/01/09 17:11 auto_snap Accelerator Energies snapshot
  • 303947 12/01/09 17:08 folts cryo /tech
  • 303946 12/01/09 17:04 ellie Magnets have been cycled. Left P0=3.660, Right P0=2.630
  • 303945 12/01/09 16:59 apar Start_Parity_Run_14362__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=NO BEAM,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303940 12/01/09 16:57 apar End_of_Parity_Run_14361__CODA_Config=Pvdis
  • 303938 12/01/09 16:04 R. Subedi Cosmic runs
  • 303937 12/01/09 16:02 jpchen target: alarm for Loop 1 temperature sensor #1 disabled
  • 303936 12/01/09 16:02 apar Start_Parity_Run_14361__CODA_Config=Pvdis_Run_type=Cosmics,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=NO BEAM,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303931 12/01/09 15:54 apar End_of_Parity_Run_14360__CODA_Config=Pvdis
  • 303930 12/01/09 15:53 Holt yellow target alarm
  • 303929 12/01/09 14:36 brads Firmware upgraded for Moller FADC board
  • 303928 12/01/09 13:11 auto_snap Left Arm HV snapshot
  • 303927 12/01/09 13:11 auto_snap Happex Charge Asym snapshot
  • 303926 12/01/09 13:11 auto_snap Hall A Tools snapshot
  • 303925 12/01/09 13:11 auto_snap Compton Polarimeter snapshot
  • 303924 12/01/09 13:11 auto_snap Beam Position and Current snapshot
  • 303923 12/01/09 13:11 auto_snap Accelerator Energies snapshot
  • 303922 12/01/09 12:56 brads Vertical polarization setup is complete; accel vacuum problem fixed, beam in early swing (hoping!)
  • 303921 12/01/09 12:54 Musson/Ahmed Cavity BPM Receivers
  • 303920 12/01/09 12:27 brads Run coordinator daily summary
  • 303919 12/01/09 11:15 brads Situation update
  • 303918 12/01/09 10:56 A. Camsonne Compton photon table check
  • 303913 12/01/09 09:11 auto_snap Left Arm HV snapshot
  • 303912 12/01/09 09:11 auto_snap Happex Charge Asym snapshot
  • 303911 12/01/09 09:11 auto_snap Hall A Tools snapshot
  • 303910 12/01/09 09:11 auto_snap Compton Polarimeter snapshot
  • 303909 12/01/09 09:11 auto_snap Beam Position and Current snapshot
  • 303908 12/01/09 09:11 auto_snap Accelerator Energies snapshot
  • 303907 12/01/09 08:33 Kai called MCC for update
  • 303906 12/01/09 07:57 Holt snapshot of target screens
  • 303905 12/01/09 07:56 deng screen short of halla tools
  • 303904 12/01/09 07:55 Deng owl shift summary
  • 303903 12/01/09 07:29 apar Start_Parity_Run_14360__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=NO BEAM,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303898 12/01/09 07:26 apar End_of_Parity_Run_14359__CODA_Config=Pvdis
  • 303897 12/01/09 06:30 apar Start_Parity_Run_14359__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=NO BEAM,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303892 12/01/09 06:26 apar End_of_Parity_Run_14358__CODA_Config=Pvdis
  • 303891 12/01/09 05:31 apar Start_Parity_Run_14358__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production.,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303886 12/01/09 05:28 apar End_of_Parity_Run_14357__CODA_Config=Pvdis
  • 303885 12/01/09 05:10 auto_snap Left Arm HV snapshot
  • 303884 12/01/09 05:10 auto_snap Happex Charge Asym snapshot
  • 303883 12/01/09 05:10 auto_snap Hall A Tools snapshot
  • 303882 12/01/09 05:10 auto_snap Compton Polarimeter snapshot
  • 303881 12/01/09 05:10 auto_snap Beam Position and Current snapshot
  • 303880 12/01/09 05:10 auto_snap Accelerator Energies snapshot
  • 303879 12/01/09 04:32 apar Start_Parity_Run_14357__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production.,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303874 12/01/09 04:29 apar End_of_Parity_Run_14356__CODA_Config=Pvdis
  • 303873 12/01/09 03:33 apar Start_Parity_Run_14356__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production.,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303868 12/01/09 03:31 apar End_of_Parity_Run_14355__CODA_Config=Pvdis
  • 303867 12/01/09 02:35 G. Russo target screen
  • 303866 12/01/09 02:35 apar Start_Parity_Run_14355__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production.,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303865 12/01/09 02:28 apar Start_Parity_Run_14354__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Production.,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303860 12/01/09 02:25 apar End_of_Parity_Run_14353__CODA_Config=Pvdis
  • 303859 12/01/09 01:29 apar Start_Parity_Run_14353__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Left arm only,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303855 12/01/09 01:26 apar End_of_Parity_Run_14352__CODA_Config=Pvdis
  • 303854 12/01/09 01:10 auto_snap Left Arm HV snapshot
  • 303853 12/01/09 01:10 auto_snap Happex Charge Asym snapshot
  • 303852 12/01/09 01:10 auto_snap Hall A Tools snapshot
  • 303851 12/01/09 01:10 auto_snap Compton Polarimeter snapshot
  • 303850 12/01/09 01:10 auto_snap Beam Position and Current snapshot
  • 303849 12/01/09 01:10 auto_snap Accelerator Energies snapshot
  • 303848 12/01/09 00:56 G. Russo target screen
  • 303847 12/01/09 00:34 apar Start_Parity_Run_14352__CODA_Config=Pvdis_Run_type=Production,beam_energy=6.08,beam_polarization=87,beam_raster=on,comment_text=Left arm only,dithering_magnet=on,dithering_magnets=,leftarm_p=3.66,leftarm_theta=12.9,parity_feedback=on,parity_ihwp=out,rightarm_p=2.63,rightarm_theta=20,target_fanspeed=60 Hz,target_type=20cm LD2
  • 303842 12/01/09 00:27 apar End_of_Parity_Run_14351__CODA_Config=Pvdis
  • 303839 12/01/09 00:11 G. Russo target screen
  • 303838 12/01/09 00:10 ellie Hall A Tools Screen Grab
  • 303837 12/01/09 00:03 ellie` Swing Shift Summary